### v3

# COADVISE: Covariate Adjustment with Variable Selection in Randomized Controlled Trials

Abstract
Adjusting for covariates in randomized controlled trials can enhance the credibility and efficiency of treatment effect estimation. However, handling numerous covariates and their complex (non-linear) transformations poses a challenge. Motivated by the case study of the Best Apnea Interventions for Research (BestAIR) trial data from the National Sleep Research Resource (NSRR), where the number of covariates ($p=114$) is comparable to the sample size ($N=196$), we propose a principled Covariate Adjustment with Variable Selection (COADVISE) framework. COADVISE enables variable selection for covariates most relevant to the outcome while accommodating both linear and nonlinear adjustments. This framework ensures consistent estimates with improved efficiency over unadjusted estimators and provides robust variance estimation, even under outcome model misspecification. We demonstrate efficiency gains through theoretical analysis, extensive simulations, and a re-analysis of the BestAIR trial data to compare alternative variable selection strategies, offering cautionary recommendations. A user-friendly R package, Coadvise, is available to facilitate practical implementation.
Key words: Complete randomization, Covariate balance, Efficiency gain, Model misspecification; High-dimensional data
1 Introduction
1.1 Background
Randomized controlled trials (RCTs) are widely regarded as the gold standard for evaluating the effects of treatments or interventions (Neyman, 1923; Rubin, 1974). Through randomization, covariates are balanced on average across treatment groups, enabling unbiased estimation of treatment effects. However, in practice, randomization does not guarantee perfect balance in finite samples, particularly when dealing with a large number of covariates. Such covariate imbalances can obscure the interpretation of trial outcomes and introduce accidental biases. Incorporating covariates into the analysis can enhance both the interpretability and efficiency of RCT results. While covariate balance can be pursued during both the design (Fisher, 1926; Taves, 1974; Pocock and Simon, 1975; Rosenberger and Sverdlov, 2008; Hu and Hu, 2012; Hu et al., 2014; Morgan and Rubin, 2012; Li et al., 2018b; Zhao and Ding, 2024) and analysis phases of an RCT, this paper focuses on methods for covariate adjustment during the analysis of already-conducted trials.
Regulatory agencies, including the U.S. Food and Drug Administration (FDA) and the European Medicines Agency (EMA), have emphasized the importance of covariate adjustment in RCTs. The FDA recently issued guidelines encouraging its use to improve the precision of treatment effect estimates (FDA, 2023). The EMA’s 2015 guideline highlights that “balancing treatment groups with respect to one or more specific prognostic covariates can enhance the credibility of the trial results” (EMA, 2015). The establishment of the American Statistical Association (ASA) Biopharmaceutical Section (BIOP) Covariate Adjustment Scientific Working Group (SWG) (https://carswg.github.io/) further underscores the growing focus on advancing statistical methods for covariate adjustment in clinical trials.
Despite these regulatory endorsements, significant challenges remain due to high-dimensional covariates and nonlinear relationships. These complexities necessitate robust and comprehensive analytical strategies. However, the comparative advantages and limitations of existing methods remain insufficiently explored, leaving practitioners with limited guidance and uncertainty in selecting the most appropriate approaches. For instance, in the case study presented in Section 5, the Best Apnea Interventions for Research (BestAIR) trial includes data from $N=169$ participants and $p=114$ covariates, making $p$ comparable to $N$. When adjusting for covariates using data from a single treatment arm—where $N_{1}=83$ (treated) and $N_{0}=86$ (controls)—$p$ exceeds the sample sizes. This high-dimensional setting can lead to biased and imprecise covariate-adjusted average treatment effect (ATE) estimates, as the sample size may be insufficient for reliable parameter estimation. Nevertheless, prior studies, such as Gao et al. (2024b), have identified baseline covariate imbalances in the BestAIR trial data, underscoring the need for efficient covariate-adjusted methods. This motivates our approach to improving estimation precision while addressing challenges aforementioned.
This paper presents a unified framework for covariate adjustment in RCT analysis, with a primary focus on selecting important covariates for reducing data dimensionality for efficiency gains. We provide a comprehensive tutorial on commonly used covariate-adjusted approaches, clarifying their underlying assumptions, strengths, and limitations. Additionally, we offer cautionary recommendations to guide practitioners in navigating potential pitfalls and selecting appropriate methods for specific scenarios. By addressing the aforementioned challenges, this framework aims to improve the reliability and interpretability of RCT findings, aligning with the evolving needs of researchers and regulatory agencies.
The rest of this section reviews related work on covariate adjustment and outlines our contributions. Section 2 covers key statistical preliminaries, including notation, assumptions, and the covariate adjustment methods employed in our framework. Section 3 introduces the Covariate Adjustment with Variable Selection (COADVISE) framework, highlighting its integration of variable selection and covariate-adjusted estimators. Numerical experiments assess finite-sample performance are presented in Section 4. Section 5 demonstrates the application of COADVISE through a re-analysis of BestAIR trial data. Finally, Section 6 concludes the paper, discussing implications, limitations, and future directions. Additional tutorial guidance and R code are provided in Appendix S.
1.2 Related work
Covariate adjustment in RCTs has a rich history in the statistical literature. One of the earliest and most widely used approaches is the analysis of covariance (ANCOVA), introduced by Fisher et al. (1966). This method estimates the ATE by fitting a linear regression model where the binary treatment indicator and covariates serve as predictors of the outcome. The ordinary least squares (OLS) estimator is then applied to the treatment coefficient to obtain the ATE estimate. ANCOVA is appealing due to its intuitive structure: by projecting the outcome onto the covariates, it accounts for variability explained by those covariates, thereby improving precision (Reluga et al., 2024). Notably, this remains true even when the actual relationship between covariates and outcomes is more complex than a linear model (Wang et al., 2019, 2023; Frank E Harrell, 2024; Harrell, 2017).
Despite its strengths, ANCOVA is not universally superior to the simple estimator. Its efficiency and credibility can diminish when the conditional outcome models for treatment and control groups differ significantly (Freedman, 2008). To address these limitations, Lin (2013) proposed including interactions between the treatment and covariates in the model. This method, known as analysis of heterogeneous covariance (ANHECOVA) (Cassel et al., 1976; Yang and Tsiatis, 2001; Ye et al., 2023b), allows the regression model to capture heterogeneity in the treatment effect across covariates. Although ANHECOVA has historical roots, Lin (2013) was instrumental in establishing its advantages through a design-based lens. This estimator is not only unbiased but also asymptotically more efficient than the simple estimator under a wide range of conditions, without relying on restrictive model assumptions (Lin, 2013). The adoption of ANHECOVA has expanded in recent years, particularly for its ability to enhance efficiency in RCTs employing covariate-adaptive randomization schemes (Bannick et al., 2023). The U.S. FDA has endorsed its use, especially in evaluating conditional treatment effects, further strengthening its role in modern controlled trials (FDA, 2023; Van Lancker et al., 2023). Recent work by Reluga et al. (2024) provided a unified perspective on the relationships between the simple estimator, ANCOVA, and ANHECOVA, providing conditions under which ANHECOVA uniformly dominates other methods in terms of efficiency and precision.
Beyond ANCOVA and ANHECOVA, a prominent method in the analysis of RCTs is augmented inverse probability weighting (AIPW) (Ye et al., 2023b, a; Robins et al., 1994; Tsiatis, 2006; Glynn and Quinn, 2010; Cao et al., 2009). The AIPW estimator combines outcome modeling with inverse probability weighting, effectively leveraging information from both observed outcomes and treatment assignment probabilities. This dual-model approach enables AIPW to achieve semiparametric efficiency, which means it attains the semiparametric variance lower bound when the outcome models are correctly specified. As a result, AIPW offers the highest asymptotic efficiency among regular ATE estimators in RCTs (Hahn, 1998; Hirano et al., 2003; Kang et al., 2007; Glynn and Quinn, 2010). One of the defining features of AIPW is its robustness to model misspecification in the context of RCTs. The randomization inherent in trial design allows AIPW to consistently estimate the ATE even when the outcome model is incorrectly specified. This robustness, combined with its efficiency, makes AIPW particularly attractive in practice. Moreover, AIPW offers flexibility by accommodating a wide range of parametric and nonparametric models for outcome estimation, making it more adaptable than ANCOVA or ANHECOVA, which depend on fixed linear model structures. Interestingly, when linear regression is used for outcome modeling within the AIPW framework, the AIPW estimator becomes equivalent to ANHECOVA. In this sense, ANHECOVA can be viewed as a special case of AIPW (Van Lancker et al., 2023). However, this equivalence also underscores a limitation of ANHECOVA: if the linear model is misspecified, the efficiency gains achieved by ANHECOVA are limited compared to those of AIPW when the latter employs correctly specified—or at least better specified—models. For this reason, AIPW is often regarded as a fully adjusted estimator in the literature, capable of optimizing efficiency without being overly reliant on strict parametric assumptions (Li et al., 2023).
As mentioned in Section 1.1, all covariate adjustment methods encounter difficulties in ensuring parameter convergence when faced with high-dimensional covariate data. As a result, a key preliminary step is variable selection—ideally choosing a set of variables smaller than the sample size $N$—that are most predictive of the outcome. Cho and Yang (2023) and Belloni et al. (2014) proposed strategies tailored to covariate adjustment in observational studies, and Yang et al. (2020) presented complementary methods in the context of data integration. Following these prior works, we aim to select covariates that strongly predict the outcome.
Several methods have been studied for data-driven variable selection procedures in RCTs. For example, the least absolute shrinkage and selection operator (Lasso) (Tibshirani, 1996) has been used in conjunction with ANHECOVA to select relevant covariates (Bloniarz et al., 2016; Schochet, 2022; Liu et al., 2023; Zhu et al., 2024). Another approach involves selecting the top $k$ covariates with the highest marginal correlations with the outcome, where $k<p$ (Senn, 1994). Alternatively, covariates can be screened using preliminary significance tests, such as two-sample $t$-tests or $z$-tests, to identify those most strongly associated with the treatment assignment or outcome (Permutt, 1990; Zhao and Ding, 2023). These methods have primarily focused on continuous outcomes and the application of ANCOVA and ANHECOVA estimators. As we noted, identifying which covariate sets are most beneficial for covariate adjustment is the objective. Cho and Yang, along with additional references, propose various strategies tailored to specific purposes of covariate adjustment in observational studies. Similarly, Yang et al. (2020) offered aligned strategies in the context of data integration. In the context of RCTs, where bias is controlled, covariate adjustment primarily aims to enhance efficiency. Therefore, consistent with previous research (Cho and Yang, 2023; Yang et al., 2020), we recommend selecting covariates that are strong prognostic factors for the outcome.
For binary or categorical outcomes, research on covariate adjustment has also made significant progress (Ye et al., 2023a; Moore and van der Laan, 2009; Jiang et al., 2017; Guo and Basse, 2023; Cohen and Fogarty, 2024), though variable selection has not been a primary focus in these studies. Notably, Van Lancker et al. (2024) introduced a unified framework for variable selection combined with the AIPW estimator. Their approach also leverages double/debiased machine learning (DML) algorithms (Chernozhukov et al., 2018), employing techniques such as sample splitting and cross-fitting for model fitting and outcome prediction during covariate adjustment. This integration of machine learning with semiparametric methods represents a promising direction for addressing high-dimensional covariates while maintaining the robustness and efficiency of AIPW-based estimators.
1.3 Our contributions
The contributions of this paper center on advancing covariate adjustment methods for RCTs through the development of a unified framework, COADVISE. Accompanying this framework is the Coadvise R package (https://github.com/yiliu1998/Coadvise), designed to make these methods accessible and user-friendly for researchers. Our work builds on existing methodologies by addressing several limitations and enhancing flexibility in the analysis of RCT data.
First, we extend traditional covariate adjustment methods, such as ANCOVA, ANHECOVA, and AIPW, by incorporating a variety of variable selection techniques, including Lasso, adaptive Lasso, marginal correlation, and preliminary testing into the analysis stage of RCTs. This integration addresses the challenges of high-dimensional covariates, enabling efficient and robust estimation of treatment effects. Unlike existing tools such as RobinCar, our framework explicitly supports variable selection, allowing users to identify covariates most predictive of outcomes, thereby improving estimation efficiency.
Second, since missing data are common in RCTs, we provide several missing data imputation methods in our Coadvise R package for user convenience. These methods allow users to impute missing values in the outcome and/or covariates before applying the COADVISE framework. Our package includes complete-case analysis, multiple imputation by chained equations (mice), random forest-based imputation (missForest), inverse probability weighting (ipw), and the use of missingness indicators (missInd). These methods ensure robustness across a range of assumptions about missing data mechanisms, such as missing completely at random (MCAR) or missing at random (MAR). Notably, while tools like RobinCar focus primarily on covariate adjustment, they lack native support for missing data imputation, distinguishing Coadvise as a more comprehensive solution.
To further highlight the differences between Coadvise and RobinCar, we provide a detailed comparison of their capabilities within the context of covariate adjustment for RCTs. Our framework not only supports multi-valued treatments, missing data imputation, and variable selection but also facilitates interoperability by enabling users to export selected covariates for subsequent analysis in RobinCar or custom code. This flexibility ensures that our framework complements rather than replaces existing tools, promoting seamless integration into diverse analytical workflows.
The following infotable (Table 1) provides an overview of our COADVISE framework and functions provided by our Coadvise R package, including available variable selection methods, missing data imputation methods, and covariate-adjusted estimators, along with references to the relevant literature and the properties of different methods. As a remark, based on empirical evidence in Section 4, we recommend the use of Lasso or adaptive Lasso as the variable selection method in general, and we set Lasso as the default choice in our R package.
Finally, we apply the COADVISE framework to a real-world dataset from the BestAIR trial, using data from the National Sleep Research Resource (NSRR). This case study illustrates the practical benefits of our framework and highlights its relevance for both academic research and regulatory applications. The open-source Coadvise package includes detailed documentation and reproducible examples, ensuring accessibility for a broad audience.
Missing data imputation Handles missing covariates Handles missing outcome Assumption
Complete-case analysis (cc) $\checkmark$ $\checkmark$ MCAR
Multiple imputation by chained equations (mice; Van Buuren and Groothuis-Oudshoorn (2011)) $\checkmark$ $\checkmark$ MAR
Random forest (missForest; Stekhoven and Bühlmann (2012)) $\checkmark$ $\checkmark$ MAR
Inverse probability weighting (ipw; Robins et al. (1994)) $\times$ $\checkmark$ MAR
Missingness indicator (missInd; Zhao et al. (2024); Zhao and Ding (2022); Song et al. (2021)) $\checkmark$ $\times$ MAR
Variable selection method Lasso (Tibshirani, 1996), Adaptive Lasso (Zou, 2006), Marginal correlation (Permutt, 1990), Preliminary testing (Senn, 1994; Zhao and Ding, 2023)
Covariate-adjusted estimator Model-based variance estimator (see Section 2) Robust to model misspecification (see Section 2) Nonlinear adjustment (see Section 3) Efficiency gain guarantee (see Section 3)
ANCOVA (Fisher et al., 1966) Sandwich $\checkmark$ $\times$ $\checkmark$ if $\pi_{1}=\pi_{0}^{\dagger}$ or under a constant treatment effect
ANHECOVA (Lin, 2013; Ye et al., 2023b) Sandwich $\checkmark$ $\times$ $\checkmark$
AIPW (Bannick et al., 2023) Delta method $\checkmark$ $\checkmark$ $\checkmark$ if (i) or (ii) in Theorem 1 holds
Software Multi-valued treatments Missing data imputation Variable selection Covariate-adjusted estimators
Coadvise (ours) $\checkmark$ $\checkmark$ $\checkmark$ ANCOVA, ANHECOVA, AIPW
RobinCar (Ye et al., 2023b) $\checkmark$ $\times$ $\times$ ANCOVA, ANHECOVA, AIPW
  • MCAR: missing completely at random; MAR: missing at random.
  • $\dagger$: $\pi_{a}=P(A=a)$ is the probability of being assigned to the treatment group $a$, where $a=0,1$.
Table 1: Overview of the COADVISE Framework: The top two sections include all available methods for imputing missing data and variable selection. The third section summarizes available covariate-adjusted estimators with associated variance estimators and their properties. The bottom section provides a comparison with the RobinCar package.
2 Preliminaries
2.1 Notation and set-up
The notation in this paper follows the “super-population” framework (Ding et al., 2017), which assumes that each observation is a random sample drawn from an infinite population under certain inclusion/exclusion criteria (Frank E Harrell, 2024). For instance, when estimating the ATE of an active drug for patients with a certain disease, it may be reasonable to assume that individuals are drawn from such a super-population, and the findings from the RCT can inform whether the drug is effective to the target population. In contrast, the “finite-population” framework (Bloniarz et al., 2016; Greenland et al., 1999; Schochet, 2010; Stuart et al., 2011) is applicable when focusing on the ATE for a policy applied to a fixed number of units, such as the states in the U.S., where the number of units is finite and fixed. A key difference between these two frameworks lies in the source of randomness: in the finite-population framework, randomness arises from treatment assignment and outcomes, whereas in the super-population framework, additional randomness comes from the covariates.
We consider a binary treatment, denoted by $A\in\{0,1\}$, where $1$ represents the treated group and 00 represents the control group. Each participant in the sample is characterized by a baseline covariate vector $\mathbf{X}$, such as clinical characteristics. We adopt the Neyman-Rubin potential outcomes framework (Neyman, 1923; Imbens and Rubin, 2015), in which each participant has two potential outcomes, $Y(0)$ and $Y(1)$, representing the potential outcomes under control and treatment, respectively.
2.2 Causal parameter, identification assumptions, and the unadjusted estimator
We are interested in estimating the ATE from RCT data, which is defined as:
$\tau=\mathbb{E}\{Y(1)-Y(0)\},$
where the expectation $\mathbb{E}\{\cdot\}$ is taken over the target population represented by the trial sample. However, we can only observe one outcome for each participant, denoted by $Y$. Given the nature and definition of potential outcomes, we assume that the observed outcome is consistent with the potential outcome under the received treatment, i.e., $Y=Y(A)=AY(1)+(1-A)Y(0)$. We further assume that there is only one version of the treatment $A$, and the potential outcome of each participant neither depends on nor affects the treatment received by others (also known as the stable-unit treatment value assumption [SUTVA]). In addition, we make the following standard identification assumptions for RCT data.
Assumption 1 (Randomization).
[…]
Assumption 2 (Positivity).
There exist two constants $c_{1}$ and $c_{2}$ such that $10<c_{1}<\pi_{1}=P(A=1)<c_{2}<$ 1 almost surely.
Assumption 1 is reasonable because, in RCTs, the treatment assignment is controlled by the experimenters. Assumption 2 is also commonly satisfied, as the probability of assigning each participant to either the treatment or control group is fixed and non-zero in RCTs, such as $\pi_{1}=0.5$ in the case of a 1:1 randomization.
Under these assumptions, the ATE $\tau$ from an RCT sample can be identified using the following formula:
$\tau=\frac{\mathbb{E}(AY)}{\mathbb{E}(A)}-\frac{\mathbb{E}\{(1-A)Y\}}{\mathbb{E}(1-A)}.$
Thus, given a set of independent and identically distributed (i.i.d.) data, denoted by $\mathcal{O}=\{(\mathbf{X}_{i},A_{i},Y_{i}),i=1,\dots,N\}$, we can estimate $\tau$ using the simple estimator:
$\widehat{\tau}_{\text{simple}}=\frac{1}{N_{1}}\sum_{i=1}^{N}A_{i}Y_{i}-\frac{1}{N_{0}}\sum_{i=1}^{N}(1-A_{i})Y_{i},$ (1)
where $N_{1}=\sum_{i=1}^{N}A_{i}$ and $N_{0}=\sum_{i=1}^{N}(1-A_{i})$. The variance estimator for $\widehat{\tau}_{\text{simple}}$, often referred to as the Neyman-type variance estimator, is given by:
$\widehat{\sigma}^{2}_{\text{simple}}=\frac{\widehat{S}_{0}^{2}}{N_{0}}+\frac{\widehat{S}_{1}^{2}}{N_{1}},$
where $\widehat{S}_{a}^{2}=(N_{a}-1)^{-1}\sum_{i=1}^{N_{a}}I(A_{i}=a)(Y_{i}-\widehat{\bar{Y}}(a))^{2}$, with $\widehat{\bar{Y}}(a)={N_{a}}^{-1}\sum_{i=1}^{N_{a}}I(A_{i}=a)Y_{i}$. The Neyman-type variance estimator is unbiased for the true variance of $\widehat{\tau}_{\text{simple}}$ under the super-population framework. Additionally, the estimator $\widehat{\tau}_{\text{simple}}$ when divided by the Neyman-type variance estimator, corresponds to the usual two-sample t-test statistic (or the F-statistic in one-way ANOVA). As a result, $\widehat{\tau}_{\text{simple}}$ is sometimes referred to as the ANOVA estimator in the literature (St et al., 1989; Ye et al., 2023b; Reluga et al., 2024).
The estimator $\widehat{\tau}_{\text{simple}}$ is consistent for $\tau$ and asymptotically normal, and it is attractive in practice due to its simplicity and ease of interpretation. However, incorporating covariate information can improve both the efficiency and credibility of ATE estimation, as discussed in Section 1.2. Therefore, we introduce several covariate-adjusted estimators in the following sections.
2.3 Linear covariate adjustment
Two historically popular methods for covariate adjustment are analysis of covariance (ANCOVA) and analysis of heterogeneous covariance (ANHECOVA) (Fisher et al., 1966; Cassel et al., 1976). Both methods adjust for covariates using linear regression models. The ANCOVA model (Fisher et al., 1966) is specified as:
$\mathbb{E}\{Y\mid\mathbf{X},A\}=\gamma_{0}+\beta_{1}A+\mathbf{X}^{\prime}\boldsymbol{\gamma}_{1},$ (2)
where the ATE estimator from ANCOVA is the ordinary least squares (OLS) estimator for the coefficient $\beta_{1}$, i.e., $\widehat{\tau}_{\text{ANCOVA}}=\widehat{\beta}_{1}^{\text{ols}}$.
A refinement of the ANCOVA model, known as the ANHECOVA model, includes interaction terms between the covariates and the treatment, and is specified as:
$\mathbb{E}\{Y\mid\mathbf{X},A\}=\gamma_{0}+\beta_{2}A+\mathbf{X}^{\prime}\boldsymbol{\gamma}_{1}+A(\mathbf{X}-\bar{\mathbf{X}})^{\prime}\boldsymbol{\gamma}_{2},$ (3)
where the ATE estimator from ANHECOVA is the OLS estimator for coefficient $\beta_{2}$, i.e., $\widehat{\tau}_{\text{ANHECOVA}}=\widehat{\beta}_{2}^{\text{ols}}$.
A robust variance estimator for both the ANCOVA and ANHECOVA estimators is the Huber-White (HW) sandwich variance estimator (Lin, 2013; Zhao and Ding, 2021). For ANCOVA, the HW variance estimator is written as
$\widehat{\mathbb{V}}(\widehat{\tau}_{\text{ANCOVA}})=(\mathbf{V}^{\prime}\mathbf{V})^{-1}\{\mathbf{V}^{\prime}\text{diag}(\widehat{\varepsilon}_{1}^{2},\dots,\widehat{\varepsilon}_{N}^{2})\mathbf{V}\}(\mathbf{V}^{\prime}\mathbf{V})^{-1},$ (4)
where $\mathbf{V}=(A,\mathbf{X})$, and $(\widehat{\varepsilon}_{1},\dots,\widehat{\varepsilon}_{N})$ is the vector of estimated regression residuals. For the ANHECOVA estimator, $\mathbf{V}$ is modified to include interaction terms as $\mathbf{V}=(A,\mathbf{X},A(\mathbf{X}-\bar{\mathbf{X}}))$. In the finite-population framework, equation (4) provides robust and consistent variance estimates for both ANCOVA and ANHECOVA estimators. However, under the super-population framework, additional uncertainty arises from estimating $\bar{\mathbf{X}}$ as an approximation of $\mathbb{E}(\mathbf{X})$. Therefore, the corrected HW variance estimator for the ANHECOVA estimator is given by (Zhao and Ding, 2021):
$\widehat{\mathbb{V}}(\widehat{\tau}_{\text{ANHECOVA}})=(\mathbf{V}^{\prime}\mathbf{V})^{-1}\{\mathbf{V}^{\prime}\text{diag}(\widehat{\varepsilon}_{1}^{2},\dots,\widehat{\varepsilon}_{N}^{2})\mathbf{V}\}(\mathbf{V}^{\prime}\mathbf{V})^{-1}+(\widehat{\alpha}_{1}-\widehat{\alpha}_{0})^{\prime}{S_{X}^{2}}(\widehat{\alpha}_{1}-\widehat{\alpha}_{0})/N,$
where $S_{X}^{2}$ is the sample covariance of $\mathbf{X}$, and $\widehat{\alpha}_{a}$ is the OLS estimator from the regression model $\mathbb{E}(Y(a)\mid\mathbf{X})=\xi_{a}+\alpha_{a}\mathbf{X}$, using data from the group with $A=a$, for $a=0,1$.
It has been shown that the ANHECOVA estimator, $\widehat{\tau}_{\text{ANHECOVA}}$, is consistent, asymptotically normal, and asymptotically more efficient than both $\widehat{\tau}_{\text{simple}}$ and $\widehat{\tau}_{\text{ANCOVA}}$ (Lin, 2013; Ye et al., 2023b). However, both $\widehat{\tau}_{\text{ANCOVA}}$ and $\widehat{\tau}_{\text{ANHECOVA}}$ have a limitation: they cannot accommodate non-linear adjustments. In some cases, non-linear adjustments may lead to greater efficiency gains, particularly when the true outcome model follows a known non-linear form given some prior knowledge. The augmented inverse probability weighted (AIPW) estimator, introduced in the next section, offers a more flexible approach to covariate adjustment by accommodating both linear and non-linear outcome regression models in the two treatment groups.
2.4 Non-linear covariate adjustment
The AIPW estimator for RCT data is given by (Bannick et al., 2023; Glynn and Quinn, 2010):
$\widehat{\tau}_{\text{AIPW}}=\widehat{\theta}_{1}-\widehat{\theta}_{0},$ (5)
where for $a=0,1$,
$\widehat{\theta}_{a}=\frac{1}{N_{a}}\sum_{i=1}^{N}I(A_{i}=a)[Y_{i}-\widehat{\mu}_{a}(\mathbf{X}_{i})]+\frac{1}{N}\sum_{i=1}^{N}\widehat{\mu}_{a}(\mathbf{X}_{i}),\text{ with }N_{a}=\sum_{i=1}^{N}I(A_{i}=a).$
In (5), $N_{a}$ is the number of subjects in group $A=a$, and $\widehat{\mu}_{a}(\mathbf{X})$ is the predicted value of $\mathbb{E}\{Y(a)\mid\mathbf{X}\}$, obtained by fitting a working model $\mu_{a}(\cdot)$ under treatment $A=a$ (e.g., a generalized linear model (GLM) $\mu_{a}(\mathbf{X})=g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})$). This model may not be correctly specified, where correct specification means $\mu_{a}(\mathbf{X})\to_{p}\mathbb{E}\{Y(a)\mid\mathbf{X}\}$, where $\to_{p}$ denotes convergence in probability. It can be shown that when $\mu_{a}(\mathbf{X})$ is a linear model for $Y(a)$, the AIPW estimator coincides with the ANHECOVA estimator, provided that the same set of covariates are used as predictors (Ye et al., 2023b; Van Lancker et al., 2023; Reluga et al., 2024). Thus, ANHECOVA can be viewed as a special case of the AIPW estimator.
Remark 1.
In practice, the working model for $\mu_{a}(\mathbf{X})$ in the AIPW estimator can be fit using black-box machine learning (ML) models and combined through ensemble methods such as SuperLearner (Chernozhukov et al., 2018; Van der Laan et al., 2007). However, our focus on parametric models is driven by the theoretical results on efficiency gain presented in Theorem 1 (Section 3.2). Specifically, we establish efficiency conditions based on the estimation approach for $\boldsymbol{\beta}_{a}$. While ML-based estimation methods are flexible, deriving unified theoretical results on efficiency gain in this setting remains an open question, which we leave for future research. Further discussion on our decision not to employ ML models in this work is provided in Remark 2, where we formally introduce the COADVISE framework and outline additional considerations.
A robust variance estimator for the AIPW estimator can be derived using the Delta method (Taylor’s expansion) (Ye et al., 2023a). Under certain regularity conditions (Bannick et al., 2023) and denoting $\boldsymbol{\theta}=(\theta_{0},\theta_{1})^{\prime}$, it can be shown that:
$\sqrt{N}(\widehat{\boldsymbol{\theta}}-\boldsymbol{\theta})\to_{d}\mathcal{N}(\boldsymbol{0},\mathbf{V}),$ (6)
where $\to_{d}$ denotes convergence in distribution, and
$\mathbf{V}=\begin{pmatrix}v_{11}&v_{12}\\ v_{21}&v_{22}\end{pmatrix},$ (7)
with
$v_{aa}$ $=P(A=a)^{-1}\mathbb{V}\{Y_{a}-\mu_{a}(\mathbf{X})\}+2\text{Cov}\{Y_{a},\mu_{a}(\mathbf{X})\}-\mathbb{V}\{\mu_{a}(\mathbf{X})\},$
$v_{ab}$ =$\text{Cov}\{Y_{a},\mu_{b}(\mathbf{X})\}+\text{Cov}\{Y_{b},\mu_{a}(\mathbf{X})\}-\text{Cov}\{\mu_{a}(\mathbf{X}),\mu_{b}(\mathbf{X})\}\text{ for }a\not=b.$
The function $\mu_{a}$ is the assumed limit of $\widehat{\mu}_{a}$ used in (5), and it satisfies $\|\widehat{\mu}_{a}-\mu_{a}\|_{2}\to_{p}0$, where $\|\cdot\|_{2}$ represents a finite $L_{2}$ norm. However, $\mu_{a}$ is not necessarily equal to the true outcome model $\mathbb{E}\{Y(a)\mid\mathbf{X}\}$, unless the working model is correctly specified.
The asymptotic variance for $\widehat{\tau}_{\text{AIPW}}=\widehat{\theta}_{1}-\widehat{\theta}_{0}=c^{\prime}\widehat{\boldsymbol{\theta}}:=f(\widehat{\boldsymbol{\theta}})$ is given by $\{\nabla f(\boldsymbol{\theta})\}^{\prime}\mathbf{V}\{\nabla f(\boldsymbol{\theta})\}=v_{11}-2v_{12}+v_{22}$, where $c=(1\ \ -1)^{\prime}$. Therefore, a consistent variance estimator is:
$N^{-1}\{\nabla f(\widehat{\boldsymbol{\theta}})\}^{\prime}\widehat{\mathbf{V}}\{\nabla f(\widehat{\boldsymbol{\theta}})\}=N^{-1}\{\widehat{v}_{11}-2\widehat{v}_{12}+\widehat{v}_{22}\},$
where $\widehat{v}_{11}$, $\widehat{v}_{12}$, and $\widehat{v}_{22}$ are sample-based estimates of $v_{11}$, $v_{12}$, and $v_{22}$, respectively, using the sample versions of the variances, covariances, and proportions.
3 Methodology
All of the covariate-adjusted estimators, as discussed in Section 2, can suffer from high-dimensional issues. When the number of available covariates is large relative to the sample size, or when overfitting occurs (Bloniarz et al., 2016), the finite-sample performance of these estimators may deteriorate, potentially leading to efficiency losses compared to simpler estimators. In addition, the consistency and robustness of variance estimators can be compromised. For example, computing matrix inverses in sandwich variance estimators may become unstable or infeasible, resulting in unavailable or unreliable standard error estimates (Matsouaka et al., 2023, 2024).
To address these concerns, we incorporate a variable selection step before performing covariate-adjusted ATE estimation. The goal is to retain only the most important variables—identified through data-driven criteria—while ensuring that the number of selected covariates remains small relative to the sample size (i.e., of lower order than $N$). This approach balances the inclusion of relevant covariates with the exclusion of less informative ones, improving finite-sample efficiency. It yields an estimator that lies between the simple estimator (using no covariates) and the fully adjusted estimator (using all covariates), achieving higher efficiency while maintaining consistency.
We also note that our COADVISE framework generates predicted mean potential outcome vectors for all participants in multi-valued treatment settings, allowing users to compute any pairwise contrasts between mean outcomes. However, for consistency with the notation in Section 2, the following Sections 3.1 and 3.2 focus on the AIPW estimator under binary treatment to illustrate our workflow. This focus is motivated by the AIPW estimator’s superior theoretical properties, intuitive interpretation, and straightforward generalization to multi-valued treatments. Additional practical guidelines for users are provided in Appendix S.
For simplicity, we assume here that the observed data contain no missing values, which allows us to concentrate on high-dimensional issues and variable selection. In practice, if missing data are present, we first impute them within the COADVISE framework to obtain a complete dataset. By separating the imputation and variable selection steps, the subsequent theoretical results remain valid without loss of generality.
3.1 The COADVISE workflow
Figure 1: An illustration of the COADVISE framework using AIPW estimation for the covariate adjustment estimator consisting of two stages: Stage I: Perform variable selection for the conditional outcome models in both treatment groups. Stage II: Re-fit the conditional outcome models using the selected variables and compute the AIPW estimator. The model re-fitting step aligns with the theoretical results on efficiency gains, ensuring that the estimator benefits from covariate adjustment with variable selection.
Figure 1 illustrates our proposed workflow using the AIPW estimator as an example. The process begins with variable selection performed separately for each treatment group. This approach allows the outcome models $\mu_{0}(\mathbf{X})$ and $\mu_{1}(\mathbf{X})$ to include different sets of variables, thereby providing greater flexibility in capturing the most relevant features for each potential outcome based on the data.
Next, we fit two parametric outcome regression models—one for $Y(1)$ and one for $Y(0)$—using a chosen link function $g(\cdot)$ from the GLM family (for example, logistic regression for a binary outcome).
Once these models are estimated, each model is applied to the data from the opposite treatment group to predict the missing potential outcomes. In other words, a model $\widehat{\mu}_{a}(\boldsymbol{X})$ fitted using data from group $A=a$ is used to predict the potential outcome $Y(a)$ for subjects in group $A=1-a$, for $a=0,1$. This procedure yields all the information necessary to compute the AIPW estimator in (5). Finally, we quantify uncertainty by employing the robust variance estimator in (6) (Ye et al., 2023a).
For a more technical illustration, Appendix S.2 provides an algorithm box detailing the Lasso variable selection combined with AIPW estimation.
Remark 2.
Although variable selection methods like Lasso and adaptive Lasso can directly produce a fitted model by selecting variables with nonzero regression coefficients, we take an additional step by refitting a GLM using only the selected variables and estimating the coefficients via maximum likelihood estimation (MLE). This refitting step is essential for two reasons. First, it facilitates achieving the efficiency gains established in Theorem 1 in the following section. Second, it enhances the interpretability of the covariate-adjusted estimator by relying on transparent parametric models rather than black-box approaches.
3.2 Theoretical results
We establish theoretical results for efficiency gains via the AIPW estimator, with detailed proofs provided in Web Appendix A of our Online Supplemental Material. The key results are summarized in the following Theorem 1.
Theorem 1.
Suppose Assumptions 12 in Section 2.2 hold, and denote the asymptotic variance of the AIPW and simple estimator as $\mathbb{V}_{\text{aipw}}$ and $\mathbb{V}_{\text{simple}}$, respectively. Then, regardless of the variable selection method used, $\mathbb{V}_{\text{aipw}}\leq\mathbb{V}_{\text{simple}}$ in any of the following cases:
  • (i) Both conditional outcome models $\mu_{0}(\mathbf{X})$ and $\mu_{1}(\mathbf{X})$, after variable selection, are specified by linear models with OLS used for estimating the regression coefficients.
  • (ii) If the outcome model for $Y(a)$ ($a=0,1$), after variable selection, is specified by a nonlinear GLM with link function $g$, then there exists a $\boldsymbol{\beta}_{a}$ as the probability limit of MLE $\widehat{\boldsymbol{\beta}}_{a}$ such that with probability 1: (a) the sign of $g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})$ matches that of $\mathbb{E}(Y(a)\mid\mathbf{X})$; and (b) $|g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})|\leq 2|\mathbb{E}(Y(a)\mid\mathbf{X})|$.
For point (i) in Theorem 1, we highlight the connection with the ANHECOVA estimator. If both $\mu_{0}(\mathbf{X})$ and $\mu_{1}(\mathbf{X})$ use the same set of covariates as ANHECOVA, then the AIPW estimator is identical to ANHECOVA, so in this case, the efficiency gain result is consistent with that established by Lin (2013) However, due to variable selection, $\mu_{0}(\mathbf{X})$ and $\mu_{1}(\mathbf{X})$ may use different covariates, making the AIPW estimator distinct but still guaranteeing efficiency gains when linear regressions are applied.
For point (ii), conditions (a) and (b) have the following interpretations: for (a), the GLM must correctly capture the sign of the true conditional outcome with probability 1; and for (b), the bias of the GLM is constrained to maintain efficiency. For example, with positive-valued outcomes, conditions (a) and (b) can be simplified to $0\leq g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})\leq 2\mathbb{E}(Y(a)\mid\mathbf{X})$, which is reasonable for some positive categorical outcomes, such as binary outcomes modeled by logistic regression. Further illustration and discussion of these conditions are provided in Web Appendix A of our Online Supplemental Material.
Comparing (i) and (ii), linear models with OLS ensure efficiency gains without additional bias constraints. However, nonlinear GLMs may offer higher efficiency gains in specific cases (e.g., binary outcomes), since the AIPW estimator’s efficiency gain depends on the bias between the posited and true outcome models. When both outcome models are correctly specified, the AIPW estimator is the most efficient, achieving the asymptotic (semiparametric) variance lower bound (Hahn, 1998; Hirano et al., 2003).
4 Numerical Experiments
In this section, we present a series of Monte Carlo simulation experiments to evaluate different configurations within our COADVISE framework. We use a comprehensive data-generating process (DGP), detailed in Section 4.1, and generate two distinct types of datasets for each DGP.
First, we generate 20 independent super-population full datasets, each containing $5\times 10^{7}$ independent units, with both potential outcomes $(Y(0),Y(1))$ available for each unit. We then compute the mean of $Y(1)-Y(0)$ across these 20 large datasets to obtain the true ATE. Given the large sample size and 20 replications, the associated uncertainty is negligible.
Second, we generate $M=500$ independent random sample datasets of size $N$, with $N$ varying between 100 and 500 to represent different sample size scenarios. In each Monte Carlo replication, the dataset consists of $N_{1}$ treated units and $N_{0}$ control units, with an expected 1:1 ratio such that $N_{0}+N_{1}=N$.
4.1 Data generating process
In line with the nature of an RCT, we generate the binary treatment assignment as $A\sim\text{Bern}(0.5)$, independent of any covariates or outcomes, without imposing any restrictions on the sample size. We consider two sets of covariates: $\mathbf{X}=(X_{1},\dots,X_{5})^{\prime}$ and $\mathbf{V}=(V_{1},\dots,V_{50})^{\prime}$. The set $\mathbf{X}$ is associated with the outcomes, while $\mathbf{V}$ consists of noise variables that are uncorrelated with both the outcomes and the treatment.
We then specify two models for the potential outcomes:
Continuous outcome: $Y(a)=50+20\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}+a\delta(\mathbf{X})+\epsilon,$
Binary outcome: $Y(a)\sim\text{Bern}(e(\mathbf{X},a)),\text{ where }e(\mathbf{X},a)=\{1+\exp(-20\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}-a[3+\delta(\mathbf{X})])\}^{-1},$
for $a=0,1$, where $\epsilon\sim\mathcal{N}(0,1)$ i.i.d., and $\delta(\mathbf{X})$ characterizes the individual treatment effect. The details of $\boldsymbol{\beta}_{0}$, different cases of $\delta(\mathbf{X})$ (linear and nonlinear), and the distributions used for $\mathbf{X}$ and $\mathbf{V}$ can be found in our Online Supplemental Material for reproducibility.
4.2 Methods for comparison
We consider the following competing methods for estimating ATE: Simple, ANCOVA, ANHECOVA, and AIPW estimators. For each of these methods, we apply the variable selection procedures described in Table 2 to select covariates from $\{\mathbf{X},\mathbf{V}\}$ defined in Section 4.1. All estimation and variable selection methods are available through our R package, Coadvise.
Table 2: Competing methods for variable selection in simulation
Method abbreviation Description
All Include all $\mathbf{X}$ and $\mathbf{V}$, i.e., no covariate selection
Lasso Select covariates with non-zero regression coefficient estimates in Lasso regression (Bloniarz et al., 2016)
Adaptive Lasso Select covariates with non-zero regression coefficient estimates in adaptive Lasso regression (Zou, 2006)
Corr ($k$) Select the $k$ covariates with the highest marginal correlations with $Y$, where $k=1,3,10$ (Permutt, 1990)
Corr ($\xi$) Select covariates whose marginal correlations with $Y$ are greater than $\xi$, where $\xi=0.10,0.25$ (Permutt, 1990)
Pre-test ($\alpha$) Perform a univariate two-sample $t$-test on each covariate to test for differences between two groups at level $\alpha$. Select covariates where the test is rejected, with $\alpha=0.05,0.10$ (Senn, 1994; Zhao and Ding, 2023)
4.3 Performance criteria
The simulation results are presented using the following criteria: bias (via boxplots), coverage probability (CP%), and empirical percent power (Power%). Denote $\tau$ as the true ATE value and $\widehat{\tau}$ as the estimated value from each Monte Carlo replication. Across $M=500$ replications, these measurements are defined as follows:
  • Bias: $\widehat{\tau}-\tau$, represented in boxplots. From the width of a box, we can evaluate the estimation efficiency.
  • CP%: The proportion of replications where $\tau$ falls within the 95% confidence interval constructed from $\widehat{\tau}$ and its estimated standard error, using a normal approximation.
  • Power%: The proportion of replications where the p-value is $\leq 0.05$ for testing $H_{0}:\tau=0$ versus $H_{a}:\tau\neq 0$.
With $M=500$ simulations, a CPis considered significantly different from the nominal 95% coverage level if it falls outside the range $95\%\pm 1.96\times\sqrt{95\%\times 5\%/500}=[93\%,97\%]$.
4.4 Results
For brevity, we present the simulation results for the linear continuous outcome under both sample sizes of 100 and 500 in this section. Complete simulation results, including figures, are available in Appendix B of our Online Supplemental Material.
Figure 2: Simulation results for the linear continuous outcome with sample sizes (a) $N=100$ and (b) $N=500$. In the Bias boxplots, the black dots represent the medians. In the CPplots, the red dashed line indicates the 95% coverage level, while the gray dashed lines indicate 93% and 97% coverage levels. In the Powerplots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power. Missing values in the CPand Powerfor ANHECOVA and ANCOVA are due to the non-obtainable sandwich variance estimators in high-dimensional settings.
Figure 2 presents the results of biases via boxplots in the upper panels, and the results of CPand Powerin the lower four panels. We used different colors to distinguish between variable selection methods: green represents the estimator without variable selection (fully adjusted including all covariates), cyan represents Lasso and adaptive Lasso, gray indicates selection based on marginal correlation (as described in Table 2 for pre-specified $k$ or $\xi$), orange represents preliminary testing, and purple represents the simple estimator (which does not include any covariates).
From Figure 2, we observe that when the sample size is small ($N=100$), there are many outliers in the boxplots for the ‘No selection” case for both AIPW and ANHECOVA estimators. This is expected, since the fully adjusted AIPW and ANHECOVA estimators include a relatively large number of covariates compared to the sample size. Across both sample sizes, the Simple, Pre-test (both $\alpha=0.05$ and $0.1$), and Corr ($k=1,3$) estimators exhibit wider bias box widths, with the Simple estimator widest, indicating that these methods are less efficient in this simulation setting.
In terms of CPresults, we find that all methods are more stable around the 95% nominal level at the larger sample size, as the confidence intervals are Wald-type and rely on the large sample normal approximation. However, even when $N=500$, we observe below nominal CPfor both ANCOVA and ANHECOVA (due to the inability to obtain HW sandwich variance estimators in high-dimensional settings). Additionally, across both sample sizes, the Simple and Pre-test (both $\alpha=0.05$ and $0.1$) estimators result in the lowest power. This finding for the Pre-test estimator is consistent with Zhao and Ding (2023), where they advise against its use in practice based on theory and simulations. In contrast, Lasso and adaptive Lasso selection methods combined with AIPW generally perform well across both sample sizes and all three estimators in terms of efficiency, variance estimation (as measured by CP%), and power.
Although we do not include the results of other DGPs, we briefly summarize them here. The results for the other continuous outcome are similar to those shown in Figure 2. For the two binary outcomes, we find that the efficiency gains observed in the boxplots are also similar to the continuous outcome cases, although the improvement relative to the Simple estimator is less pronounced due to the binary outcome being restricted to values in $\{0,1\}$. Moreover, the issue of non-obtainable sandwich variance estimators for ANHECOVA and ANCOVA is also present in the binary outcome cases.
5 Case Study
We apply the COADVISE framework to analyze the Best Apnea Interventions for Research (BestAIR) trial data, available through the National Sleep Research Resources (NSRR) (https://sleepdata.org/datasets/bestair/files/datasets). The dataset includes 169 participants, with 83 assigned to the active treatment group (continuous positive airway pressure [CPAP]) and 86 to a combined control arm. The primary objective of the study is to assess the impact of CPAP on 24-hour systolic blood pressure (SBP). The data includes 114 covariates, making it high-dimensional given the sample size ($N=169$).
Both the outcome and covariates contain missing values. Since only 25 observations are complete, we use the random forest method (Stekhoven and Bühlmann, 2012) (missForest option in our package) to impute all missing values. We consider two outcomes: (i) continuous SBP values (in mmHg); and (ii) binary SBP values, classifying SBP as high ($\geq 130$ mmHg) or low ($<130$ mmHg) based on a 130 mmHg threshold.
Table 3 presents the results of the analysis using three approaches to variable selection: (i) no covariate selection (all covariates included), (ii) Lasso, and (iii) adaptive Lasso. When no variable selection is applied, only the simple estimator provides complete results for both outcomes. The standard errors (SEs) for the other methods are unavailable (NA), except for the AIPW estimator in the binary SBP case. This is due to the instability of the HW sandwich variance estimator for ANCOVA and ANHECOVA, which, given the sample size relative to the number of covariates, becomes unstable when inverting high-dimensional matrices, causing it to fail to converge. Additionally, in the case of continuous SBP, the variance estimated for AIPW using the Delta method was negative, leading to an unavailable (NA) SE. Therefore, without variable selection, the results of all covariate-adjusted estimators are not reliable.
However, when variable selection is applied, all covariate-adjusted methods return complete results, including SEs, CIs, and p-values. Interestingly, the results from the covariate-adjusted methods often differ notably from those of the simple estimator. For the continuous SBP outcome, the simple estimator indicates a significantly non-zero ATE with p-values less than 0.05, whereas none of the variable selection methods identify a significant ATE at the 0.05 level. Additionally, with variable selection, the SEs from the covariate-adjusted methods are smaller than those from the simple estimator for both outcomes. This suggests that proper covariate adjustment can enhance efficiency.
Table 3: Estimated average treatment effects by CPAP on blood SBP by data from the BestAIR trial
Variable selection Estimator Continuous SBP Binary SBP
Estimate (CI) SE p-value Estimate (CI) SE p-value
Simple -3.82 (-7.03, -0.62) 1.64 0.02 -0.12 (-0.25, 0.00) 0.06 0.06
No selection ANCOVA -0.19 (NA, NA) NA NA 0.08 (NA, NA) NA NA
ANHECOVA 48.68 (NA, NA) NA NA 0.48 (NA, NA) NA NA
AIPW 29.84 (NA, NA) NA NA -0.05 (-0.12, 0.02) 0.04 0.18
Simple -3.82 (-7.03, -0.62) 1.64 0.02 -0.12 (-0.25, 0.00) 0.06 0.06
Lasso ANCOVA -0.01 (-0.32, 0.30) 0.16 0.94 -0.00 (-0.08, 0.07) 0.04 0.93
ANHECOVA -0.08 (-0.42, 0.27) 0.18 0.67 0.00 (-0.08, 0.09) 0.04 0.92
AIPW -0.39 (-0.90, 0.12) 0.26 0.13 -0.04 (-0.13, 0.04) 0.04 0.30
Simple -3.82 (-7.03, -0.62) 1.64 0.02 -0.12 (-0.25, 0.00) 0.06 0.06
Adaptive Lasso ANCOVA -0.09 (-0.44, 0.27) 0.18 0.63 -0.03 (-0.11, 0.05) 0.04 0.45
ANHECOVA -0.29 (-1.05, 0.48) 0.39 0.46 -0.03 (-0.11, 0.05) 0.04 0.41
AIPW -0.32 (-0.79, 0.15) 0.24 0.19 -0.09 (-0.16, -0.01) 0.04 0.03
  • CI: confidence interval; SE: standard error; NA: not available.
6 Concluding Remarks
6.1 Summary
COADVISE is a comprehensive and flexible framework for covariate adjustment with variable selection, designed to enhance the reliability and efficiency of treatment effect estimation in RCTs. The framework accommodates several variable selection procedures, including no selection, Lasso, adaptive Lasso, marginal correlation methods, and preliminary testing, prior to covariate adjustment. We also implement three widely used covariate-adjustment methods—ANCOVA, ANHECOVA, and AIPW—in the estimation stage. For AIPW, we provide a broad range of parametric models for the conditional outcome models used in the estimator. Notably, we do not implement machine learning (ML) models in the final model fitting step after variable selection due to their black-box nature, which complicates theoretical guarantees for efficiency gains and interpretability. Instead, we focus on transparent parametric models to ensure clarity and alignment with the theoretical conditions established in our work.
Our framework stands out from existing methods for several reasons. To our knowledge, this tutorial is the first to offer comprehensive guidance on variable selection preceding covariate adjustment, while systematically comparing various variable selection and covariate-adjusted methods from a practical statistical perspective. Moreover, the data-driven variable selection techniques identify key predictors in a balanced manner, optimizing both efficiency and robustness and ensuring that the number of selected variables remains manageable given the sample size. Although some relevant variables may be omitted, this approach strikes a careful balance between the extremes of unadjusted and fully adjusted methods, thereby enhancing efficiency without compromising consistency.
Furthermore, our empirical studies suggest that Lasso and adaptive Lasso should be considered as priority in practice, since they generally outperform other methods for variable selection in our framework. Their key advantages include (i) flexibility in model specification aligning with the AIPW estimator’s outcome model(s), and (ii) the absence of some tuning parameters that must be selected by users, such as thresholds $k$ and $\xi$ in marginal correlation methods, or preliminary testing level $\alpha$, making them more practical for application. Although Lasso and adaptive Lasso do require a tuning parameter $\lambda$ in the penalty term (see Algorithm 1 in Appendix S.2), our package implements lambda.min, which selects the $\lambda$ value that minimizes the cross-validated error, making this process straightforward for users. Adaptive Lasso further has adaptive weights in the penalty term, which are inversely proportional to an initial estimate of the coefficients, obtained from an unpenalized OLS regression in our package.
6.2 Discussion
We acknowledge several limitations of the COADVISE framework. One notable concern is the issue of post-selection inference resulting from variable selection prior to ATE estimation (Fithian et al., 2014). Although our simulations suggest that our approach yields valid inferences, further theoretical investigation is warranted. In support of our method, we note that our model-based variance estimators (such as the sandwich or Delta methods) are robust to model misspecification, ensuring valid uncertainty quantification regardless of the variables selected for the outcome model(s). Furthermore, fitting the post-selection models using MLE or OLS improves efficiency, as demonstrated by the theory established in Section 3.2.
A related concern was raised in a recent commentary (Frank E Harrell, 2024), which argued that complex methods like ANHECOVA and AIPW, unlike ANCOVA, rely on the assumption that RCT participants are randomly sampled from the clinical population. This assumption ignores inclusion/exclusion criteria and may limit external validity. We respectfully contend that this issue is not unique to these complex estimators; even simpler estimators, including ANCOVA, target the ATE for the population defined by the trial’s criteria, since the RCT sample represents only those who met the criteria. Our primary contribution is to address the high-dimensional covariate challenges that affect internal validity, thereby ensuring robust ATE estimates for the trial population across varying sample sizes. Future extensions might explore data integration and external information borrowing to further enhance the generalizability of RCT findings (Lee et al., 2022, 2023, 2024a, 2024b).
Additionally, while our Coadvise R package includes a few commonly used missing data imputation methods for user convenience, we have not conducted a comprehensive investigation into their optimal integration within our framework, particularly in conjunction with variable selection. Future work could explore more robust approaches to handling missing data, including tailored imputation strategies that align with the underlying variable selection mechanisms to enhance both interpretability and statistical efficiency.
There are also several promising avenues for extending our framework. For instance, alternative adjustment methods could be incorporated to further improve efficiency. Existing literature suggests that integrating propensity score modeling, $e(\mathbf{X})=P(A=1\mid\mathbf{X})$, into the AIPW estimator (Gao et al., 2024b), or applying propensity score weighting methods without outcome modeling (such as inverse probability weights (Shen et al., 2014), overlap weights (Li et al., 2018a; Yang et al., 2021; Zeng et al., 2021), or other schemes (Li and Greene, 2013; Liu et al., 2024b)) can enhance efficiency, even when the propensity score is known (e.g., $e(\mathbf{X})=0.5$ for 1:1 randomization across all $\mathbf{X}$) (Williamson et al., 2014; Shen et al., 2014). However, extending our framework to include propensity score modeling introduces an additional variable selection challenge, effectively leading to a double selection problem for the AIPW estimator (Koch et al., 2018; Belloni et al., 2014). Moreover, although logistic regression is commonly used for propensity score estimation (Thomas et al., 2020; Hirano and Imbens, 2001; Connors et al., 1996) it remains uncertain whether it consistently enhances efficiency or if alternative methods with clearer regularity conditions might be preferable. Another potential extension involves prognostic covariate adjustment (PROCOVA), which employs machine learning techniques to generate prognostic scores based on baseline characteristics (Harrell Jr et al., 1984, 1996; Vanderbeek et al., 2022; Rubin and Thomas, 2000; Schuler et al., 2021; Li et al., 2024). Future work could extend our variable selection methods to PROCOVA or explore the joint use of propensity and prognostic scores for covariate adjustment (Yang and Zhang, 2023; Zhang et al., 2022).
Furthermore, our framework can be extended to a wide range of more general and complex settings. These include covariate-adaptive randomization (Bannick et al., 2023; Kernan et al., 1999; Broglio, 2018; Zhu et al., 2024; Yu et al., 2024), rerandomization (Zhu and Liu, 2023; Morgan and Rubin, 2012), time-to-event outcomes (Hua et al., 2024; Yi et al., 2020; Benkeser et al., 2021), restricted mean survival times (Han, 2023; Zhong et al., 2022; Karrison and Kocherginsky, 2018; Hanada et al., 2024), cluster-randomized trials (Murray, 1998; Donner et al., 2000; Stephens et al., 2012; Zhou et al., 2022; Balzer et al., 2023), and ordinal or composite outcomes with novel estimands such as the win ratio or generalized odds ratio (Agresti, 1980; Pocock et al., 2012; Barnhart et al., 2024; Matsouaka, 2022; Oakes, 2016; Zhang and Wu, 2024), as well as the Mann-Whitney rank-based statistic (Follmann et al., 2020). Other promising directions include conformal inference on individual or heterogeneous treatment effects (Yang et al., 2024; Wang et al., 2024; Gao et al., 2024a), the integration of multi-source RCT and observational data (Zhuang et al., 2024; Lee et al., 2024b, 2023; Yang et al., 2023; Han et al., 2025; Liu et al., 2024a; Yang and Wang, 2022; Liu et al., 2025), synthetic controls (Abadie et al., 2015, 2010; Ben-Michael et al., 2021), and external controls (Li et al., 2023; Lee et al., 2024c; Gao et al., 2023; Wang et al., 2025). Additionally, alternative strategies for estimating regression coefficients, such as debiased machine learning algorithms (Chernozhukov et al., 2018; Banik and Shao, 2022; Carlsson et al., 2014), debiased regressions (Lu et al., 2023; Chang et al., 2024; Chiang et al., 2023), and linear calibration (Bannick et al., 2023; Guo and Basse, 2023; Cohen and Fogarty, 2024) are of interest for future investigation.
Data and software availability
The BestAIR trial data used in Section 5 is available upon request at https://sleepdata.org. The R package Coadvise implementing the methodology of this paper is available at https://github.com/yiliu1998/Coadvise.
Acknowledgments
The authors are grateful to Marlena Bannick for insightful discussions during the early stages of this research endeavor. Liu was supported by the National Heart, Lung, and Blood Institute (NHLBI) of the National Institutes of Health (NIH) under Award Number T32HL079896. The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH.
Appendix S Appendix: R Code and Tutorial to COADVISE
S.1 R code for BestAIR trial study
We show R code of the analysis for our case study in Section 5, as a tutorial for the use of our R package Coadvise. For the demonstration purpose, we only present the code by Lasso selection. For other methods, readers can visit the package website (https://github.com/yiliu1998/Coadvise) or type code ?Coadvise in R (assuming the package is correctly installed) to see more details.
In the code below, A is the binary treatment (CPAP) indicator vector, Y is the continuous SPB value, and X is the covariate matrix (169 rows and 114 columns). We illustrate the additional arguments used in the code in the following Table S.1.
#### Install the Coadvise package and load the dataset
if(!require("devtools")) install.packages("devtools")
devtools::install_github("yiliu1998/Coadvise")
library(Coadvise) # load the R package
load("BestAIR.Rdata") # the original dataset
#### Analysis using Coadvise() function
result.cont <- Coadvise(Y=Y, A=A, X=X, trt.name=1, ctrl.name=0,
conti.out=TRUE, var.sel.method="Lasso",
lasso.family="gaussian", out1.model.aipw="linear",
out0.model.aipw="linear", MI.method="missForest",
seed=4399)
#### Print the ATE results
print(result.cont\$df.fit)
method tau se ci.lwr ci.upr p
1 Simple -3.82127278 1.6350819 -7.0259744 -0.6165711 0.01943644
2 ANCOVA -0.01156769 0.1567089 -0.3187115 0.2955761 0.94115646
3 ANHECOVA -0.07524842 0.1776052 -0.4233482 0.2728513 0.67179650
4 AIPW -0.38878980 0.2600589 -0.8984959 0.1209163 0.13491273
#### Print the mean outcome vector and covariance matrix by AIPW
print(result.cont\$AIPW.out.means)
\$mu
tau0 tau1
1 125.0551 124.6664
\$sigma
[,1] [,2]
[1,] 176.1196 106.39670
[2,] 106.3967 48.10336
Arguments Description Required Default Available options
A The treatment vector $\checkmark$
Y The outcome vector $\checkmark$
X The covariate vector/matrix/data frame $\checkmark$
trt.name The value of treatment in A $\checkmark$
ctrl.name The value of control group in A $\checkmark$
conti.out Whether the outcome Y is continuous $\checkmark$ TRUE, FALSE
var.sel.method Variable selection method $\times$ "Lasso" "No", "Lasso", "A.Lasso", "Corr.k", "Corr.xi", "Pre.test"
lasso.family Family used in Lasso regression $\times$ "gaussian" "gaussian", "binomial", "poisson", "multinomial" (Same as the glmnet package (Friedman et al., 2021))
A.lasso.family Family used in adaptive Lasso regression $\times$ "gaussian" "gaussian", "binomial", "poisson", "multinomial"
out1.model.aipw $\mathbb{E}\{Y(1)\mid\mathbf{X}\}$ model in AIPW $\times$ "linear" "linear", "logit", "probit" "log", "cloglog", "identity"
out0.model.aipw $\mathbb{E}\{Y(0)\mid\mathbf{X}\}$ model in AIPW $\times$ "linear" "linear", "logit", "probit" "log", "cloglog", "identity"
k Number of covariates with highest correlations with outcome $\times$ 1 Any positive integers less than the number of covariates
xi Lowest correlation with outcome $\times$ 0.25 Numbers between 0 and 1
pre.alpha Level of preliminary testing $\times$ 0.05 Numbers between 0 and 1
conf.level Level of returned confidence interval for ATE $\times$ 0.05 Numbers between 0 and 1
MI.method Missing data imputation method $\times$ "cc" "cc", "mice", "missForest", "ipw", "missInd"
seed seed used in set.seed() function $\times$ 4399 Any real numbers
Table S.1: Arguments in Coadvise() R function.
Both trt.name and ctrl.name need to be specified without default, which are the value of treatment and control groups in A, respectively. Note that A can have more than two levels, but we require users to input the treated (trt.name) and control (ctrl.name) each time for a pair-wise comparison. The conti.out must be specified from TRUE and FALSE as well, which allows the program to identify whether it needs to be processed as a categorical outcome for valid missing data imputation purpose. The var.sel.method argument is for the variable selection method, and we specify "Lasso" for its value. We also need to specify the model family via lasso.family if "Lasso" is used. The default is "gaussian" for using linear model with Gaussian random error assumption in Lasso regression, but users can choose other families, such as "binomial" for binary outcome if using logistic regression, and any families provided in the glmnet package.Friedman et al. (2021); Hastie et al. (2021) The arguments out0.model.aipw and out1.model.aipw are for specifying the link functions used in outcome models $\mu_{0}(\mathbf{X})$ and $\mu_{1}(\mathbf{X})$, respectively, in the AIPW estimator. For example, in the binary outcome case, we use "logit" to specify logistic regression models for both potential outcomes.
The function Coadvise() returns a list comprising three components: (i) a data frame containing results for the Simple, ANCOVA, ANHECOVA, and AIPW estimators, including point estimates (tau), standard errors (se), confidence intervals (ci.lwr and ci.upr), and p-values (p); (ii) the covariate data after variable selection, used in all ANCOVA, ANHECOVA, and AIPW estimators. For the AIPW estimator, two sets of covariate data are returned, as variable selection is performed separately for the two treatment groups; and (iii) the estimated mean outcomes vector and covariance matrix from the AIPW estimator. In this example, we display components (i) and (iii) from the returned list.
S.2 Algorithm example
In this section, we present the proposed COADVISE framework, providing technical details in Algorithm 1. For illustration, we use Lasso for variable selection and AIPW for estimation.
Let $\mathcal{R}^{p}$ denote the $p$-dimensional Euclidean space $(p>0)$. Define $N_{a}=\sum_{i=1}^{N}I(A_{i}=a)$ as the number of units with treatment assignment $A_{i}=a$, $s_{N}^{2}(\mathcal{W})$ as the sample variance of the i.i.d. data $\mathcal{W}=\{W_{1},\dots,W_{N}\}$, and $\text{cov}_{N}(\mathcal{W},\mathcal{V})$ as the sample covariance of the data $\mathcal{W}=\{W_{1},\dots,W_{N}\}$ and $\mathcal{V}=\{V_{1},\dots,V_{N}\}$. Similarly, $s_{N_{a}}^{2}(\mathcal{W})$ represents the sample variance for all $W_{i}\in\mathcal{W}$ with $A_{i}=a$, and $\text{cov}_{N_{a}}(\mathcal{W},\mathcal{V})$ denotes the sample covariance for all $W_{i}\in\mathcal{W}$ and $V_{i}\in\mathcal{V}$ with $A_{i}=a$, where $a=0,1$. Denote $\mathbb{P}_{N}[f(W)]=N^{-1}\sum_{i=1}^{N}f(W_{i})$ the empirical mean of any function of the i.i.d. observed data $\mathcal{W}$.
Algorithm 1 COADVISE by using Lasso for the variable selection and AIPW for ATE estimation
  Input: I.i.d. observed data $\mathcal{O}=\{(A_{i},Y_{i},\mathbf{X}_{i}),i=1,\dots,N\}$, where $\mathbf{X}_{i}\in\mathcal{R}^{p}$, $p>1$; a chosen generalized linear model (GLM) $g(\mathbf{X}^{\prime}\boldsymbol{\beta})$ with given link $g$, where $\boldsymbol{\beta}\in\mathcal{R}^{p}$ is the regression coefficient vector; a deviance function $D_{g}(\boldsymbol{\beta})$ related to the given link $g$; tuning parameters $\lambda_{0}$ and $\lambda_{1}$ in Lasso.
  Output: The estimate $\widehat{\tau}$, the variance estimation $\widehat{\mathbb{V}}(\tau)$, and the 95% confidence interval $\widehat{\mathcal{C}}_{0.95}(\tau)$ for ATE $\tau=\mathbb{E}\{Y(1)-Y(0)\}$.
  for $a=0,1$ do
$\widehat{\boldsymbol{\beta}}_{a}^{\text{lasso}}=\text{arg}\min_{\boldsymbol{\beta}_{a}}\left[\frac{1}{N_{a}}D_{g}(\boldsymbol{\beta}_{a})+\lambda_{a}\sum_{j=1}^{p}|\beta_{a}^{(j)}|\right],$
and let $\widehat{S}_{a}=\{j:\widehat{\beta}^{\text{lasso},(j)}_{a}\not=0,j=1,\dots,p\}$.
  end for
  for $a=0,1$ do
     $\widetilde{\boldsymbol{\beta}}_{a}=(\beta_{a}^{(1)}I(1\in\widehat{S}_{a}),\dots,\beta_{a}^{(p)}I(p\in\widehat{S}_{a}))^{\prime}$;
     Find $\widehat{\boldsymbol{\beta}}_{a}$ that solves the MLE equation of $\boldsymbol{\beta}_{a}$, e.g., $\mathbb{P}_{N}\left[I(A=a)\{Y-g(\mathbf{X}^{\prime}\widetilde{\boldsymbol{\beta}}_{a})\}\mathbf{X}^{\prime}\right]=\boldsymbol{0}$ for linear, logistic or Poisson regressions;
     Calculate
$\widehat{\theta}_{a}=\frac{1}{N}\sum_{i=1}^{N}\left[\frac{I(A_{i}=a)}{\widehat{P}(A=a)}\{Y_{i}-g(\mathbf{X}_{i}^{\prime}\widehat{\boldsymbol{\beta}}_{a})\}+g(\mathbf{X}_{i}^{\prime}\widehat{\boldsymbol{\beta}}_{a})\right],$
and let $\widehat{\mu}_{a}(\mathbf{X})=g(\mathbf{X}^{\prime}\widehat{\boldsymbol{\beta}}_{a})$.
  end for
  Calculate
$\widehat{v}_{11}$ $=(N/N_{1})s_{N_{1}}^{2}(Y-\widehat{\mu}_{1}(\mathbf{X}))+2\text{cov}_{N_{1}}(Y,\widehat{\mu}_{1}(\mathbf{X}))-s_{N}^{2}(\widehat{\mu}_{1}(\mathbf{X})),$
$\widehat{v}_{00}$ $=(N/N_{0})s_{N_{0}}^{2}(Y-\widehat{\mu}_{0}(\mathbf{X}))+2\text{cov}_{N_{0}}(Y,\widehat{\mu}_{0}(\mathbf{X}))-s_{N}^{2}(\widehat{\mu}_{1}(\mathbf{X})),$
$\widehat{v}_{10}$ $=\text{cov}_{N_{0}}(Y,\widehat{\mu}_{0}(\mathbf{X}))+\text{cov}_{N_{1}}(Y,\widehat{\mu}_{1}(\mathbf{X}))-\text{cov}_{N}\{\widehat{\mu}_{1}(\mathbf{X}),\widehat{\mu}_{0}(\mathbf{X})\}.$
  Return: $\widehat{\tau}=\widehat{\theta}_{1}-\widehat{\theta}_{0},\widehat{\mathbb{V}}(\tau)=N^{-1}(\widehat{v}_{11}-2\widehat{v}_{10}+\widehat{v}_{00}),\widehat{\mathcal{C}}_{0.95}(\tau)=\widehat{\tau}\pm 1.96\cdot\widehat{\mathbb{V}}(\tau).$
S.3 Comparison of Coadvise and RobinCar
In the code below, we use a toy example to compare and illustrate our Coadvise package and the well-known RobinCar package.
#### We generate a small data with continuous outcome
n <- 300
X <- matrix(rnorm(10*n), ncol=10) # related covariates
V <- matrix(rt(200*n, df=10), ncol=200) # noisy variables
A <- rbinom(n, size=1, prob=0.5) # treatment
Y <- X%*%(1:10)/10 + A*(4+X%*%(1:10)/50) + rnorm(n) # outcome
#### load R packages
if(!require("RobinCar")) { install.packages("RobinCar") }
library(RobinCar)
library(Coadvise)
?robincar_linear # the function in RobinCar we will use later
#### We first use Coadvise to perform a variable selection, using the Corr.k method with k=5, and results are saved in the following "step1" object
step1 <- Coadvise(Y=Y, X=X, A=A, trt.name=1, ctrl.name=0,
conti.out=TRUE, var.sel.method="Corr.k", k=5,
out1.model.aipw="linear", out0.model.aipw="linear",
seed=1335)
#### We then use covariates selected in step1 for ANCOVA for further analyses
X.selected <- step1\$stage1.covars\$ANCOVA
covar.names <- colnames(X.selected)
df.RobinCar <- data.frame(X.selected, A=as.factor(A), Y)
#### The simple estimator by RobinCar
rc.simple <- robincar_linear(df=df.RobinCar, treat_col="A", response_col="Y",
contrast_h = "diff", adj_method="ANOVA")
print(rc.simple)
contrast estimate se pval (2-sided)‘
<chr> <dbl> <dbl> <dbl>
1 treat 1 - 0 4.06 0.264 2.59e-53
#### The ANCOVA estimator by RobinCar
rc.ancova <- robincar_linear(df=df.RobinCar, treat_col="A", response_col="Y",
covariate_cols=covar.names, contrast_h = "diff", adj_method="ANCOVA")
print(rc.ancova)
contrast estimate se pval (2-sided)‘
<chr> <dbl> <dbl> <dbl>
1 treat 1 - 0 3.95 0.151 1.27e-151
#### The ANHECOVA estimator by RobinCar
rc.anhecova <- robincar_linear(df=df.RobinCar, treat_col="A", response_col="Y",
contrast_h = "diff", covariate_cols=covar.names, adj_method="ANHECOVA")
print(rc.anhecova)
contrast estimate se pval (2-sided)‘
<chr> <dbl> <dbl> <dbl>
1 treat 1 - 0 3.95 0.151 1.54e-151
#### The ATE results by Coadvise using the selected covariates from step1 with no variable selection (var.sel.method="No")
step2 <- Coadvise(Y=Y, X=X.selected, A=A, trt.name=1, ctrl.name=0,
conti.out=TRUE, var.sel.method="No",
out1.model.aipw="linear", out0.model.aipw="linear",
seed=1335)
print(step2\$df.fit)
method tau se ci.lwr ci.upr p
1 Simple 4.058923 0.2640761 3.541344 4.576503 0
2 ANCOVA 3.952416 0.1498078 3.658798 4.246034 0
3 ANHECOVA 3.951386 0.1497789 3.657825 4.244948 0
4 AIPW 3.951386 0.1494168 3.658535 4.244238 0
#### We finally check if using var.sel.method="No" in step2 is the same as the output of step1 (for Simple, ANCOVA, and ANHECOVA)
print(step1\$df.fit)
method tau se ci.lwr ci.upr p
1 Simple 4.058923 0.2640761 3.541344 4.576503 0
2 ANCOVA 3.952416 0.1498078 3.658798 4.246034 0
3 ANHECOVA 3.951386 0.1497789 3.657825 4.244948 0
4 AIPW 4.021081 0.1557283 3.715859 4.326303 0
From the results above, we observe that Coadvise and RobinCar yield the same point estimates and standard errors when using identical estimation methods. Additionally, our Coadvise package allows for variable selection and outputs the selected covariates. Re-running the Coadvise() function with the selected covariates, but without further variable selection, produces results consistent with the initial variable selection process.
Note that in the last two results (step1\$df.fit and step2\$df.fit), the AIPW estimators produce different point estimates. We would like to clarify that this occurs because for step2\$df.fit, we used the variables selected by ANCOVA for AIPW, whereas in step1\$df.fit, the variables selected in AIPW are different than ANCOVA, and the two outcome models in AIPW may use different sets of variables as predictors.
Appendix S Technical Proofs
In this section, we present theoretical investigations into the efficiency gain achieved by the AIPW estimator for the ATE in RCT data. We adhere to the notation, setups, and assumptions established in the main text of the paper.
S.1 Minimizing the variance over a class of augmented estimators
First, consider a general class of augmented estimators for $\tau$ defined as follows:
$\mathcal{G}=\{\widehat{\tau}^{\mathcal{G}}(g_{0},g_{1}):\mathbb{E}\{g_{a}(\mathbf{X})^{2}\}<\infty,a=0,1\},$ (B.1.1)
where
$\widehat{\tau}^{\mathcal{G}}=\widehat{\tau}^{\mathcal{G}}(g_{0},g_{1})=\frac{1}{N}\sum_{i=1}^{N}\left\{\frac{A_{i}Y_{i}}{\widehat{\pi}_{1}}-\frac{A_{i}-\widehat{\pi}_{1}}{\widehat{\pi}_{1}}g_{1}(\mathbf{X}_{i})\right\}-\frac{1}{N}\sum_{i=1}^{N}\left\{\frac{(1-A_{i})Y_{i}}{1-\widehat{\pi}_{1}}-\frac{\widehat{\pi}_{1}-A_{i}}{1-\widehat{\pi}_{1}}g_{0}(\mathbf{X}_{i})\right\}+o_{p}(N^{-1/2}),$
with $\widehat{\pi}_{1}=N_{1}/N$, the proportion of individuals with $A=1$ in the sample.
We note that when $g_{1}(\mathbf{X})=g_{0}(\mathbf{X})=0$, the estimator $\widehat{\tau}^{\mathcal{G}}$ reduces to $\widehat{\tau}_{\text{simple}}$. Furthermore, it is straightforward to show that for any choice of $g_{1}$ and $g_{0}$, $\widehat{\tau}^{\mathcal{G}}\to_{p}\tau$, using standard conditioning arguments and large sample theory.
The estimator $\widehat{\tau}^{\mathcal{G}}$ provides a general form of the AIPW estimator, as it consists of $\widehat{\tau}_{\text{simple}}$ augmented by functions $g_{0}$ and $g_{1}$. The term $o_{p}(N^{-1/2})$ indicates that two estimators in the class $\mathcal{G}$ differ by a negligible term that converges to 0 in probability with zero asymptotic variance. As can be observed, $\mathcal{G}$ is a large (uncountable) set, containing functions from an infinite-dimensional space.
When $g_{a}(\mathbf{X})=\mathbb{E}(Y(a)\mid\mathbf{X})$, representing the fully specified model for $Y(a)$, $a=0,1$, the estimator $\widehat{\tau}^{\mathcal{G}}$ achieves the semiparametric efficiency bound (Hahn, 1998; Hirano et al., 2003). However, since the true model of $Y(a)$ as a function of $\mathbf{X}$ is unknown, achieving this bound is generally not expected. Nevertheless, we can still improve efficiency through “appropriate augmentation,” even without precise model specification, compared to $\widehat{\tau}_{\text{simple}}$. This is our objective in leveraging covariate information to gain efficiency (Bannick et al., 2023).
In practice, we focus on a subclass of $\mathcal{G}$, denoted as $\mathcal{G}_{\boldsymbol{\beta}}^{g}\subset\mathcal{G}$, where $g_{a}(\mathbf{X})=g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})$ follows a generalized linear model (GLM) for $Y(a)$, for $a=0,1$. This subclass is characterized by the given function $g$, for all vectors $\boldsymbol{\beta}\in\mathcal{R}^{p}$ such that $\mathbb{E}\{g(\mathbf{X}^{\prime}\boldsymbol{\beta})^{2}\}<\infty$. If the true model for $Y(a)$ belongs to $\mathcal{G}_{\boldsymbol{\beta}}^{g}$, then there exists a fixed $\boldsymbol{\beta}_{a}^{0}\in\mathcal{R}^{p}$ such that $g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{0})=\mathbb{E}(Y(a)\mid\mathbf{X})$. This subclass $\mathcal{G}_{\boldsymbol{\beta}}^{g}$ has practical appeal because, in many applications, we often fit both $Y(0)$ and $Y(1)$ using the same GLM with an appropriate link function $g$.
Next, we denote an AIPW estimator belonging to $\mathcal{G}_{\boldsymbol{\beta}}^{g}$ by $\widehat{\tau}^{g}$, highlighting the dependence on the link function $g$. Using argument $\mathbb{V}(\cdot)=\mathbb{E}\{\mathbb{V}(\cdot\mid\mathbf{X},Y(0),Y(1))\}+\mathbb{V}\{\mathbb{E}(\cdot\mid\mathbf{X},Y(0),Y(1))\}$, we have the asymptotic variance of $\widehat{\tau}^{g}$ as follows:
$\mathbb{V}(\widehat{\tau}^{g})$ $=\mathbb{V}\left\{\frac{AY}{\pi_{1}}-\frac{A-\pi_{1}}{\pi_{1}}g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{1}^{*})-\frac{(1-A)Y}{1-\pi_{1}}+\frac{\pi_{1}-A}{1-\pi_{1}}g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}^{*})\right\}$
$=\mathbb{V}\left\{\frac{AY(1)}{\pi_{1}}-\frac{A-\pi_{1}}{\pi_{1}}g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{1}^{*})-\frac{(1-A)Y(0)}{1-\pi_{1}}+\frac{\pi_{1}-A}{1-\pi_{1}}g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}^{*})\right\}$
$=\mathbb{E}\left\{\frac{1-\pi_{1}}{\pi_{1}}\{Y(1)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{1}^{*})\}^{2}+\frac{\pi_{1}}{1-\pi_{1}}\{Y(0)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}^{*})\}^{2}\right\}+\mathbb{V}\{Y(1)-Y(0)\},$ (B.1.2)
where $\boldsymbol{\beta}_{a}^{*}$ is the probability limit of $\widehat{\boldsymbol{\beta}}_{a}$, and $\widehat{\boldsymbol{\beta}}_{a}$ is an estimator (e.g., the ordinary least square [OLS] estimator) used to estimate $\boldsymbol{\beta}_{a}$ ($a=0,1$).
Therefore, to minimize (S.1) over choices of $({\boldsymbol{\beta}_{0}^{*}}^{\prime},{\boldsymbol{\beta}_{1}^{*}}^{\prime})^{\prime}$, it is equivalent to minimize
$\mathbb{E}\left\{\frac{1-\pi_{1}}{\pi_{1}}\{Y(1)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{1}^{*})\}^{2}+\frac{\pi_{1}}{1-\pi_{1}}\{Y(0)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}^{*})\}^{2}\right\}.$ (B.1.3)
Then, it is natural to consider setting the derivative with respect to $({\boldsymbol{\beta}_{0}^{*}}^{\prime},{\boldsymbol{\beta}_{1}^{*}}^{\prime})^{\prime}$ to 0, i.e., to solve equations
00 0 $=\frac{1-\pi_{1}}{\pi_{1}}\mathbb{E}\left\{g_{\boldsymbol{\beta}_{1}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{1}^{*})[Y(1)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{1}^{*})]\right\},$
00 0 $=\frac{\pi_{1}}{1-\pi_{1}}\mathbb{E}\left\{g_{\boldsymbol{\beta}_{0}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}^{*})[Y(0)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}^{*})]\right\},$ (B.1.4)
where $g_{\boldsymbol{\beta}}(\mathbf{X}^{\prime}\boldsymbol{\beta}^{*})$ denotes $\dfrac{\partial}{\partial\boldsymbol{\beta}}g(\mathbf{X}^{\prime}\boldsymbol{\beta})\bigg{|}_{\boldsymbol{\beta}=\boldsymbol{\beta}^{*}}$.
Observe that the solution of the following equation, say, $\widehat{\boldsymbol{\beta}}_{a}$,
$\frac{1}{N}\sum_{i=1}^{N}I(A_{i}=a)g_{\boldsymbol{\beta}_{a}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})[Y_{i}-g(\mathbf{X}_{i}^{\prime}\boldsymbol{\beta}_{a})]=0,$ (B.1.5)
converges to the solution of (S.1), for $a=0,1$. This is because when the outcome model is possibly misspecified, i.e, when $g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})\not=\mathbb{E}(Y(a)\mid\mathbf{X})$, the left-hand side of the above equation converges in probability to
$\pi_{a}\mathbb{E}\left\{g_{\boldsymbol{\beta}_{a}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})[\mathbb{E}(Y(a)\mid\mathbf{X})-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})]\right\}=\pi_{a}\mathbb{E}\left\{g_{\boldsymbol{\beta}_{a}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})[Y(a)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})]\right\},$ (B.1.6)
for $a=0,1$, which differs from (S.1) only by a constant. Therefore, setting (B.1.6) to be 0 results in the same solution to (S.1). We refer to this $\widehat{\boldsymbol{\beta}}_{a}$ as the OLS estimator in the following text.
Moreover, to verify whether solving (S.1) always uniquely minimizes (B.1.3), we check if the second-order derivative of (B.1.3) with respect to $({\boldsymbol{\beta}_{0}^{*}}^{\prime},{\boldsymbol{\beta}_{1}^{*}}^{\prime})^{\prime}$ is semi-positive definite (a sufficient but not necessary condition). The second-order derivative can be expressed as $\text{diag}(\boldsymbol{\Sigma}_{0},\boldsymbol{\Sigma}_{1})$ (diagonal matrix of two blocks), where
$\boldsymbol{\Sigma}_{a}$ $=2\frac{1-\pi_{a}}{\pi_{a}}\mathbb{E}\left\{-\{Y(a)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})\}g_{\boldsymbol{\beta}_{a}\boldsymbol{\beta}_{a}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})+g_{\boldsymbol{\beta}_{a}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})g_{\boldsymbol{\beta}_{a}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})^{\prime}\right\}$
$=2\frac{1-\pi_{a}}{\pi_{a}}\mathbb{E}\left\{g_{\boldsymbol{\beta}_{a}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})g_{\boldsymbol{\beta}_{a}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})^{\prime}\right\}-2\frac{1-\pi_{a}}{\pi_{a}}\mathbb{E}\left\{\{\mathbb{E}(Y(a)\mid\mathbf{X})-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})\}g_{\boldsymbol{\beta}_{a}\boldsymbol{\beta}_{a}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})\right\},$ (B.1.7)
for $a=0,1$, where $g_{\boldsymbol{\beta}\boldsymbol{\beta}}(\mathbf{X}^{\prime}\boldsymbol{\beta}^{*})$ denotes $\dfrac{\partial^{2}}{\partial\boldsymbol{\beta}\partial\boldsymbol{\beta}^{\prime}}g(\mathbf{X}^{\prime}\boldsymbol{\beta})\bigg{|}_{\boldsymbol{\beta}=\boldsymbol{\beta}^{*}}$. We note that the first term of (S.1) is always positive definite, while for the second term, further discussion is provided in Sections S.2 and S.3.
S.2 Guaranteed efficiency gain using linear models
Consider the special case where $g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})=\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}$, the identical link function. In this case, the second term in (S.1) is always zero because $g_{\boldsymbol{\beta}_{a}\boldsymbol{\beta}_{a}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})=\boldsymbol{0}$. Thus, the minimum of (B.1.3) is always achieved by the OLS estimator $\widehat{\boldsymbol{\beta}}_{a}$ for $\boldsymbol{\beta}_{a}$. The efficiency gain is guaranteed, as $\widehat{\tau}_{\text{simple}}\in\mathcal{G}_{\boldsymbol{\beta}}^{g}$ by setting $\boldsymbol{\beta}_{0}=\boldsymbol{\beta}_{1}=\boldsymbol{0}$. Consequently, the estimator in the class $\mathcal{G}_{\boldsymbol{\beta}}^{g}$, which minimizes the variance, must be at least as efficient as $\widehat{\tau}_{\text{simple}}$.
Furthermore, this result can easily be extended situations where a variable selection step is performed before AIPW estimation. Suppose we select $s$ variables from $\mathbf{X}$ for $Y(a)$ model, where $0<s<p$. When fitting the OLS estimator, the only difference is that $\boldsymbol{\beta}\in\mathcal{R}^{s}$ instead of $\mathcal{R}^{p}$ for estimators in $\mathcal{G}^{\boldsymbol{\beta}}_{g}$. Here, $\widehat{\tau}_{\text{simple}}$ remains a special case of estimators in $\mathcal{G}^{\boldsymbol{\beta}}_{g}$. Therefore, the proof remains straightforward.
Moreover, this illustrates that with more covariates (i.e., higher dimensions of $\boldsymbol{\beta}_{a}$), there is a greater opportunity to achieve higher efficiency. Without loss of generality, assume the first $s$ variables of $\mathbf{X}$ are selected. We can view the lower dimension $s$ as $\boldsymbol{\beta}_{a}\in\mathcal{R}^{p}$ where $\beta_{a}^{(s+1)}=\dots=\beta_{a}^{(p)}=0$ are fixed, with $\beta_{a}^{(j)}$ denoting the $j$-th element of $\boldsymbol{\beta}_{a}$. Therefore, the class of estimators $\mathcal{G}^{\boldsymbol{\beta}}_{g}$ when $\boldsymbol{\beta}_{a}\in\mathcal{R}^{s}$ can be viewed as a subset of $\mathcal{G}^{\boldsymbol{\beta}}_{g}$ when $\boldsymbol{\beta}_{a}\in\mathcal{R}^{p}$. Naturally, the variance by a minimizer found from a larger class must be at least as low as that by a minimizer found from a smaller subclass. However, when the sample size is not sufficiently large, including many covariates can negatively impact the finite-sample performance of the estimator, as all of the aforementioned results are based on large-sample asymptotics.
S.3 Conditions on efficiency gain under a nonlinear GLM
nvestigating efficiency gains with a general link function $g$ is more challenging due to several factors: (i) the usual maximum likelihood estimator (MLE) for regression coefficients in a general GLM is not obtained by solving (B.1.5) to minimize the squared loss in (B.1.3); (ii) for a general link function $g$, the second-order derivative (S.1) is not always positive-definite; (iii) $\widehat{\tau}_{\text{simple}}$ may not even belong to $\mathcal{G}_{\boldsymbol{\beta}}^{g}$ for certain link functions.
For (iii), consider logistic regression. The logistic link function is given by $g(u)=\{1+e^{-u}\}^{-1}>0$, which implies that the class $\mathcal{G}_{\boldsymbol{\beta}}^{g}$, defined by the logistic function $g$, does not include $\widehat{\tau}_{\text{simple}}$. This is because the influence function of $\widehat{\tau}_{\text{simple}}$ cannot be approximated by any influence functions of estimators in $\mathcal{G}_{\boldsymbol{\beta}}^{g}$ (in the sense that the difference between the two influence functions can be $O_{p}(N^{-r})$ with an $r\geq 1/2$). Achieving such an approximation would require $g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})\to_{p}0$ for all $\mathbf{X}\in\mathcal{X}$, which in trun would imply that there exists some fixed $\boldsymbol{\beta}_{a}^{*}$ such that $\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*}\to_{p}-\infty$ for all $\mathbf{X}\in\mathcal{X}$. Consequently, $\widehat{\tau}_{\text{simple}}$ cannot even lie on the boundary (or closure) of $\mathcal{G}_{\boldsymbol{\beta}}^{g}$.
To further explore the conditions for efficiency gain, we examine the difference between the asymptotic variances of $\widehat{\tau}_{\text{simple}}$ and $\widehat{\tau}^{g}$, given by:
$\mathbb{V}(\widehat{\tau}_{\text{simple}})-\mathbb{V}(\widehat{\tau}^{g})$ $=\mathbb{E}\left\{\frac{1-\pi_{1}}{\pi_{1}}Y(1)^{2}+\frac{\pi_{1}}{1-\pi_{1}}Y(0)^{2}\right\}-\mathbb{E}\left\{\frac{1-\pi_{1}}{\pi_{1}}\{Y(1)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{1}^{*})\}^{2}+\frac{\pi_{1}}{1-\pi_{1}}\{Y(0)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}^{*})\}^{2}\right\}$
$=\mathbb{E}\left\{\frac{1-\pi_{1}}{\pi_{1}}g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{1})\{2Y(1)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{1})\}+\frac{\pi_{1}}{1-\pi_{1}}g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{0})\{2Y(0)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{0})\}\right\}$
$=\frac{1-\pi_{1}}{\pi_{1}}\underbrace{\mathbb{E}\left\{g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{1})[2Y(1)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{1})]\right\}}_{I_{1}}+\frac{\pi_{1}}{1-\pi_{1}}\underbrace{\mathbb{E}\left\{g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{0})[2Y(0)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{0})]\right\}}_{I_{0}}.$ (B.3.1)
For both $a=0,1$, we note that $I_{a}=\mathbb{E}\left\{g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})[2Y(a)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})]\right\}=\mathbb{E}\left\{g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})[2\mathbb{E}(Y(a)\mid\mathbf{X})-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})]\right\}=\mathbb{E}\{I_{a}(\mathbf{X})\}$, where $I_{a}(\mathbf{X})=g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})[2\mathbb{E}(Y(a)\mid\mathbf{X})-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})]$. Therefore, $I_{a}(\mathbf{X})$ achieves its maximum when $g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})=\mathbb{E}(Y(a)\mid\mathbf{X})$. This implies that when $g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})$ is the fully specified model for $\mathbb{E}(Y(a)\mid\mathbf{X})$, $\mathbb{V}(\widehat{\tau}_{\text{simple}})-\mathbb{V}(\widehat{\tau}^{g})$ reaches its maximal value $\mathbb{E}\{\mathbb{E}\{Y(a)^{2}\mid\mathbf{X}\}\}=\mathbb{E}\{Y(a)^{2}\}$, corresponding to the highest efficiency gain. Additionally, in this case, the second term in the (S.1) of the squared loss is always zero, which ensures $\boldsymbol{\beta}_{a}^{*}$ is the variance minimizer. However, this is typically unrealistic in practice, and if the model is not well-specified, it is possible that (S.3) can be negative. In the following, we outline some conditions for achieving efficiency gains.
Condition S.3.1.
For a nonlinear link function $g$, with probability 1, and there exists some $\boldsymbol{\beta}_{a}^{*}$ ($a=0,1$), we have: (a) the sign of $g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})$ must match that of $\mathbb{E}(Y(a)\mid\mathbf{X})$; and (b) $|g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})|\leq 2|\mathbb{E}(Y(a)\mid\mathbf{X})|$.
If Condition S.3.1 holds, then it is clear that (S.3) is always non-negative. This condition may be mild for some positive categorical outcomes. In the following illustration, we consider logistic regression as a more specific example, as it is commonly used in binary outcome modelling. For both $a=0,1$, we further write
$g_{\boldsymbol{\beta}_{a}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})$ $=\frac{\exp(-\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})}{\{1+\exp(-\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})\}^{2}}\mathbf{X}^{\prime}=g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})\{1-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})\}\mathbf{X}^{\prime},$
$g_{\boldsymbol{\beta}_{a}\boldsymbol{\beta}_{a}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})$ $=\frac{\exp(-\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})\{\exp(-\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})-1\}}{\{1+\exp(-\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})\}^{3}}\mathbf{X}\mathbf{X}^{\prime}=g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})\{1-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})\}\{1-2g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})\}\mathbf{X}\mathbf{X}^{\prime}.$ (B.3.2)
Therefore, for any general $\mathbf{X}\not=\boldsymbol{0}$, we have $g_{\boldsymbol{\beta}_{a}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})\not=\boldsymbol{0}$, since the logistic function satisfies $10<g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})<$ 1. Additionally, $g_{\boldsymbol{\beta}_{a}\boldsymbol{\beta}_{a}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})=\boldsymbol{0}$ only when $\boldsymbol{\beta}_{a}^{*}=\boldsymbol{0}$ (so $g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})=1/2$); for other $\boldsymbol{\beta}_{a}^{*}$, the sign of $g_{\boldsymbol{\beta}_{a}\boldsymbol{\beta}_{a}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})$ can be either positive or negative. We now introduce Condition S.3.2 for using logistic regression to achieve efficiency gain.
Condition S.3.2 (Logistic outcome models).
(i) There exists a constant $\delta>0$ such that $\mathbb{E}(Y(1)\mid\mathbf{X})=1-\mathbb{E}(Y(0)\mid\mathbf{X})\geq\delta$ with probability 1 over $\mathbf{X}\in\mathcal{X}$; (ii) For logistic regression model $g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})$ we assigned to $Y(a)$, $|f_{a}(\mathbf{X})-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})|<g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})\{1-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})\}|1-2g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})|^{-1}$ with probability 1.
If Condition S.3.2 is satisfied, it is straightforward to verify that the OLS estimator $\widehat{\boldsymbol{\beta}}_{a}$ minimizes the variance. We also have additional intuitions about Condition S.3.2 as follows. Condition S.3.2(i) requires the true proportion of positive responses under both treatment and control is non-zero, which is a mild assumption for many binary outcomes in practice. In Condition S.3.2(ii), the left-hand side of the inequality in (ii) can be interpreted as the (absolute local) bias of the model $g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})$ at $\mathbf{X}$. This means that if the bias can be controlled by the right-hand side almost surely, along with (i), the proof in Section S.2 regarding the minimizer still applies to logistic regression.
Figure S.3.1 illustrates the right-hand side of inequality in (ii). From the figure, if $\boldsymbol{\beta}_{a}^{*}$ is such that $g(\mathbf{X}^{\prime}\boldsymbol{\beta}^{*})\approx 1/2$, the (absolute value of) bias is allowed to be relatively large. However, when $g(\mathbf{X}^{\prime}\boldsymbol{\beta}^{*})$ is either smaller or larger, the bias must be smaller, approaching zero when $g(\mathbf{X}^{\prime}\boldsymbol{\beta}^{*})\approx 0$ or $1$. In other words, when the true proportion of $Y(a)$ is very small or vary large, the model must better approximate the true value. But when the true proportion is around $50\%$, more error in the posited model is acceptable. The intuition is when the model $g(\mathbf{X}^{\prime}\boldsymbol{\beta}^{*})$ approaches a medium value, the uncertainty regarding whether the true $Y(a)$ is 00 or $1$ increases.
Figure S.3.1: An illustration for the bound of absolute bias in Condition S.3.2(ii) (right-hand side of the inequality). Panel (a) provides the absolute biases over $g(\mathbf{X}^{\prime}\boldsymbol{\beta})\in[0.1,0.9]$, which ranges from $[0.11,\infty)$. When $g(\mathbf{X}^{\prime}\boldsymbol{\beta})=0.5$, the absolute bias is actually $\infty$; Panel (b) provides a further look for $g(\mathbf{X}^{\prime}\boldsymbol{\beta})\in[0.1,0.3]$.
Moreover, it may be natural to consider an alternative coefficient estimation strategy to gain efficiency without requiring the conditions mentioned above for a general GLM. Revisiting (S.3), for $a=0,1$, write
$I_{a}$ $=\mathbb{E}\{g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})[2Y(a)-2g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})+g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})]\}=2\mathbb{E}\{g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})[Y(a)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})]\}+\mathbb{E}\{g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})^{2}\}.$
Therefore, to ensure $I_{a}\geq 0$, consider $\widetilde{\boldsymbol{\beta}}_{a}$ as the solution to the following joint equations
$\frac{1}{N}\sum_{i=1}^{N}I(A_{i}=a)g(\mathbf{X}_{i}^{\prime}\boldsymbol{\beta}_{a})[Y_{i}-g(\mathbf{X}_{i}^{\prime}\boldsymbol{\beta}_{a})]\mathbf{X}^{\prime}=\boldsymbol{0}.$ (B.3.3)
Here $\mathbf{X}$ should include an intercept term so that setting (B.3.3) to zero implies $\dfrac{1}{N}\displaystyle\sum_{i=1}^{N}I(A_{i}=a)g(\mathbf{X}_{i}^{\prime}\boldsymbol{\beta}_{a})[Y_{i}-g(\mathbf{X}_{i}^{\prime}\boldsymbol{\beta}_{a})]=0$. Therefore, the probability limit of $\widetilde{\boldsymbol{\beta}}_{a}$, denoted by ${\boldsymbol{\beta}}_{a}^{**}$, can satisfy $\mathbb{E}\{g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{**})[Y(a)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{**})]\}=0$, and thus, for $a=0,1$, we have
$I_{a}=2\mathbb{E}\{g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{**})[Y(a)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{**})]\}+\mathbb{E}\{g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{**})^{2}\}=\mathbb{E}\{g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{**})^{2}\}\geq 0.$
However, if we plug this ${\boldsymbol{\beta}}_{a}^{**}$ into (S.1), in general, $\mathbb{E}\{g_{\boldsymbol{\beta}_{a}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{**})[Y(a)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{**})]\}\not=0$, meaning that $\widetilde{\boldsymbol{\beta}}_{a}$ is not the global (optimal) variance minimizer among all $\widehat{\tau}^{g}\in\mathcal{G}_{\boldsymbol{\beta}}^{g}$. Another issue with this approach is that $\widetilde{\boldsymbol{\beta}}_{a}$ is not the MLE for the regression coefficients of a nonlinear GLM, so it does not guarantee unbiased predictions of the outcomes (Van Lancker et al., 2024). Finally, in practice, solving (B.3.3) can be unstable. For example, in logistic regression, a $\widetilde{\boldsymbol{\beta}}_{a}$ with some large elements can make $g(\mathbf{X}^{\prime}\widetilde{\boldsymbol{\beta}}_{a})\approx 0$, but this $\widetilde{\boldsymbol{\beta}}_{a}$ might deviate significantly from the true value, leading to poor statistical properties. For these reasons, we did not incorporate this approach into our framework.
Appendix S Additional Simulation Results
S.1 Additional details of the data generating process
Recall that we generate the binary treatment assignment by $A\sim\text{Bern}(0.5)$, and we consider two sets of covariates $\mathbf{X}=(X_{1},\dots,X_{5})^{\prime}$ and $\mathbf{V}=(V_{1},\dots,V_{50})^{\prime}$, where $(X_{1},X_{2})\sim\mathcal{N}_{2}(\boldsymbol{\mu}_{X},\boldsymbol{\Sigma}_{X})$ is a bivariate normal distribution with $\boldsymbol{\mu}_{X}=(0,0)^{\prime}$, $\boldsymbol{\Sigma}_{X}=\begin{pmatrix}1&0.8\\ 0.8&1\end{pmatrix}$, $X_{3}\sim\mathcal{N}(0,1)$, $X_{4}\sim t_{df=10}$, and $X_{5}\sim\text{Bin}(10,0.2)-2$. Additionally, $\mathbf{V}\sim\mathcal{N}_{50}(\boldsymbol{\mu}_{V},\boldsymbol{\Sigma}_{V})$, where $\boldsymbol{\mu}_{V}=(\underbrace{1,\dots,1}_{50\text{ elements}})^{\prime}$, and $\boldsymbol{\Sigma}_{V}$ is the correlation matrix of matrix $\boldsymbol{B}$, with
$\boldsymbol{B}=\underbrace{\begin{pmatrix}0.10&0.10&\dots&0.10\\ 0.11&0.11&\dots&0.11\\ \vdots&\vdots&\ddots&\vdots\\ 0.59&0.59&\dots&0.59\end{pmatrix}}_{50\text{ columns}}+2\boldsymbol{I}_{50},$
where $\boldsymbol{I}_{50}$ is the $50\times 50$ identity matrix.
We specify the following two models for potential outcomes:
Continuous outcome: $Y(a)=50+20\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}+a\delta(\mathbf{X})+\epsilon,$
Binary outcome: $Y(a)\sim\text{Bern}(e(\mathbf{X},a)),\text{ with }e(\mathbf{X},a)=\{1+\exp(-20\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}-a[3+\delta(\mathbf{X})])\}^{-1},$
and we consider two specifications of $\delta(\mathbf{X})$:
  • linear: $\delta(\mathbf{X})=c_{1}\mathbf{X}\boldsymbol{\beta}_{1}$, and
  • nonlinear: $\delta(\mathbf{X})=c_{2}\mathbf{W}\boldsymbol{\beta}_{1}$, with $\mathbf{W}=(X_{1}^{2},\sin(X_{2}^{2}),X_{3}X_{4},X_{4}X_{5},X_{5}^{2})$,
where $\boldsymbol{\beta}_{0}=(1,1,1,1,1)^{\prime}$, and $\boldsymbol{\beta}_{1}=(2,2,3,3,3)^{\prime}$. For continuous outcomes, $(c_{1},c_{2})=(8.15,1)$; for binary outcomes, $(c_{1},c_{2})=(10,1)$. The true ATEs, evaluating by the super-population data ($N=10^{8}$), are $\tau=8.15$ for both continuous outcomes and $\tau=0.11$ for both binary outcomes.
S.2 Additional simulation results
Figure S.2.1: Simulation results under continuous outcome, linear $\delta(\mathbf{X})$ and $N=100$. In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
Figure S.2.2: Simulation results under continuous outcome, linear $\delta(\mathbf{X})$ and $N=500$. In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
Figure S.2.3: Simulation results under continuous outcome, nonlinear $\delta(\mathbf{X})$ and $N=100$. In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
Figure S.2.4: Simulation results under continuous outcome, nonlinear $\delta(\mathbf{X})$ and $N=500$. In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
Figure S.2.5: Simulation results under binary outcome, linear $\delta(\mathbf{X})$ and $N=100$. In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
Figure S.2.6: Simulation results under binary outcome, linear $\delta(\mathbf{X})$ and $N=500$. In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
Figure S.2.7: Simulation results under binary outcome, nonlinear $\delta(\mathbf{X})$ and $N=100$. In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
Figure S.2.8: Simulation results under binary outcome, nonlinear $\delta(\mathbf{X})$ and $N=500$. In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.

### v4

@@
 The establishment of the American Statistical Association (ASA) Biopharmaceutical Section (BIOP) Covariate Adjustment Scientific Working Group (SWG) (https://carswg.github.io/) further underscores the growing focus on advancing statistical methods for covariate adjustment in clinical trials.
-Despite these regulatory endorsements, significant challenges remain due to high-dimensional covariates and nonlinear relationships.
-These complexities necessitate robust and comprehensive analytical strategies.
-However, the comparative advantages and limitations of existing methods remain insufficiently explored, leaving practitioners with limited guidance and uncertainty in selecting the most appropriate approaches.
-For instance, in the case study presented in Section 5, the Best Apnea Interventions for Research (BestAIR) trial includes data from $N=169$ participants and $p=114$ covariates, making $p$ comparable to $N$.
-When adjusting for covariates using data from a single treatment arm—where $N_{1}=83$ (treated) and $N_{0}=86$ (controls)—$p$ exceeds the sample sizes.
-This high-dimensional setting can lead to biased and imprecise covariate-adjusted average treatment effect (ATE) estimates, as the sample size may be insufficient for reliable parameter estimation.
-Nevertheless, prior studies, such as Gao et al.
-(2024b), have identified baseline covariate imbalances in the BestAIR trial data, underscoring the need for efficient covariate-adjusted methods.
-This motivates our approach to improving estimation precision while addressing challenges aforementioned.
+[…]
 This paper presents a unified framework for covariate adjustment in RCT analysis, with a primary focus on selecting important covariates for reducing data dimensionality for efficiency gains.
@@
 (2024) provided a unified perspective on the relationships between the simple estimator, ANCOVA, and ANHECOVA, providing conditions under which ANHECOVA uniformly dominates other methods in terms of efficiency and precision.
-Beyond ANCOVA and ANHECOVA, a prominent method in the analysis of RCTs is augmented inverse probability weighting (AIPW) (Ye et al., 2023b, a; Robins et al., 1994; Tsiatis, 2006; Glynn and Quinn, 2010; Cao et al., 2009).
-The AIPW estimator combines outcome modeling with inverse probability weighting, effectively leveraging information from both observed outcomes and treatment assignment probabilities.
-This dual-model approach enables AIPW to achieve semiparametric efficiency, which means it attains the semiparametric variance lower bound when the outcome models are correctly specified.
-As a result, AIPW offers the highest asymptotic efficiency among regular ATE estimators in RCTs (Hahn, 1998; Hirano et al., 2003; Kang et al., 2007; Glynn and Quinn, 2010).
-One of the defining features of AIPW is its robustness to model misspecification in the context of RCTs.
-The randomization inherent in trial design allows AIPW to consistently estimate the ATE even when the outcome model is incorrectly specified.
+Beyond ANCOVA and ANHECOVA, a prominent method in the analysis of RCTs is augmented inverse probability weighting (AIPW) (Robins et al., 1994; Tsiatis, 2006; Glynn and Quinn, 2010; Cao et al., 2009; Kang et al., 2007).
+The AIPW estimator combines outcome modeling with inverse treatment probability weighting, effectively leveraging information from both observed outcomes and treatment assignment probabilities.
+This dual-model approach allows AIPW to achieve semiparametric efficiency—that is, to attain the semiparametric variance lower bound—when both the treatment and outcome models are correctly specified.
+In the setting of RCTs considered in this paper, the treatment assignment mechanism is known by design, so only correct specification of the outcome model is required.
+As a result, AIPW can be asymptotically most efficient among regular estimators of the ATE in this context.
+Moreover, AIPW remains consistent even when the outcome model is misspecified, owing to the robustness conferred by randomization.
 This robustness, combined with its efficiency, makes AIPW particularly attractive in practice.
+Recent work has extended AIPW methodology to accommodate more complex randomization schemes.
+For example, Bannick et al.
+(2023) consider covariate-adaptive randomization, in which treatment assignment probabilities are stratified and potentially correlated with baseline covariates.
+In certain structured settings, such as when allocation probabilities are constant within strata, AIPW can still achieve semiparametric efficiency with only a correctly specified outcome model.
+However, under more general covariate-adaptive designs, correct modeling of the treatment assignment mechanism becomes necessary, mirroring challenges found in observational studies.
+Additionally, Ye et al.
+(2023a, b) propose robust variance estimators for AIPW under multi-arm RCT settings, which we adopt in our implementation (see Section 2.4).
 Moreover, AIPW offers flexibility by accommodating a wide range of parametric and nonparametric models for outcome estimation, making it more adaptable than ANCOVA or ANHECOVA, which depend on fixed linear model structures.
-Interestingly, when linear regression is used for outcome modeling within the AIPW framework, the AIPW estimator becomes equivalent to ANHECOVA.
+Interestingly, when linear regression is used for outcome modeling within the AIPW framework, the AIPW estimator becomes equivalent to ANHECOVA in RCTs considered in this paper.
 In this sense, ANHECOVA can be viewed as a special case of AIPW (Van Lancker et al., 2023).
 However, this equivalence also underscores a limitation of ANHECOVA: if the linear model is misspecified, the efficiency gains achieved by ANHECOVA are limited compared to those of AIPW when the latter employs correctly specified—or at least better specified—models.
-For this reason, AIPW is often regarded as a fully adjusted estimator in the literature, capable of optimizing efficiency without being overly reliant on strict parametric assumptions (Li et al., 2023).
+For this reason, AIPW is also regarded as a fully adjusted estimator in the literature, capable of optimizing efficiency without being overly reliant on strict parametric assumptions (Li et al., 2023).
 As mentioned in Section 1.1, all covariate adjustment methods encounter difficulties in ensuring parameter convergence when faced with high-dimensional covariate data.
@@
 Our work builds on existing methodologies by addressing several limitations and enhancing flexibility in the analysis of RCT data.
-First, we extend traditional covariate adjustment methods, such as ANCOVA, ANHECOVA, and AIPW, by incorporating a variety of variable selection techniques, including Lasso, adaptive Lasso, marginal correlation, and preliminary testing into the analysis stage of RCTs.
+First, under a complete randomization scheme, we extend traditional covariate adjustment methods, such as ANCOVA, ANHECOVA, and AIPW, by incorporating a variety of variable selection techniques, including Lasso, adaptive Lasso, marginal correlation, and preliminary testing into the analysis stage of RCTs.
 This integration addresses the challenges of high-dimensional covariates, enabling efficient and robust estimation of treatment effects.
-Unlike existing tools such as RobinCar, our framework explicitly supports variable selection, allowing users to identify covariates most predictive of outcomes, thereby improving estimation efficiency.
-Second, since missing data are common in RCTs, we provide several missing data imputation methods in our Coadvise R package for user convenience.
-These methods allow users to impute missing values in the outcome and/or covariates before applying the COADVISE framework.
-Our package includes complete-case analysis, multiple imputation by chained equations (mice), random forest-based imputation (missForest), inverse probability weighting (ipw), and the use of missingness indicators (missInd).
-These methods ensure robustness across a range of assumptions about missing data mechanisms, such as missing completely at random (MCAR) or missing at random (MAR).
-Notably, while tools like RobinCar focus primarily on covariate adjustment, they lack native support for missing data imputation, distinguishing Coadvise as a more comprehensive solution.
-To further highlight the differences between Coadvise and RobinCar, we provide a detailed comparison of their capabilities within the context of covariate adjustment for RCTs.
-Our framework not only supports multi-valued treatments, missing data imputation, and variable selection but also facilitates interoperability by enabling users to export selected covariates for subsequent analysis in RobinCar or custom code.
-This flexibility ensures that our framework complements rather than replaces existing tools, promoting seamless integration into diverse analytical workflows.
-The following infotable (Table 1) provides an overview of our COADVISE framework and functions provided by our Coadvise R package, including available variable selection methods, missing data imputation methods, and covariate-adjusted estimators, along with references to the relevant literature and the properties of different methods.
-As a remark, based on empirical evidence in Section 4, we recommend the use of Lasso or adaptive Lasso as the variable selection method in general, and we set Lasso as the default choice in our R package.
+Furthermore, due to concerns regarding the validity of post-selection inference arising from using the same dataset for both variable selection and covariate adjustment (Van Lancker et al., 2024; Belloni et al., 2017), in Section 3.3, we outline regularity conditions that variable selection procedures must satisfy to ensure valid post-selection inference, and we justify that the AIPW estimator remains valid when using Lasso and adaptive Lasso for variable selection.
+Second, as detailed in Section 3.4, we introduce two key extensions to the COADVISE framework beyond complete randomization.
+The first extension incorporates externally trained super-covariates—or outcome foundation models—into the variable selection process.
+The second extension adapts the framework to covariate-adaptive randomization, allowing treatment allocation to be stratified based on a factor-type variable to achieve improved balance across prognostic strata.
+All of the above contributions are implemented in our Coadvise R package.
+Unlike existing tools such as RobinCar, Coadvise explicitly supports variable selection, enabling users to identify covariates most predictive of outcomes and improve estimation efficiency.
+To clarify the distinctions between the two packages, we provide a detailed overview of the COADVISE framework and a comparison with RobinCar within the context of covariate adjustment for RCTs (see Table 1).
+For user convenience, the package includes several missing data imputation methods for handling missing covariates and outcomes, which are common in RCTs; details are provided in Appendix S.1.
+Our package also promotes interoperability by allowing users to export selected covariates for use in RobinCar or other tools and custom analyses.
+This flexibility ensures that Coadvise complements rather than replaces existing tools.
+Based on empirical evidence in Section 4, we recommend using Lasso or adaptive Lasso for variable selection and set Lasso as the default in our package.
 Finally, we apply the COADVISE framework to a real-world dataset from the BestAIR trial, using data from the National Sleep Research Resource (NSRR).
@@
 The open-source Coadvise package includes detailed documentation and reproducible examples, ensuring accessibility for a broad audience.
-Missing data imputation Handles missing covariates Handles missing outcome Assumption
-Complete-case analysis (cc) $\checkmark$ $\checkmark$ MCAR
-Multiple imputation by chained equations (mice; Van Buuren and Groothuis-Oudshoorn (2011)) $\checkmark$ $\checkmark$ MAR
-Random forest (missForest; Stekhoven and Bühlmann (2012)) $\checkmark$ $\checkmark$ MAR
-Inverse probability weighting (ipw; Robins et al. (1994)) $\times$ $\checkmark$ MAR
-Missingness indicator (missInd; Zhao et al. (2024); Zhao and Ding (2022); Song et al. (2021)) $\checkmark$ $\times$ MAR
-Variable selection method Lasso (Tibshirani, 1996), Adaptive Lasso (Zou, 2006), Marginal correlation (Permutt, 1990), Preliminary testing (Senn, 1994; Zhao and Ding, 2023)
+Variable selection Lasso (Tibshirani, 1996), Adaptive Lasso (Zou, 2006), Marginal correlation (Permutt, 1990), Preliminary testing (Senn, 1994; Zhao and Ding, 2023)
 Covariate-adjusted estimator Model-based variance estimator (see Section 2) Robust to model misspecification (see Section 2) Nonlinear adjustment (see Section 3) Efficiency gain guarantee (see Section 3)
@@
 AIPW (Bannick et al., 2023) Delta method $\checkmark$ $\checkmark$ $\checkmark$ if (i) or (ii) in Theorem 1 holds
-Software Multi-valued treatments Missing data imputation Variable selection Covariate-adjusted estimators
-Coadvise (ours) $\checkmark$ $\checkmark$ $\checkmark$ ANCOVA, ANHECOVA, AIPW
-RobinCar (Ye et al., 2023b) $\checkmark$ $\times$ $\times$ ANCOVA, ANHECOVA, AIPW
-  • MCAR: missing completely at random; MAR: missing at random.
+Software Multi-valued treatments Missing data imputation Variable selection Covariate-adaptive randomization
+Coadvise (ours) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+RobinCar (Ye et al., 2023b) $\checkmark$ $\times$ $\times$ $\checkmark$
   • $\dagger$: $\pi_{a}=P(A=a)$ is the probability of being assigned to the treatment group $a$, where $a=0,1$.
-Table 1: Overview of the COADVISE Framework: The top two sections include all available methods for imputing missing data and variable selection.
-The third section summarizes available covariate-adjusted estimators with associated variance estimators and their properties.
-The bottom section provides a comparison with the RobinCar package.
+Table 1: Overview of the COADVISE Framework.
+The top two sections summarize all available methods for variable selection and covariate-adjusted estimation, including key properties of the supported estimators.
+The bottom section compares COADVISE with the RobinCar package.
+Details on missing data imputation methods are provided in Table S.1 in Appendix S.1.
 2 Preliminaries
@@
 A key difference between these two frameworks lies in the source of randomness: in the finite-population framework, randomness arises from treatment assignment and outcomes, whereas in the super-population framework, additional randomness comes from the covariates.
-We consider a binary treatment, denoted by $A\in\{0,1\}$, where $1$ represents the treated group and 00 represents the control group.
+We consider a binary treatment, denoted by $A\in\{0,1\}$, where 111 represents the treated group and 0 represents the control group.
 Each participant in the sample is characterized by a baseline covariate vector $\mathbf{X}$, such as clinical characteristics.
@@
 However, we can only observe one outcome for each participant, denoted by $Y$.
-Given the nature and definition of potential outcomes, we assume that the observed outcome is consistent with the potential outcome under the received treatment, i.e., $Y=Y(A)=AY(1)+(1-A)Y(0)$.
-We further assume that there is only one version of the treatment $A$, and the potential outcome of each participant neither depends on nor affects the treatment received by others (also known as the stable-unit treatment value assumption [SUTVA]).
+Throughout, we assume that (i) the observed outcome is consistent with the potential outcome under the received treatment, i.e., $Y=AY(1)+(1-A)Y(0)$; and (ii) there is only one version of the treatment $A$, and the potential outcome of each participant neither depends on nor affects the treatment received by others (also known as the stable-unit treatment value assumption [SUTVA]).
 In addition, we make the following standard identification assumptions for RCT data.
@@
 Assumption 2 (Positivity).
-There exist two constants $c_{1}$ and $c_{2}$ such that $10<c_{1}<\pi_{1}=P(A=1)<c_{2}<$ 1 almost surely.
+There exist a constant $\eta$ such that $0<\eta\leq\pi_{a}=P(A=a)\leq 1-\eta<1$ almost surely, for $a=0,1$.
 Assumption 1 is reasonable because, in RCTs, the treatment assignment is controlled by the experimenters.
@@
 where $\to_{d}$ denotes convergence in distribution, and
-$\mathbf{V}=\begin{pmatrix}v_{11}&v_{12}\\ v_{21}&v_{22}\end{pmatrix},$ (7)
+$\mathbf{V}=\begin{pmatrix}v_{00}&v_{01}\\ v_{10}&v_{11}\end{pmatrix},$ (7)
 with
-$v_{aa}$ $=P(A=a)^{-1}\mathbb{V}\{Y_{a}-\mu_{a}(\mathbf{X})\}+2\text{Cov}\{Y_{a},\mu_{a}(\mathbf{X})\}-\mathbb{V}\{\mu_{a}(\mathbf{X})\},$
+$v_{aa}$ $=\pi_{a}^{-1}\mathbb{V}\{Y_{a}-\mu_{a}(\mathbf{X})\}+2\text{Cov}\{Y_{a},\mu_{a}(\mathbf{X})\}-\mathbb{V}\{\mu_{a}(\mathbf{X})\},\text{ where }\pi_{a}=P(A=a),$
 $v_{ab}$ =$\text{Cov}\{Y_{a},\mu_{b}(\mathbf{X})\}+\text{Cov}\{Y_{b},\mu_{a}(\mathbf{X})\}-\text{Cov}\{\mu_{a}(\mathbf{X}),\mu_{b}(\mathbf{X})\}\text{ for }a\not=b.$
@@
 However, $\mu_{a}$ is not necessarily equal to the true outcome model $\mathbb{E}\{Y(a)\mid\mathbf{X}\}$, unless the working model is correctly specified.
-The asymptotic variance for $\widehat{\tau}_{\text{AIPW}}=\widehat{\theta}_{1}-\widehat{\theta}_{0}=c^{\prime}\widehat{\boldsymbol{\theta}}:=f(\widehat{\boldsymbol{\theta}})$ is given by $\{\nabla f(\boldsymbol{\theta})\}^{\prime}\mathbf{V}\{\nabla f(\boldsymbol{\theta})\}=v_{11}-2v_{12}+v_{22}$, where $c=(1\ \ -1)^{\prime}$.
+The asymptotic variance for $\widehat{\tau}_{\text{AIPW}}=\widehat{\theta}_{1}-\widehat{\theta}_{0}=c^{\prime}\widehat{\boldsymbol{\theta}}:=f(\widehat{\boldsymbol{\theta}})$ is given by $\{\nabla f(\boldsymbol{\theta})\}^{\prime}\mathbf{V}\{\nabla f(\boldsymbol{\theta})\}=v_{00}-2v_{01}+v_{11}$, where $c=(1\ \ -1)^{\prime}$.
 Therefore, a consistent variance estimator is:
-$N^{-1}\{\nabla f(\widehat{\boldsymbol{\theta}})\}^{\prime}\widehat{\mathbf{V}}\{\nabla f(\widehat{\boldsymbol{\theta}})\}=N^{-1}\{\widehat{v}_{11}-2\widehat{v}_{12}+\widehat{v}_{22}\},$
-where $\widehat{v}_{11}$, $\widehat{v}_{12}$, and $\widehat{v}_{22}$ are sample-based estimates of $v_{11}$, $v_{12}$, and $v_{22}$, respectively, using the sample versions of the variances, covariances, and proportions.
+$N^{-1}\{\nabla f(\widehat{\boldsymbol{\theta}})\}^{\prime}\widehat{\mathbf{V}}\{\nabla f(\widehat{\boldsymbol{\theta}})\}=N^{-1}\{\widehat{v}_{00}-2\widehat{v}_{01}+\widehat{v}_{11}\},$
+where $\widehat{v}_{00}$, $\widehat{v}_{01}$, and $\widehat{v}_{11}$ are sample-based estimates of $v_{00}$, $v_{01}$, and $v_{11}$, respectively, using the sample versions of the variances, covariances, and proportions.
 3 Methodology
@@
 Once these models are estimated, each model is applied to the data from the opposite treatment group to predict the missing potential outcomes.
-In other words, a model $\widehat{\mu}_{a}(\boldsymbol{X})$ fitted using data from group $A=a$ is used to predict the potential outcome $Y(a)$ for subjects in group $A=1-a$, for $a=0,1$.
+In other words, a model $\widehat{\mu}_{a}(\mathbf{X})$ fitted using data from group $A=a$ is used to predict the potential outcome $Y(a)$ for subjects in group $A=1-a$, for $a=0,1$.
 This procedure yields all the information necessary to compute the AIPW estimator in (5).
@@
 When both outcome models are correctly specified, the AIPW estimator is the most efficient, achieving the asymptotic (semiparametric) variance lower bound (Hahn, 1998; Hirano et al., 2003).
+3.3 Post-selection inference
+Although Theorem 1 guarantees efficiency gain regardless of the choice of variable selection method, the uncertainty introduced by variable selection must be accounted for when conducting inference with the final AIPW estimator.
+This challenge is commonly referred to as the post-selection inference problem.
+In this section, we outline the theoretical conditions under which a variable selection method enables valid inference using the variance estimator applied to the AIPW estimator constructed from the selected covariates.
+Design-based theoretical results related to Lasso-based variable selection, in conjunction with the ANHECOVA estimator, have been studied in Bloniarz et al.
+(2016).
+Suppose we have $p>1$ covariates, denoted by $\mathbf{X}=(X_{1},\dots,X_{p})^{\prime}$.
+Let $\mathcal{S}_{0}\subset\{1,\dots,p\}$ and $\mathcal{S}_{1}\subset\{1,\dots,p\}$ denote the sets of indices corresponding to the relevant covariates for the potential outcomes $Y(0)$ and $Y(1)$, respectively.
+Define $s_{a}=|\mathcal{S}_{a}|$ for $a=0,1$.
+We draw upon the theoretical results of Farrell (2015) and Van Lancker et al.
+(2024) to summarize the following technical conditions within our framework.
+The former reference considers the more general setting of observational studies, whereas our RCT framework represents a special case that avoids conditions related to the propensity score model.
+We assume that the estimated outcome regression coefficient $\widehat{\boldsymbol{\beta}}_{a}$ converges in probability to a limiting value $\boldsymbol{\beta}_{a}^{*}$, which may differ from the true coefficient vector $\boldsymbol{\beta}_{a}$ in the outcome model $\mu_{a}(\mathbf{X})=\mu_{a}(\mathbf{X};\boldsymbol{\beta}_{a})$, as model misspecification is allowed in this setting.
+Condition 1.
+$s_{a}\cdot\log\{\max(p,N)\}/\sqrt{N}=o(1)$ for $a=0,1.$
+Condition 2 (Convergence and error rates of estimated coefficients).
+$G_{N}^{a}(\mathbf{X})=\frac{1}{N}\sum_{i=1}^{N}\{\mu_{a}(\mathbf{X}_{i};\widehat{\boldsymbol{\beta}}_{a})-\mu_{a}(\mathbf{X}_{i};\boldsymbol{\beta}_{a}^{*})\}^{2}$
+is uniformly integrable with $\sqrt{G_{N}^{a}(\mathbf{X})}=o_{P}(1)$, and $\|\widehat{\boldsymbol{\beta}}_{a}-\boldsymbol{\beta}_{a}^{*}\|_{1}=O_{P}\left(s_{a}\sqrt{\log\{\max(p,N)\}/N}\right)$ for $a=0,1$.
+Here, $P$ is a data generating process that obeys the following regularity and boundedness conditions (below, let $U_{a}=Y(a)-\mu_{a}(\mathbf{X})$ for $a=0,1$):
+  • $\{(A_{i},\mathbf{X}_{i},Y_{i})\}_{i=1}^{N}$ is an i.i.d. sample from population $\mathcal{O}=(A,\mathbf{X},Y)$;
+  • The covariates $\mathbf{X}$ have bounded support with $\max_{j\in\{1,\dots,p\}}|X_{j}|\leq M<\infty$;
+  • $\mathbb{E}\{|U_{a}|^{4}\mid\mathbf{X}\}\leq M<\infty$, $\min_{(j,a)\in\{1,\dots,p\}\times\{0,1\}}\mathbb{E}\{X_{j}^{2}U_{a}^{2}\}$ is bounded away from zero, and for some $r>0$, $\min\{\mathbb{E}\{|\mu_{1}(\mathbf{X})\mu_{0}(\mathbf{X})|^{1+r}\},\mathbb{E}\{|U_{0}|^{4+r}\},\mathbb{E}\{|U_{1}|^{4+r}\}\}\leq M<\infty$, for $a=0,1$.
+Condition 1 concerns the relationship among the sparsity $s_{a}$, the number of covariates $p$, and the sample size $N$.
+This condition is common in the high-dimensional inference literature (Belloni and Chernozhukov, 2013; Belloni et al., 2014, 2017; Bloniarz et al., 2016).
+Condition 2 requires that the fitted outcome models converge to a bounded limit, along with some regular and mild assumptions on the data-generating process.
+Both conditions may be strong in more general observational studies, but in our setting, where variable selection is employed and model misspecification is permitted, the dimensionality $p$ can be effectively limited and the existence of the limiting value $\boldsymbol{\beta}_{a}^{*}$ is reasonable.
+Outcome models by Lasso and adaptive Lasso variable selections with linear and logistic link functions have been shown to satisfy these conditions (Belloni and Chernozhukov, 2013; Belloni et al., 2017), supporting the validity of post-selection inference using the variance estimator adopted in our framework when either of these two variable selection methods is employed.
+Lasso also has a standardized implementation in R via the glmnet package, which can be fully pre-specified to prevent data dredging.
+For user convenience, we have incorporated this function for both Lasso and adaptive Lasso in our Coadvise R package.
+These methods also address the overfitting issues associated with OLS when the number of covariates is large relative to the sample size.
+Finally, we remark on the robustness of variance estimators for covariate-adjusted estimators.
+Thanks to the RCT design, the propensity score (treatment model) is known and correctly specified.
+In our framework, variable selection is applied only to the outcome models used in the covariate-adjusted estimators.
+For all estimators considered—ANCOVA, ANHECOVA, and AIPW with Lasso-based variable selection—Condition 1 ensures that the residuals in their HW sandwich or model-based variance estimators capture the unexplained variation, including uncertainty due to variable selection (Liu et al., 2023; Bloniarz et al., 2016; Zhu et al., 2024).
+As a result, the robust variance estimators constructed from these residuals implicitly account for this uncertainty.
+3.4 Extensions
+Thus far, we have discussed our framework under the setting of complete (simple) randomization.
+In this section, we outline two possible extensions of the COADVISE framework that address emerging challenges in RCT analysis and showcase how our framework may be combined with some advanced techniques.
+While these extensions are included to illustrate practical implementation possibilities, they are not central contributions of this paper and require more rigorous theoretical investigation in future work.
+We present them here to help motivate future research directions.
+3.4.1 Pre-specified super-covariate
+The first extension considers the recently emerging methodology of leveraging super-covariates to enhance RCT analyses.
+A super-covariate is typically defined or termed as a transformation, summary score, model output, or projection of the existing baseline covariates (potentially high-dimensional) that captures key prognostic information about the outcome, often learned from external data (can be randomized or observational).
+As noted in several studies (Gagnon-Bartsch et al., 2023; De Bartolomeis et al., 2025), the limited sample sizes in RCTs, together with the increasing availability of large external observational datasets containing the same covariates and outcomes, have motivated growing interest in constructing super-covariates from external data to improve the efficiency of ATE estimation.
+Prior research has investigated the super-covariate adjustment approaches.
+Gagnon-Bartsch et al.
+(2023) proposed applying machine learning algorithms on multiple external data to obtain foundation models.
+De Bartolomeis et al.
+(2025) introduced a hybrid AIPW estimator that integrates available foundation models to improve the efficiency of the AIPW estimator.
+Schuler et al.
+(2022); Højbjerre-Frandsen et al.
+(2025); Liao et al.
+(2025) and Sun (2024) considered leveraging historical data to learn prognostic regression models.
+Poulet et al.
+(2025) extended the idea of prediction powered inference to RCTs, providing externally trained prognostic scores as a function of baseline covariates for all RCT participants.
+However, a key concern is that the external data may differ from the RCT data due to distributional shifts, such as covariate shift or outcome shift (Liu et al., 2025; Lee et al., 2023, 2022).
+As a result, it is generally uncertain whether the super-covariates constructed from external data are truly predictive of the outcomes in the RCT population.
+To address this issue, we propose a general, data-driven strategy that incorporates both the super-covariates trained from external data and the original covariates in the RCT data for covariate adjustment.
+Let $\mathcal{O}=(\mathbf{X},A,Y)$ denote a copy of the RCT data, and in general, suppose we have $K\geq 1$ other mutually independent external data sources for super-covariates.
+The combination of the COADVISE framework and super-covariates is illustrated in the following Figure 2.
+We first obtain or train $K$ super-covariates (or models) from each of the $K$ external data sources, denoted by $f_{1}(\mathbf{X}),\dots,f_{K}(\mathbf{X})$.
+These super-covariates are then applied to the RCT participants to generate corresponding representations, e.g., by predictions.
+The resulting features are combined with the original covariates $\mathbf{X}$ to form an augmented dataset $\mathcal{O}^{\text{aug}}$, which is subsequently analyzed using the COADVISE framework.
+Both the original covariates and the super-covariates are treated as candidate variables for selection in the adjustment process.
+Figure 2: Illustration of incorporating super-covariates into the COADVISE framework in Figure 1.
+In practice, we recommend that users first obtain the $K$ super-covariates and apply them to the RCT data.
+They can then use our Coadvise R package to perform the analysis based on the augmented RCT dataset, as illustrated in Figure 2.
+Our proposal offers several flexible features.
+First, we allow each external dataset to construct super-covariates of arbitrary dimensions, which are then aggregated and incorporated into the augmented dataset as additional covariate information.
+Second, since COADVISE employs a data-driven variable selection strategy—such as Lasso—even if some of the $K$ super-covariates are not predictive of the outcome in the RCT data, they may be excluded prior to the final adjustment step.
+This enables COADVISE to automatically guard against the inclusion of irrelevant covariates or super-covariates and to maintain an appropriate balance between sample size and model complexity.
+3.4.2 Covariate-adaptive randomization
+The preceding discussion in this paper focused on complete (simple) randomization.
+However, more complex randomization schemes are commonly employed in practice to ensure balance in treatment allocation across key subgroups.
+For instance, stratified permuted block randomization often assigns participants within predefined blocks (e.g., based on age or clinical site) to treatment arms in a way that maintains a predetermined allocation ratio, thereby achieving better balance within each block (Zelen, 1974).
+Our second extension centers around the covariate-adaptive randomization (Bugni et al., 2018; Ye et al., 2022; Rafi, 2023; Bannick et al., 2023) including stratified permuted block randomization (Zelen, 1974) and Pocock and Simon’s minimization (Pocock and Simon, 1975; Taves, 1974) as special cases.
+In general, we let $\mathbf{Z}$ be a discrete baseline covariate vector used in covariate-adaptive randomization, where $\mathbf{Z}$ has $L<\infty$ distinct joint levels $\mathbf{z}_{1},\dots,\mathbf{z}_{L}$.
+Also, $\mathbf{Z}$ and $\mathbf{X}$ can share some common components.
+For example, the gender can be a factor for stratified randomization at the design stage, but it can also be incorporated into the covariate adjustment in the analysis stage.
+Following Bannick et al.
+(2023), we impose the following mild assumption on treatment assignment in covariate-adaptive RCTs.
+We use $\pi_{a}(\mathbf{z})=P(A=a\mid\mathbf{Z}=\mathbf{z})$ to denote the treatment assignment probability for arm $a$ within stratum $\mathbf{z}$.
+Assumption 3.
+[…]
+Assumption 3(ii) ensures that each treatment is assigned with the same marginal probability across all strata defined by $\mathbf{Z}$, which is a primary goal of covariate-adaptive randomization.
+The estimand of interest remains the ATE, $\tau=\mathbb{E}\{Y(1)-Y(0)\}$.
+As shown in Bannick et al.
+(2023), the AIPW estimator (5) remains consistent for $\tau$ even when the outcome model is misspecified.
+To incorporate covariate-adaptive randomization into the COADVISE framework, we propose using the same AIPW estimation procedure outlined in Figure 1 following variable selection.
+The only modification lies in the final step, where a generalized variance expression derived by Bannick et al.
+(2023) is applied.
+We first review key aspects of the asymptotic variance expression.
+Recalling the notation in Equation (6) from Section 2.4, we now present its modification under covariate-adaptive randomization, following Bannick et al.
+(2023), with stratification by $\mathbf{Z}$:
+$\sqrt{N}(\widehat{\boldsymbol{\theta}}-\boldsymbol{\theta})\to_{d}\mathcal{N}(\mathbf{0},\mathbf{V}^{*}),$ (8)
+where
+$\mathbf{V}^{*}=\mathbf{V}-\mathbb{E}[\mathbf{Q}(\mathbf{Z})\left\{\mathbf{\Omega}_{\text{SR}}-\mathbf{\Omega}(\mathbf{Z})\right\}\mathbf{Q}(\mathbf{Z})],$ (9)
+with
+$\mathbf{Q}(\mathbf{z})=\text{diag}\left\{\pi_{a}^{-1}\mathbb{E}\{Y(a)-\theta_{a}\mid\mathbf{Z}=\mathbf{z}\}-\pi_{a}^{-1}\mathbb{E}[\mu_{a}(\mathbf{X})-\mathbb{E}\{\mu_{a}(\mathbf{X})\}\mid\mathbf{Z}=\mathbf{z}]\right\},\quad\text{for }a=0,1.$
+Furthermore, $\mathbf{\Omega}(\mathbf{z})=\text{diag}\left\{\pi_{0}(\mathbf{z}),\pi_{1}(\mathbf{z})\right\}-\left(\pi_{0}(\mathbf{z})~~\pi_{1}(\mathbf{z})\right)^{\prime}\left(\pi_{0}(\mathbf{z})~~\pi_{1}(\mathbf{z})\right)$ is a $\times 2$ 2 covariance matrix determined by the covariate-adaptive randomization design.
+In contrast, $\mathbf{\Omega}_{\text{SR}}=\text{diag}\{\pi_{0},\pi_{1}\}-(\pi_{0}~~\pi_{1})^{\prime}(\pi_{0}~~\pi_{1})$ denotes the corresponding matrix under simple randomization.
+Under Assumption 3, the covariate-adaptive randomization mechanism assigns equal treatment probabilities within each stratum, implying $\pi_{a}(\mathbf{z})=\pi_{a}$ for all $\mathbf{z}$ and $a$, and thus $\mathbf{\Omega}(\mathbf{Z})=\mathbf{\Omega}_{\text{SR}}$.
+Consequently, $\mathbb{E}\left\{\mathbf{Q}(\mathbf{Z})(\mathbf{\Omega}_{\text{SR}}-\mathbf{\Omega}(\mathbf{Z}))\mathbf{Q}(\mathbf{Z})\right\}=\mathbf{0}$ and the asymptotic variance still simplifies to $\mathbf{V}^{*}=\mathbf{V}$.
+Nevertheless, to accommodate potential deviations in finite samples, we estimate and retain this correction term in our variance estimator.
+Specifically, we estimate the treatment probabilities by $\widehat{\pi}_{a}(\mathbf{z})=N_{a}(\mathbf{z})/N(\mathbf{z})$ and $\widehat{\pi}_{a}=N_{a}/N$ for $a=0,1$.
+Consistent with the estimator of $\mathbf{V}$ in Equation (7), we use sample-based estimates for all components, including proportions, variances, and covariances, to construct the finite-sample version of the variance estimator.
+Finally, by the Delta method, the variance estimator for the AIPW estimate of the ATE can be obtained accordingly, following the same derivation presented at the end of Section 2.4.
+Additionally, we include two basic estimators without covariate adjustment: the simple unadjusted estimator (1), which remains consistent under Assumption 3, and a stratified unadjusted estimator following Bugni et al.
+(2018), defined as
+$\widehat{\tau}_{\text{strata}}=\sum_{\mathbf{z}\in\mathcal{Z}}\frac{N(\mathbf{z})}{N}\left\{\bar{Y}(1;\mathbf{z})-\bar{Y}(0;\mathbf{z})\right\},$ (10)
+where $\mathbf{z}$ indexes the strata defined by $\mathbf{Z}$, $\mathcal{Z}$ denotes the support of $\mathbf{Z}$, $N(\mathbf{z})$ is the number of participants in stratum $\mathbf{z}$, and $\bar{Y}(a;\mathbf{z})$ is the average outcome in treatment arm $a$ within stratum $\mathbf{z}$.
+A natural plug-in variance estimator for $\widehat{\tau}_{\text{strata}}$ is given by:
+$\widehat{\mathbb{V}}\left(\widehat{\tau}_{\text{strata}}\right)=\sum_{\mathbf{z}\in\mathcal{Z}}\left(\frac{N(\mathbf{z})}{N}\right)^{2}\left\{\frac{\widehat{S}_{1}^{2}(\mathbf{z})}{N_{1}(\mathbf{z})}+\frac{\widehat{S}_{0}^{2}(\mathbf{z})}{N_{0}(\mathbf{z})}\right\},$ (11)
+where $N_{a}(\mathbf{z})$ is the number of participants receiving treatment $a$ in stratum $\mathbf{z}$, and $\widehat{S}_{a}^{2}(\mathbf{z})$ is the sample variance of outcomes in treatment arm $a$ within stratum $\mathbf{z}$, for $a=0,1$.
+We provide an R function, CoadviseCAR(), in our Coadvise R package to implement the proposed framework for covariate-adaptive randomized trials, where “CAR” stands for covariate-adaptive randomization.
+A toy example implementing the above three estimators using Lasso variable selection (for AIPW) is given in Appendix S.5.
 4 Numerical Experiments
@@
 We use a comprehensive data-generating process (DGP), detailed in Section 4.1, and generate two distinct types of datasets for each DGP.
-First, we generate 20 independent super-population full datasets, each containing $5\times 10^{7}$ independent units, with both potential outcomes $(Y(0),Y(1))$ available for each unit.
-We then compute the mean of $Y(1)-Y(0)$ across these 20 large datasets to obtain the true ATE.
-Given the large sample size and 20 replications, the associated uncertainty is negligible.
-Second, we generate $M=500$ independent random sample datasets of size $N$, with $N$ varying between 100 and 500 to represent different sample size scenarios.
+[…]
+Second, we generate $M=500$ independent random sample datasets of size $N$, with $N$ varying in {40, 100, 200, 500} to represent different sample size scenarios.
+The choice of 200 is is intended to mimic the sample size in the case study of Section 5, where $N=169$.
 In each Monte Carlo replication, the dataset consists of $N_{1}$ treated units and $N_{0}$ control units, with an expected 1:1 ratio such that $N_{0}+N_{1}=N$.
@@
 We then specify two models for the potential outcomes:
-Continuous outcome: $Y(a)=50+20\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}+a\delta(\mathbf{X})+\epsilon,$
-Binary outcome: $Y(a)\sim\text{Bern}(e(\mathbf{X},a)),\text{ where }e(\mathbf{X},a)=\{1+\exp(-20\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}-a[3+\delta(\mathbf{X})])\}^{-1},$
+Continuous outcome: $Y(a)=30+20\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}+a\delta(\mathbf{X})+\epsilon,$
+Binary outcome: $Y(a)\sim\text{Bern}(e(\mathbf{X},a)),\text{ where }$
+$e(\mathbf{X},a)=\{1+\exp(-20\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}-a\delta(\mathbf{X}))\}^{-1},$
 for $a=0,1$, where $\epsilon\sim\mathcal{N}(0,1)$ i.i.d., and $\delta(\mathbf{X})$ characterizes the individual treatment effect.
-The details of $\boldsymbol{\beta}_{0}$, different cases of $\delta(\mathbf{X})$ (linear and nonlinear), and the distributions used for $\mathbf{X}$ and $\mathbf{V}$ can be found in our Online Supplemental Material for reproducibility.
+The details of $\boldsymbol{\beta}_{0}$, different cases of $\delta(\mathbf{X})$ (linear and nonlinear), and the distributions used for $\mathbf{X}$ and $\mathbf{V}$ can be found in our Online Supplemental Material (Section B) for reproducibility.
 4.2 Methods for comparison
@@
 Method abbreviation Description
-All Include all $\mathbf{X}$ and $\mathbf{V}$, i.e., no covariate selection
-Lasso Select covariates with non-zero regression coefficient estimates in Lasso regression (Bloniarz et al., 2016)
-Adaptive Lasso Select covariates with non-zero regression coefficient estimates in adaptive Lasso regression (Zou, 2006)
-Corr ($k$) Select the $k$ covariates with the highest marginal correlations with $Y$, where $k=1,3,10$ (Permutt, 1990)
-Corr ($\xi$) Select covariates whose marginal correlations with $Y$ are greater than $\xi$, where $\xi=0.10,0.25$ (Permutt, 1990)
-Pre-test ($\alpha$) Perform a univariate two-sample $t$-test on each covariate to test for differences between two groups at level $\alpha$.
-Select covariates where the test is rejected, with $\alpha=0.05,0.10$ (Senn, 1994; Zhao and Ding, 2023)
+All
+Include all $\mathbf{X}$ and $\mathbf{V}$, i.e., no covariate selection
+Lasso
+Select covariates with non-zero regression
+coefficient estimates in Lasso regression (Bloniarz et al., 2016)
+Adaptive Lasso
+Select covariates with non-zero regression
+coefficient estimates in adaptive Lasso regression (Zou, 2006)
+Corr ($k$)
+Select the $k$ covariates with the highest
+marginal correlations with $Y$, where $k=1,3,10$ (Permutt, 1990)
+Corr ($\xi$)
+Select covariates whose marginal correlations
+with $Y$ are greater than $\xi$, where $\xi=0.10,0.25$ (Permutt, 1990)
+Pre-test ($\alpha$)
+Perform a univariate two-sample $t$-test on each covariate to test
+for differences between two groups at level $\alpha$. Select covariates
+where the test is rejected, with $\alpha=0.05,0.10$
+(Senn, 1994; Zhao and Ding, 2023)
 4.3 Performance criteria
@@
   • CP%: The proportion of replications where $\tau$ falls within the 95% confidence interval constructed from $\widehat{\tau}$ and its estimated standard error, using a normal approximation.
-  • Power%: The proportion of replications where the p-value is $\leq 0.05$ for testing $H_{0}:\tau=0$ versus $H_{a}:\tau\neq 0$.
+  • Power%: The proportion of replications where the p-value is 0.05$\leq$ 0.05≤ 0.05 for testing $H_{0}:\tau=0$ versus $H_{a}:\tau\neq 0$.
 With $M=500$ simulations, a CPis considered significantly different from the nominal 95% coverage level if it falls outside the range $95\%\pm 1.96\times\sqrt{95\%\times 5\%/500}=[93\%,97\%]$.
 4.4 Results
-For brevity, we present the simulation results for the linear continuous outcome under both sample sizes of 100 and 500 in this section.
-Complete simulation results, including figures, are available in Appendix B of our Online Supplemental Material.
-Figure 2: Simulation results for the linear continuous outcome with sample sizes (a) $N=100$ and (b) $N=500$.
+For brevity, we present the simulation results for the nonlinear continuous and nonlinear outcomes under sample sizes of 40 and 200 in this section.
+Complete simulation results, including figures, are available in our Online Supplemental Material (Section B).
+Nonlinear continuous outcome Refer to caption
+Nonlinear binary outcome Refer to caption
+Figure 3: Simulation results for the nonlinear and binary continuous outcomes with sample sizes (a) $N=40$ and (b) $N=200$.
 In the Bias boxplots, the black dots represent the medians.
@@
 Missing values in the CPand Powerfor ANHECOVA and ANCOVA are due to the non-obtainable sandwich variance estimators in high-dimensional settings.
-Figure 2 presents the results of biases via boxplots in the upper panels, and the results of CPand Powerin the lower four panels.
+Figure 4: Distributions of the AIPW estimates across 500 Monte Carlo replicates under nonlinear continuous and binary outcomes, with sample sizes $N=40$ and $N=200$.
+The red dashed lines represent the theoretical normal distributions based on the ATE truth and standard deviations of the 500 point estimates.
+The blue lines indicate the true ATE.
+Figure 3 presents the results of biases via boxplots in the upper panels, and the results of CPand Powerin the lower four panels under each outcome.
 We used different colors to distinguish between variable selection methods: green represents the estimator without variable selection (fully adjusted including all covariates), cyan represents Lasso and adaptive Lasso, gray indicates selection based on marginal correlation (as described in Table 2 for pre-specified $k$ or $\xi$), orange represents preliminary testing, and purple represents the simple estimator (which does not include any covariates).
-From Figure 2, we observe that when the sample size is small ($N=100$), there are many outliers in the boxplots for the ‘No selection” case for both AIPW and ANHECOVA estimators.
-This is expected, since the fully adjusted AIPW and ANHECOVA estimators include a relatively large number of covariates compared to the sample size.
-Across both sample sizes, the Simple, Pre-test (both $\alpha=0.05$ and $0.1$), and Corr ($k=1,3$) estimators exhibit wider bias box widths, with the Simple estimator widest, indicating that these methods are less efficient in this simulation setting.
-In terms of CPresults, we find that all methods are more stable around the 95% nominal level at the larger sample size, as the confidence intervals are Wald-type and rely on the large sample normal approximation.
-However, even when $N=500$, we observe below nominal CPfor both ANCOVA and ANHECOVA (due to the inability to obtain HW sandwich variance estimators in high-dimensional settings).
-Additionally, across both sample sizes, the Simple and Pre-test (both $\alpha=0.05$ and $0.1$) estimators result in the lowest power.
+From Figure 3, we observe that when the sample size is very small ($N=40$), there are many outliers in the boxplots for the “No selection” case for all covariate-adjusted estimators.
+This is expected, since they include a relatively large number of covariates compared to the sample size.
+Across both sample sizes, the Simple, Pre-test (both $\alpha=0.05$ and 0.10.10.1) estimators exhibit wider bias box widths, with the Simple estimator widest.
+The Corr method does not exhibit consistent performance across different parameter choices of $k$ and $\xi$.
+For example, Corr with $k=1$ or $k=3$ is notably less efficient when the sample size is $N=200$, yet performs relatively well under $N=40$.
+In contrast, Corr with $\xi=0.10$ shows poor performance under $N=40$ but improves substantially with a larger sample size of $N=200$.
+Lasso and adaptive Lasso exhibit similar performance patterns under the same estimator and are overall the most efficient and consistently performing methods across simulation settings.
+An exception is observed with the ANHECOVA estimator under $N=40$ for both outcomes, where some outliers appear.
+Nevertheless, in all other scenarios—particularly when paired with the AIPW estimator—these methods consistently demonstrate stable gains in efficiency and statistical power.
+These findings suggest that Simple, Pre-test, and Corr methods are generally less efficient in our simulation setting, particularly when compared to Lasso and adaptive Lasso.
+In terms of CPand power, all methods exhibit greater stability around the 95% nominal level at the larger sample size ($N=200$), reflecting improved consistency of the variance estimators with increasing sample size.
+However, even when $N=200$, we observe below nominal or missing CPfor all three covariate-adjusted estimators (due to the inability to obtain HW sandwich variance estimators in high-dimensional settings).
+Interestingly, across both sample sizes, the Simple and Pre-test (both $\alpha=0.05$ and 0.10.10.1) estimators result in the lowest power.
 This finding for the Pre-test estimator is consistent with Zhao and Ding (2023), where they advise against its use in practice based on theory and simulations.
-In contrast, Lasso and adaptive Lasso selection methods combined with AIPW generally perform well across both sample sizes and all three estimators in terms of efficiency, variance estimation (as measured by CP%), and power.
-Although we do not include the results of other DGPs, we briefly summarize them here.
-The results for the other continuous outcome are similar to those shown in Figure 2.
-For the two binary outcomes, we find that the efficiency gains observed in the boxplots are also similar to the continuous outcome cases, although the improvement relative to the Simple estimator is less pronounced due to the binary outcome being restricted to values in $\{0,1\}$.
-Moreover, the issue of non-obtainable sandwich variance estimators for ANHECOVA and ANCOVA is also present in the binary outcome cases.
+Lasso and adaptive Lasso, when combined with all estimators, yield satisfactory CPunder the continuous outcome with $N=200$.
+However, under the binary outcome setting with $N=200$, they exhibit under-coverage when used with the AIPW estimator, though this performance improves when the sample size increases to $N=500$, as shown in our Online Supplemental Material.
+This empirical finding suggests that, for binary outcomes using logistic regression as the outcome model in AIPW, it is valuable to investigate more robust variance estimation methods for smaller sample sizes in future research.
+Overall, Lasso and adaptive Lasso selection methods combined with AIPW generally perform well (in terms of both point and variance estimations) across both sample sizes and both outcomes.
+Futhermore, although we do not include the results of other DGPs, we briefly summarize them here.
+The results for other two outcomes (linear continuous and linear binary) are similar to those shown in Figure 3 and our comments above.
+At the largest sample size ($N=500$), the variance estimators following Lasso and adaptive Lasso selections perform more reliably with similar efficiency gains, yielding CPvalues that are closer to the 95% nominal level across all outcomes.
+However, the issue of non-obtainable sandwich variance estimators for ANHECOVA and ANCOVA without variable selection persists even at $N=500$.
+This indicates that, despite a seemingly adequate sample size relative to the data dimension, the sandwich variance estimator may remain computationally unstable or fail to converge when the number of covariates is large.
+This finding further highlights the advantage of applying variable selection to control the effective dimensionality, thereby facilitating the computation of model-based variance estimators.
+Finally, to empirically assess the asymptotic normality of post-selection estimators following Lasso and adaptive Lasso, we present histograms of AIPW estimates after variable selection in Figure 4.
+Additional results for ANCOVA, ANHECOVA, and other simulation scenarios are provided in the Online Supplemental Material.
+As shown in Figure 4, the empirical distributions closely align with the corresponding theoretical normal densities (indicated by red dashed lines).
+As sample size increases, the estimates become more concentrated around the mean, and the agreement with the theoretical curves improves under both continuous and binary outcomes.
+Supplemental figures further support these findings, demonstrating strong normality even at moderate sample sizes (e.g., $N=100$ and $N=500$).
+Taken together, these results provide empirical evidence for the asymptotic normality of covariate-adjusted AIPW estimators following variable selection via Lasso or adaptive Lasso.
 5 Case Study
@@
 The data includes 114 covariates, making it high-dimensional given the sample size ($N=169$).
+We consider two outcomes: (i) continuous SBP values (in mmHg); and (ii) binary SBP values, classifying SBP as high (130$\geq$ 130≥ 130 mmHg) or low (<130<130< 130 mmHg) based on a 130 mmHg threshold.
 Both the outcome and covariates contain missing values.
 Since only 25 observations are complete, we use the random forest method (Stekhoven and Bühlmann, 2012) (missForest option in our package) to impute all missing values.
-We consider two outcomes: (i) continuous SBP values (in mmHg); and (ii) binary SBP values, classifying SBP as high ($\geq 130$ mmHg) or low ($<130$ mmHg) based on a 130 mmHg threshold.
+To justify the missingness pattern, we note that missForest assumes an MAR mechanism, which, while untestable in practice, is less restrictive than the MCAR assumption.
+For missing outcomes, we include all baseline covariates (e.g., age, gender, race, study site) in the imputation model, making the MAR assumption more plausible.
+Since treatment is randomized and all covariates are collected prior to treatment assignment, the outcome missingness mechanism is unlikely to depend on unobserved outcomes.
+Furthermore, missingness in baseline covariates is less concerning in randomized trials, as these covariates are independent of both treatment and outcome by design.
+In this application, which primarily aims to demonstrate the COADVISE framework, we impute all missing covariates under the MAR assumption and include them in the variable selection step to take full advantage of the high-dimensional structure of the data.
 Table 3 presents the results of the analysis using three approaches to variable selection: (i) no covariate selection (all covariates included), (ii) Lasso, and (iii) adaptive Lasso.
@@
 We acknowledge several limitations of the COADVISE framework.
-One notable concern is the issue of post-selection inference resulting from variable selection prior to ATE estimation (Fithian et al., 2014).
-Although our simulations suggest that our approach yields valid inferences, further theoretical investigation is warranted.
+First, we need to further investigate the post-selection inference arising from variable selection methods beyond Lasso and adaptive Lasso.
+Although our simulations suggest that all variable selection methods included in our package yield valid inferences under larger sample sizes, additional theoretical work is warranted to support these findings.
 In support of our method, we note that our model-based variance estimators (such as the sandwich or Delta methods) are robust to model misspecification, ensuring valid uncertainty quantification regardless of the variables selected for the outcome model(s).
@@
 Future extensions might explore data integration and external information borrowing to further enhance the generalizability of RCT findings (Lee et al., 2022, 2023, 2024a, 2024b).
-Additionally, while our Coadvise R package includes a few commonly used missing data imputation methods for user convenience, we have not conducted a comprehensive investigation into their optimal integration within our framework, particularly in conjunction with variable selection.
-Future work could explore more robust approaches to handling missing data, including tailored imputation strategies that align with the underlying variable selection mechanisms to enhance both interpretability and statistical efficiency.
+Additionally, while our Coadvise R package includes several commonly used missing data imputation methods for user convenience, we have not undertaken a comprehensive investigation into their theoretical robustness or optimal integration within our framework, particularly in combination with variable selection.
+Future work could explore more robust imputation strategies that align with different variable selection mechanisms.
+We also plan to incorporate more advanced and recently developed imputation methods with theoretical guarantees that are compatible with our framework in future versions of the package.
 There are also several promising avenues for extending our framework.
@@
 Additionally, alternative strategies for estimating regression coefficients, such as debiased machine learning algorithms (Chernozhukov et al., 2018; Banik and Shao, 2022; Carlsson et al., 2014), debiased regressions (Lu et al., 2023; Chang et al., 2024; Chiang et al., 2023), and linear calibration (Bannick et al., 2023; Guo and Basse, 2023; Cohen and Fogarty, 2024) are of interest for future investigation.
-Data and software availability
-The BestAIR trial data used in Section 5 is available upon request at https://sleepdata.org.
+Data and code availability statement
+The BestAIR trial data used in Section 5 is available upon reasonable request at https://sleepdata.org.
 The R package Coadvise implementing the methodology of this paper is available at https://github.com/yiliu1998/Coadvise.
@@
 The authors are grateful to Marlena Bannick for insightful discussions during the early stages of this research endeavor.
-Liu was supported by the National Heart, Lung, and Blood Institute (NHLBI) of the National Institutes of Health (NIH) under Award Number T32HL079896.
-The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH.
+Yi Liu was supported by the National Heart, Lung, and Blood Institute (NHLBI) of the National Institutes of Health (NIH) under Award Number T32HL079896.
+This project is supported by the Food and Drug Administration (FDA) of the U.S.
+Department of Health and Human Services (HHS) as part of a financial assistance award U01FD007934 totaling \$1,674,013 over two years funded by FDA/HHS.
+It is also supported by the National Science Foundation under Award Number SES 2242776, totaling \$225,000 over three years.
+The contents are those of the authors and do not necessarily represent the official views of, nor an endorsement by, FDA/HHS, the NIH, or the U.S.
+Government.
 Appendix S Appendix: R Code and Tutorial to COADVISE
-S.1 R code for BestAIR trial study
+S.1 Summary of data imputation methods in COADVISE
+Imputation method
+Handles missing
+covariates
+Handles missing
+outcome
+Assumption
+Complete-case analysis (cc) $\checkmark$ $\checkmark$ MCAR
+Multiple imputation by chained equations
+(mice; Van Buuren and Groothuis-Oudshoorn (2011))
+$\checkmark$ $\checkmark$ MAR
+Random forest
+(missForest; Stekhoven and Bühlmann (2012))
+$\checkmark$ $\checkmark$ MAR
+Inverse probability weighting
+(ipw; Robins et al. (1994))
+$\times$ $\checkmark$ MAR
+Missingness indicator
+(missInd; Zhao et al. (2024);
+Zhao and Ding (2022); Song et al. (2021))
+$\checkmark$ $\times$ MAR
+Table S.1: Summary of missing data imputation methods available in the current Coadvise R package.
+S.2 R code for BestAIR trial study
 We show R code of the analysis for our case study in Section 5, as a tutorial for the use of our R package Coadvise.
@@
 In this example, we display components (i) and (iii) from the returned list.
-S.2 Algorithm example
+S.3 Algorithm example
 In this section, we present the proposed COADVISE framework, providing technical details in Algorithm 1. For illustration, we use Lasso for variable selection and AIPW for estimation.
-Let $\mathcal{R}^{p}$ denote the $p$-dimensional Euclidean space $(p>0)$.
-Define $N_{a}=\sum_{i=1}^{N}I(A_{i}=a)$ as the number of units with treatment assignment $A_{i}=a$, $s_{N}^{2}(\mathcal{W})$ as the sample variance of the i.i.d.
-data $\mathcal{W}=\{W_{1},\dots,W_{N}\}$, and $\text{cov}_{N}(\mathcal{W},\mathcal{V})$ as the sample covariance of the data $\mathcal{W}=\{W_{1},\dots,W_{N}\}$ and $\mathcal{V}=\{V_{1},\dots,V_{N}\}$.
-Similarly, $s_{N_{a}}^{2}(\mathcal{W})$ represents the sample variance for all $W_{i}\in\mathcal{W}$ with $A_{i}=a$, and $\text{cov}_{N_{a}}(\mathcal{W},\mathcal{V})$ denotes the sample covariance for all $W_{i}\in\mathcal{W}$ and $V_{i}\in\mathcal{V}$ with $A_{i}=a$, where $a=0,1$.
-Denote $\mathbb{P}_{N}[f(W)]=N^{-1}\sum_{i=1}^{N}f(W_{i})$ the empirical mean of any function of the i.i.d.
-observed data $\mathcal{W}$.
+[…]
 Algorithm 1 COADVISE by using Lasso for the variable selection and AIPW for ATE estimation
@@
 Return: $\widehat{\tau}=\widehat{\theta}_{1}-\widehat{\theta}_{0},\widehat{\mathbb{V}}(\tau)=N^{-1}(\widehat{v}_{11}-2\widehat{v}_{10}+\widehat{v}_{00}),\widehat{\mathcal{C}}_{0.95}(\tau)=\widehat{\tau}\pm 1.96\cdot\widehat{\mathbb{V}}(\tau).$
-S.3 Comparison of Coadvise and RobinCar
+S.4 Comparison of Coadvise and RobinCar
 In the code below, we use a toy example to compare and illustrate our Coadvise package and the well-known RobinCar package.
@@
 #### The ANCOVA estimator by RobinCar
-rc.ancova <- robincar_linear(df=df.RobinCar, treat_col="A", response_col="Y",
-covariate_cols=covar.names, contrast_h = "diff", adj_method="ANCOVA")
+rc.ancova <- robincar_linear(df=df.RobinCar, treat_col="A",
+response_col="Y", covariate_cols=covar.names,
+contrast_h="diff", adj_method="ANCOVA")
 print(rc.ancova)
@@
 #### The ANHECOVA estimator by RobinCar
-rc.anhecova <- robincar_linear(df=df.RobinCar, treat_col="A", response_col="Y",
-contrast_h = "diff", covariate_cols=covar.names, adj_method="ANHECOVA")
+rc.anhecova <- robincar_linear(df=df.RobinCar, treat_col="A",
+response_col="Y", contrast_h="diff",
+covariate_cols=covar.names,
+adj_method="ANHECOVA")
 print(rc.anhecova)
@@
 We would like to clarify that this occurs because for step2\$df.fit, we used the variables selected by ANCOVA for AIPW, whereas in step1\$df.fit, the variables selected in AIPW are different than ANCOVA, and the two outcome models in AIPW may use different sets of variables as predictors.
+S.5 A toy example for COADVISE with covariate-adaptive randomization
+In this section, we present a toy example to illustrate the use of the COADVISE framework for covariate-adaptive randomization via the CoadviseCAR() function.
+As detailed in Section 3.4, our Coadvise package includes three estimators for ATE estimation: the simple unadjusted estimator (Simple), the stratified unadjusted estimator (Strata), and the AIPW estimator.
+For AIPW, all variable selection methods supported in the Coadvise() function are available.
+Below, we provide the R code used to generate the toy example.
+We simulate data with $N=200$ observations and 100 covariates.
+Stratified randomization is performed across three strata—Low, Medium, and High—defined by a categorical variable.
+The potential outcomes are generated from a linear model involving the first three covariates, and the true ATE is set to $\tau=1$.
+#### Data generation
+n <- 200 # sample size
+p <- 100 # number of covariates
+strata_labels <- c("Low", "Medium", "High") # levels of strata
+X <- matrix(rnorm(n*p), nrow=n, ncol=p) # covariates
+colnames(X) <- paste0("X", 1:p)
+A <- numeric(n)
+# Generate treatment assignment with equal probability within each stratum
+for (z in strata_labels) {
+idx <- which(strata==z)
+A[idx] <- rbinom(length(idx), 1, 0.5)
+tau <- 1 # treatment effect (true ATE)
+base <- X[,1] - 0.5*X[,2] + 0.3*X[,3]
+Y0 <- base + rnorm(n) # potential outcome Y(0)
+Y1 <- base + tau + rnorm(n) # potential outcome Y(1)
+Y <- ifelse(A==1, Y1, Y0) # observed outcome
+#### Analysis using CoadviseCAR() function and results
+# Without variable selection for AIPW estimator
+CoadviseCAR(Y=Y, A=A,
+trt.name=1, ctrl.name=0, strata=strata,
+X=X, conti.out=TRUE, var.sel.method="No")\$df.fit
+method tau se ci.lwr ci.upr p
+1 Simple 1.078399 0.2258623 0.6357165 1.521081 1.800795e-06
+2 Strata 1.063202 0.2298618 0.6126817 1.513723 3.738767e-06
+3 AIPW 3.796362 NaN NaN NaN NaN
+# Lasso variable selection for AIPW estimator
+CoadviseCAR(Y=Y, A=A,
+trt.name=1, ctrl.name=0, strata=strata,
+X=X, conti.out=TRUE, var.sel.method="Lasso")\$df.fit
+method tau se ci.lwr ci.upr p
+1 Simple 1.078399 0.2258623 0.6357165 1.521081 1.800795e-06
+2 Strata 1.063202 0.2298618 0.6126817 1.513723 3.738767e-06
+3 AIPW 1.085008 0.1412289 0.8082046 1.361812 1.554312e-14
+From the analysis results, we observe that the Simple and Strata estimators yield similar point estimates and variance estimates.
+In contrast, the AIPW estimator without variable selection is highly biased and produces an unstable or unavailable variance estimate.
+However, after applying Lasso variable selection, the AIPW estimator provides an accurate point estimate with a smaller variance than both the Simple and Strata estimators, indicating a clear efficiency gain.
+Online Supplemental Material
 Appendix S Technical Proofs
-In this section, we present theoretical investigations into the efficiency gain achieved by the AIPW estimator for the ATE in RCT data.
+In this section, we present our theoretical investigations into the efficiency gain achieved by the AIPW estimator for the ATE in RCT data.
 We adhere to the notation, setups, and assumptions established in the main text of the paper.
@@
 Consequently, the estimator in the class $\mathcal{G}_{\boldsymbol{\beta}}^{g}$, which minimizes the variance, must be at least as efficient as $\widehat{\tau}_{\text{simple}}$.
-Furthermore, this result can easily be extended situations where a variable selection step is performed before AIPW estimation.
-Suppose we select $s$ variables from $\mathbf{X}$ for $Y(a)$ model, where $0<s<p$.
-When fitting the OLS estimator, the only difference is that $\boldsymbol{\beta}\in\mathcal{R}^{s}$ instead of $\mathcal{R}^{p}$ for estimators in $\mathcal{G}^{\boldsymbol{\beta}}_{g}$.
-Here, $\widehat{\tau}_{\text{simple}}$ remains a special case of estimators in $\mathcal{G}^{\boldsymbol{\beta}}_{g}$.
-Therefore, the proof remains straightforward.
+[…]
 Moreover, this illustrates that with more covariates (i.e., higher dimensions of $\boldsymbol{\beta}_{a}$), there is a greater opportunity to achieve higher efficiency.
@@
 $g_{\boldsymbol{\beta}_{a}\boldsymbol{\beta}_{a}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})$ $=\frac{\exp(-\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})\{\exp(-\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})-1\}}{\{1+\exp(-\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})\}^{3}}\mathbf{X}\mathbf{X}^{\prime}=g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})\{1-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})\}\{1-2g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})\}\mathbf{X}\mathbf{X}^{\prime}.$ (B.3.2)
-Therefore, for any general $\mathbf{X}\not=\boldsymbol{0}$, we have $g_{\boldsymbol{\beta}_{a}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})\not=\boldsymbol{0}$, since the logistic function satisfies $10<g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})<$ 1.
+Therefore, for any general $\mathbf{X}\not=\boldsymbol{0}$, we have $g_{\boldsymbol{\beta}_{a}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})\not=\boldsymbol{0}$, since the logistic function satisfies $0<g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})<1$.
 Additionally, $g_{\boldsymbol{\beta}_{a}\boldsymbol{\beta}_{a}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})=\boldsymbol{0}$ only when $\boldsymbol{\beta}_{a}^{*}=\boldsymbol{0}$ (so $g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})=1/2$); for other $\boldsymbol{\beta}_{a}^{*}$, the sign of $g_{\boldsymbol{\beta}_{a}\boldsymbol{\beta}_{a}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})$ can be either positive or negative.
@@
 This means that if the bias can be controlled by the right-hand side almost surely, along with (i), the proof in Section S.2 regarding the minimizer still applies to logistic regression.
-Figure S.3.1 illustrates the right-hand side of inequality in (ii).
-From the figure, if $\boldsymbol{\beta}_{a}^{*}$ is such that $g(\mathbf{X}^{\prime}\boldsymbol{\beta}^{*})\approx 1/2$, the (absolute value of) bias is allowed to be relatively large.
-However, when $g(\mathbf{X}^{\prime}\boldsymbol{\beta}^{*})$ is either smaller or larger, the bias must be smaller, approaching zero when $g(\mathbf{X}^{\prime}\boldsymbol{\beta}^{*})\approx 0$ or $1$.
-In other words, when the true proportion of $Y(a)$ is very small or vary large, the model must better approximate the true value.
-But when the true proportion is around $50\%$, more error in the posited model is acceptable.
-The intuition is when the model $g(\mathbf{X}^{\prime}\boldsymbol{\beta}^{*})$ approaches a medium value, the uncertainty regarding whether the true $Y(a)$ is 00 or $1$ increases.
+[…]
 Figure S.3.1: An illustration for the bound of absolute bias in Condition S.3.2(ii) (right-hand side of the inequality).
-Panel (a) provides the absolute biases over $g(\mathbf{X}^{\prime}\boldsymbol{\beta})\in[0.1,0.9]$, which ranges from $[0.11,\infty)$.
+Panel (a) provides the absolute biases over $g(\mathbf{X}^{\prime}\boldsymbol{\beta})\in[0.1,0.9]$, which ranges from [0.11,)[0.11$\infty$)[ 0.11 , ∞ ).
 When $g(\mathbf{X}^{\prime}\boldsymbol{\beta})=0.5$, the absolute bias is actually $\infty$; Panel (b) provides a further look for $g(\mathbf{X}^{\prime}\boldsymbol{\beta})\in[0.1,0.3]$.
@@
 For these reasons, we did not incorporate this approach into our framework.
-Appendix S Additional Simulation Results
+Appendix S Additional Simulation Details
 S.1 Additional details of the data generating process
@@
 $\boldsymbol{B}=\underbrace{\begin{pmatrix}0.10&0.10&\dots&0.10\\ 0.11&0.11&\dots&0.11\\ \vdots&\vdots&\ddots&\vdots\\ 0.59&0.59&\dots&0.59\end{pmatrix}}_{50\text{ columns}}+2\boldsymbol{I}_{50},$
-where $\boldsymbol{I}_{50}$ is the $50\times 50$ identity matrix.
+where $\boldsymbol{I}_{50}$ is the $\times 50$ 50 identity matrix.
 We specify the following two models for potential outcomes:
-Continuous outcome: $Y(a)=50+20\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}+a\delta(\mathbf{X})+\epsilon,$
-Binary outcome: $Y(a)\sim\text{Bern}(e(\mathbf{X},a)),\text{ with }e(\mathbf{X},a)=\{1+\exp(-20\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}-a[3+\delta(\mathbf{X})])\}^{-1},$
+Continuous outcome: $Y(a)=30+20\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}+a\delta(\mathbf{X})+\epsilon,$
+Binary outcome: $Y(a)\sim\text{Bern}(e(\mathbf{X},a)),\text{ with }e(\mathbf{X},a)=\{1+\exp(-20\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}-a\delta(\mathbf{X}))\}^{-1},$
 and we consider two specifications of $\delta(\mathbf{X})$:
-  • linear: $\delta(\mathbf{X})=c_{1}\mathbf{X}\boldsymbol{\beta}_{1}$, and
-  • nonlinear: $\delta(\mathbf{X})=c_{2}\mathbf{W}\boldsymbol{\beta}_{1}$, with $\mathbf{W}=(X_{1}^{2},\sin(X_{2}^{2}),X_{3}X_{4},X_{4}X_{5},X_{5}^{2})$,
-where $\boldsymbol{\beta}_{0}=(1,1,1,1,1)^{\prime}$, and $\boldsymbol{\beta}_{1}=(2,2,3,3,3)^{\prime}$.
-For continuous outcomes, $(c_{1},c_{2})=(8.15,1)$; for binary outcomes, $(c_{1},c_{2})=(10,1)$.
-The true ATEs, evaluating by the super-population data ($N=10^{8}$), are $\tau=8.15$ for both continuous outcomes and $\tau=0.11$ for both binary outcomes.
-S.2 Additional simulation results
+  • linear: $\delta(\mathbf{X})=d_{1}+c_{1}\mathbf{X}\boldsymbol{\beta}_{1}$, and
+  • nonlinear: $\delta(\mathbf{X})=d_{2}+c_{2}\mathbf{W}\boldsymbol{\beta}_{1}$, with $\mathbf{W}=(X_{1}^{2},-X_{2}^{2},|X_{3}|,X_{4}X_{5},X_{5})$,
+where $\boldsymbol{\beta}_{0}=(1,1,1,1,1)^{\prime}$, and $\boldsymbol{\beta}_{1}=(2,4,6,2,4)^{\prime}$.
+For continuous outcomes, $(c_{1},d_{1},c_{2},d_{2})=(1,8.15,2.92,0)$; for binary outcomes, $(c_{1},c_{2})=(10,2,2,2)$.
+The true ATEs, evaluating by the super-population data ($N=10^{8}$), are approximately $\tau=8.15$ for both continuous outcomes and $\tau=0.08$ for both binary outcomes.
+S.2 Additional simulation results of methods performance
+In Figures S.2.1S.2.16, we present full results (Bias, CPand Power%) of simulation in Section 4 of the main text.
+Figure S.2.1: Simulation results under continuous outcome, linear $\delta(\mathbf{X})$ and $N=40$.
+In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
+In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
 Figure S.2.2: Simulation results under continuous outcome, linear $\delta(\mathbf{X})$ and $N=100$.
+In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
+In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
+Figure S.2.3: Simulation results under continuous outcome, linear $\delta(\mathbf{X})$ and $N=200$.
 In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
@@
 In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
+Figure S.2.5: Simulation results under continuous outcome, nonlinear $\delta(\mathbf{X})$ and $N=40$.
+In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
+In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
 Figure S.2.6: Simulation results under continuous outcome, nonlinear $\delta(\mathbf{X})$ and $N=100$.
+In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
+In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
+Figure S.2.7: Simulation results under continuous outcome, nonlinear $\delta(\mathbf{X})$ and $N=200$.
 In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
@@
 In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
+Figure S.2.9: Simulation results under binary outcome, linear $\delta(\mathbf{X})$ and $N=40$.
+In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
+In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
 Figure S.2.10: Simulation results under binary outcome, linear $\delta(\mathbf{X})$ and $N=100$.
+In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
+In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
+Figure S.2.11: Simulation results under binary outcome, linear $\delta(\mathbf{X})$ and $N=200$.
 In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
@@
 In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
+Figure S.2.13: Simulation results under binary outcome, nonlinear $\delta(\mathbf{X})$ and $N=40$.
+In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
+In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
 Figure S.2.14: Simulation results under binary outcome, nonlinear $\delta(\mathbf{X})$ and $N=100$.
+In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
+In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
+Figure S.2.15: Simulation results under binary outcome, nonlinear $\delta(\mathbf{X})$ and $N=200$.
 In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
@@
 In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
+S.3 Additional simulation results of estimate distributions
+Figure S.3.1: Simulation results for estimation distributions by Lasso and adaptive Lasso variable selections, under continuous outcome, linear $\delta(\mathbf{X})$ and $N=40$.
+The red dashed curves are the theoretical normal densities curves, and the blue lines indicate the true ATE.
+Figure S.3.2: Simulation results for estimation distributions by Lasso and adaptive Lasso variable selections, under continuous outcome, linear $\delta(\mathbf{X})$ and $N=100$.
+The red dashed curves are the theoretical normal densities curves, and the blue lines indicate the true ATE.
+Figure S.3.3: Simulation results for estimation distributions by Lasso and adaptive Lasso variable selections, under continuous outcome, linear $\delta(\mathbf{X})$ and $N=200$.
+The red dashed curves are the theoretical normal densities curves, and the blue lines indicate the true ATE.
+Figure S.3.4: Simulation results for estimation distributions by Lasso and adaptive Lasso variable selections, under continuous outcome, linear $\delta(\mathbf{X})$ and $N=500$.
+The red dashed curves are the theoretical normal densities curves, and the blue lines indicate the true ATE.
+Figure S.3.5: Simulation results for estimation distributions by Lasso and adaptive Lasso variable selections, under continuous outcome, nonlinear $\delta(\mathbf{X})$ and $N=40$.
+The red dashed curves are the theoretical normal densities curves, and the blue lines indicate the true ATE.
+Figure S.3.6: Simulation results for estimation distributions by Lasso and adaptive Lasso variable selections, under continuous outcome, nonlinear $\delta(\mathbf{X})$ and $N=100$.
+The red dashed curves are the theoretical normal densities curves, and the blue lines indicate the true ATE.
+Figure S.3.7: Simulation results for estimation distributions by Lasso and adaptive Lasso variable selections, under continuous outcome, nonlinear $\delta(\mathbf{X})$ and $N=200$.
+The red dashed curves are the theoretical normal densities curves, and the blue lines indicate the true ATE.
+Figure S.3.8: Simulation results for estimation distributions by Lasso and adaptive Lasso variable selections, under continuous outcome, nonlinear $\delta(\mathbf{X})$ and $N=500$.
+The red dashed curves are the theoretical normal densities curves, and the blue lines indicate the true ATE.
+Figure S.3.9: Simulation results for estimation distributions by Lasso and adaptive Lasso variable selections, under binary outcome, linear $\delta(\mathbf{X})$ and $N=40$.
+The red dashed curves are the theoretical normal densities curves, and the blue lines indicate the true ATE.
+Figure S.3.10: Simulation results for estimation distributions by Lasso and adaptive Lasso variable selections, under binary outcome, linear $\delta(\mathbf{X})$ and $N=100$.
+The red dashed curves are the theoretical normal densities curves, and the blue lines indicate the true ATE.
+Figure S.3.11: Simulation results for estimation distributions by Lasso and adaptive Lasso variable selections, under binary outcome, linear $\delta(\mathbf{X})$ and $N=200$.
+The red dashed curves are the theoretical normal densities curves, and the blue lines indicate the true ATE.
+Figure S.3.12: Simulation results for estimation distributions by Lasso and adaptive Lasso variable selections, under binary outcome, linear $\delta(\mathbf{X})$ and $N=500$.
+The red dashed curves are the theoretical normal densities curves, and the blue lines indicate the true ATE.
+Figure S.3.13: Simulation results for estimation distributions by Lasso and adaptive Lasso variable selections, under binary outcome, nonlinear $\delta(\mathbf{X})$ and $N=40$.
+The red dashed curves are the theoretical normal densities curves, and the blue lines indicate the true ATE.
+Figure S.3.14: Simulation results for estimation distributions by Lasso and adaptive Lasso variable selections, under binary outcome, nonlinear $\delta(\mathbf{X})$ and $N=100$.
+The red dashed curves are the theoretical normal densities curves, and the blue lines indicate the true ATE.
+Figure S.3.15: Simulation results for estimation distributions by Lasso and adaptive Lasso variable selections, under binary outcome, nonlinear $\delta(\mathbf{X})$ and $N=200$.
+The red dashed curves are the theoretical normal densities curves, and the blue lines indicate the true ATE.
+Figure S.3.16: Simulation results for estimation distributions by Lasso and adaptive Lasso variable selections, under binary outcome, nonlinear $\delta(\mathbf{X})$ and $N=500$.
+The red dashed curves are the theoretical normal densities curves, and the blue lines indicate the true ATE.

### v5

@@
 The establishment of the American Statistical Association (ASA) Biopharmaceutical Section (BIOP) Covariate Adjustment Scientific Working Group (SWG) (https://carswg.github.io/) further underscores the growing focus on advancing statistical methods for covariate adjustment in clinical trials.
-[…]
+Despite these regulatory endorsements, significant challenges remain due to high-dimensional covariates and nonlinear relationships.
+These complexities necessitate robust and comprehensive analytical strategies.
+However, the comparative advantages and limitations of existing methods remain insufficiently explored, leaving practitioners with limited guidance and uncertainty in selecting the most appropriate approaches.
+For instance, in the case study presented in Section 5, the Best Apnea Interventions for Research (BestAIR) trial includes data from $N=169$ participants and $p=114$ covariates, making $p$ comparable to $N$.
+When adjusting for covariates using data from a single treatment arm—where $N_{1}=83$ (treated) and $N_{0}=86$ (controls)—$p$ exceeds the sample sizes.
+This high-dimensional setting can lead to biased and imprecise covariate-adjusted average treatment effect (ATE) estimates, as the sample size may be insufficient for reliable parameter estimation.
+Nevertheless, prior studies, such as Gao et al.
+(2024b), have identified baseline covariate imbalances in the BestAIR trial data, underscoring the need for efficient covariate-adjusted methods.
+This motivates our approach to improving estimation precision while addressing challenges aforementioned.
 This paper presents a unified framework for covariate adjustment in RCT analysis, with a primary focus on selecting important covariates for reducing data dimensionality for efficiency gains.
@@
 This estimator is not only unbiased but also asymptotically more efficient than the simple estimator under a wide range of conditions, without relying on restrictive model assumptions (Lin, 2013).
-The adoption of ANHECOVA has expanded in recent years, particularly for its ability to enhance efficiency in RCTs employing covariate-adaptive randomization schemes (Bannick et al., 2023).
+The adoption of ANHECOVA has expanded in recent years, particularly for its ability to enhance efficiency in RCTs employing covariate-adaptive randomization schemes (Bannick et al., 2025).
 The U.S.
@@
 (2024) provided a unified perspective on the relationships between the simple estimator, ANCOVA, and ANHECOVA, providing conditions under which ANHECOVA uniformly dominates other methods in terms of efficiency and precision.
-Beyond ANCOVA and ANHECOVA, a prominent method in the analysis of RCTs is augmented inverse probability weighting (AIPW) (Robins et al., 1994; Tsiatis, 2006; Glynn and Quinn, 2010; Cao et al., 2009; Kang et al., 2007).
+Beyond ANCOVA and ANHECOVA, a prominent method in the analysis of RCTs is augmented inverse probability weighting (AIPW) (Robins et al., 1994; Tsiatis, 2006; Glynn and Quinn, 2010; Cao et al., 2009; Kang et al., 2007; Yuan et al., 2012; Zhang et al., 2008; Tsiatis et al., 2008).
 The AIPW estimator combines outcome modeling with inverse treatment probability weighting, effectively leveraging information from both observed outcomes and treatment assignment probabilities.
@@
 As mentioned in Section 1.1, all covariate adjustment methods encounter difficulties in ensuring parameter convergence when faced with high-dimensional covariate data.
-As a result, a key preliminary step is variable selection—ideally choosing a set of variables smaller than the sample size $N$—that are most predictive of the outcome.
+As a result, a key preliminary step is variable selection—ideally choosing a set of variables smaller than the sample size NN—that are most predictive of the outcome.
 Cho and Yang (2023) and Belloni et al.
@@
 (2024) introduced a unified framework for variable selection combined with the AIPW estimator.
-Their approach also leverages double/debiased machine learning (DML) algorithms (Chernozhukov et al., 2018), employing techniques such as sample splitting and cross-fitting for model fitting and outcome prediction during covariate adjustment.
+Their approach also leverages double/debiased machine learning algorithms (Chernozhukov et al., 2018), employing techniques such as sample splitting and cross-fitting for model fitting and outcome prediction during covariate adjustment.
 This integration of machine learning with semiparametric methods represents a promising direction for addressing high-dimensional covariates while maintaining the robustness and efficiency of AIPW-based estimators.
@@
 ANHECOVA (Lin, 2013; Ye et al., 2023b) Sandwich $\checkmark$ $\times$ $\checkmark$
-AIPW (Bannick et al., 2023) Delta method $\checkmark$ $\checkmark$ $\checkmark$ if (i) or (ii) in Theorem 1 holds
+AIPW (Bannick et al., 2025) Delta method $\checkmark$ $\checkmark$ $\checkmark$ if (i) or (ii) in Theorem 1 holds
 Software Multi-valued treatments Missing data imputation Variable selection Covariate-adaptive randomization
@@
 A key difference between these two frameworks lies in the source of randomness: in the finite-population framework, randomness arises from treatment assignment and outcomes, whereas in the super-population framework, additional randomness comes from the covariates.
-We consider a binary treatment, denoted by $A\in\{0,1\}$, where 111 represents the treated group and 0 represents the control group.
+We consider a binary treatment, denoted by $A\in\{0,1\}$, where $1$ represents the treated group and 0 represents the control group.
 Each participant in the sample is characterized by a baseline covariate vector $\mathbf{X}$, such as clinical characteristics.
@@
 However, we can only observe one outcome for each participant, denoted by $Y$.
-Throughout, we assume that (i) the observed outcome is consistent with the potential outcome under the received treatment, i.e., $Y=AY(1)+(1-A)Y(0)$; and (ii) there is only one version of the treatment $A$, and the potential outcome of each participant neither depends on nor affects the treatment received by others (also known as the stable-unit treatment value assumption [SUTVA]).
+Throughout, we assume that (i) the observed outcome is consistent with the potential outcome under the received treatment, i.e., Y=AY(1)+(1A)Y(0)Y=AY(1)+(1-A)Y(0); and (ii) there is only one version of the treatment $A$, and the potential outcome of each participant neither depends on nor affects the treatment received by others (also known as the stable-unit treatment value assumption [SUTVA]).
 In addition, we make the following standard identification assumptions for RCT data.
@@
 2.4 Non-linear covariate adjustment
-The AIPW estimator for RCT data is given by (Bannick et al., 2023; Glynn and Quinn, 2010):
+The AIPW estimator for RCT data is given by (Bannick et al., 2025; Glynn and Quinn, 2010):
 $\widehat{\tau}_{\text{AIPW}}=\widehat{\theta}_{1}-\widehat{\theta}_{0},$ (5)
@@
 A robust variance estimator for the AIPW estimator can be derived using the Delta method (Taylor’s expansion) (Ye et al., 2023a).
-Under certain regularity conditions (Bannick et al., 2023) and denoting $\boldsymbol{\theta}=(\theta_{0},\theta_{1})^{\prime}$, it can be shown that:
+Under certain regularity conditions (Bannick et al., 2025) and denoting $\boldsymbol{\theta}=(\theta_{0},\theta_{1})^{\prime}$, it can be shown that:
 $\sqrt{N}(\widehat{\boldsymbol{\theta}}-\boldsymbol{\theta})\to_{d}\mathcal{N}(\boldsymbol{0},\mathbf{V}),$ (6)
@@
 However, $\mu_{a}$ is not necessarily equal to the true outcome model $\mathbb{E}\{Y(a)\mid\mathbf{X}\}$, unless the working model is correctly specified.
-The asymptotic variance for $\widehat{\tau}_{\text{AIPW}}=\widehat{\theta}_{1}-\widehat{\theta}_{0}=c^{\prime}\widehat{\boldsymbol{\theta}}:=f(\widehat{\boldsymbol{\theta}})$ is given by $\{\nabla f(\boldsymbol{\theta})\}^{\prime}\mathbf{V}\{\nabla f(\boldsymbol{\theta})\}=v_{00}-2v_{01}+v_{11}$, where $c=(1\ \ -1)^{\prime}$.
-Therefore, a consistent variance estimator is:
+[…]
 $N^{-1}\{\nabla f(\widehat{\boldsymbol{\theta}})\}^{\prime}\widehat{\mathbf{V}}\{\nabla f(\widehat{\boldsymbol{\theta}})\}=N^{-1}\{\widehat{v}_{00}-2\widehat{v}_{01}+\widehat{v}_{11}\},$
@@
 This approach allows the outcome models $\mu_{0}(\mathbf{X})$ and $\mu_{1}(\mathbf{X})$ to include different sets of variables, thereby providing greater flexibility in capturing the most relevant features for each potential outcome based on the data.
-Next, we fit two parametric outcome regression models—one for $Y(1)$ and one for $Y(0)$—using a chosen link function $g(\cdot)$ from the GLM family (for example, logistic regression for a binary outcome).
+Next, we fit two parametric outcome regression models—one for $Y(1)$ and one for Y(0)Y(0)—using a chosen link function $g(\cdot)$ from the GLM family (for example, logistic regression for a binary outcome).
 Once these models are estimated, each model is applied to the data from the opposite treatment group to predict the missing potential outcomes.
-In other words, a model $\widehat{\mu}_{a}(\mathbf{X})$ fitted using data from group $A=a$ is used to predict the potential outcome $Y(a)$ for subjects in group $A=1-a$, for $a=0,1$.
+In other words, a model $\widehat{\mu}_{a}(\mathbf{X})$ fitted using data from group $A=a$ is used to predict the potential outcome $Y(a)$ for subjects in group A=1aA=1-a, for $a=0,1$.
 This procedure yields all the information necessary to compute the AIPW estimator in (5).
@@
 Then, regardless of the variable selection method used, $\mathbb{V}_{\text{aipw}}\leq\mathbb{V}_{\text{simple}}$ in any of the following cases:
-  • (i) Both conditional outcome models $\mu_{0}(\mathbf{X})$ and $\mu_{1}(\mathbf{X})$, after variable selection, are specified by linear models with OLS used for estimating the regression coefficients.
-  • (ii) If the outcome model for $Y(a)$ ($a=0,1$), after variable selection, is specified by a nonlinear GLM with link function $g$, then there exists a $\boldsymbol{\beta}_{a}$ as the probability limit of MLE $\widehat{\boldsymbol{\beta}}_{a}$ such that with probability 1: (a) the sign of $g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})$ matches that of $\mathbb{E}(Y(a)\mid\mathbf{X})$; and (b) $|g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})|\leq 2|\mathbb{E}(Y(a)\mid\mathbf{X})|$.
+  1. (i)
+    Both conditional outcome models $\mu_{0}(\mathbf{X})$ and $\mu_{1}(\mathbf{X})$, after variable selection, are specified by linear models with OLS used for estimating the regression coefficients.
+  2. (ii)
+    If the outcome model for $Y(a)$ ($a=0,1$), after variable selection, is specified by a nonlinear GLM with link function $g$, then there exists a $\boldsymbol{\beta}_{a}$ as the probability limit of MLE $\widehat{\boldsymbol{\beta}}_{a}$ such that with probability 1:
+    1. (a)
+      The sign of $g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})$ matches that of $\mathbb{E}(Y(a)\mid\mathbf{X})$; and
+    2. (b)
+      $|g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})|\leq 2|\mathbb{E}(Y(a)\mid\mathbf{X})|$.
 For point (i) in Theorem 1, we highlight the connection with the ANHECOVA estimator.
@@
 When both outcome models are correctly specified, the AIPW estimator is the most efficient, achieving the asymptotic (semiparametric) variance lower bound (Hahn, 1998; Hirano et al., 2003).
+Remark 3 (Linear calibration for nonlinear outcome models).
+When outcome models are specified via nonlinear GLMs, Condition (ii) for efficiency gains in Theorem 1 may fail, because it requires the model-induced bias to be sufficiently small—a condition that is typically untestable in practice.
+While Condition (i) avoids this bias requirement by using linear regressions with OLS coefficient estimates, such an approach is not always optimal; for example, nonlinear GLMs are often more natural for binary or categorical outcomes.
+To address this limitation, we adopt a linear calibration step motivated by recent literature (Bannick et al., 2025; Cohen and Fogarty, 2024).
+Specifically, after obtaining the fitted outcome predictions $\widehat{\mu}_{1}(\mathbf{X})$ and $\widehat{\mu}_{0}(\mathbf{X})$ from nonlinear GLMs, we perform an additional calibration regression within each treatment arm.
+For treated participants $(A=1)$, we fit
+$\texttt{lm}(Y\sim\widehat{\mu}_{1}(\mathbf{X})+\widehat{\mu}_{0}(\mathbf{X})),$
+using the OLS estimates for regression coefficients, and we do the same for control participants $(A=0)$.
+Then, let $\widehat{\mu}_{0}^{\text{update}}(\mathbf{X})$ and $\widehat{\mu}_{1}^{\text{update}}(\mathbf{X})$ denote the predicted values for potential outcomes $Y(0)$ and $Y(1)$, respectively, from these regressions.
+We then replace the original $\widehat{\mu}_{a}(\mathbf{X})$ by $\widehat{\mu}_{a}^{\text{update}}(\mathbf{X})$ in the AIPW estimator $\widehat{\tau}_{\text{AIPW}}=\widehat{\theta}_{1}-\widehat{\theta}_{0}$, where for $a=0,1$, we update
+$\widehat{\theta}_{a}=\frac{1}{N_{a}}\sum_{i=1}^{N}I(A_{i}=a)[Y_{i}-\widehat{\mu}_{a}^{\text{update}}(\mathbf{X}_{i})]+\frac{1}{N}\sum_{i=1}^{N}\widehat{\mu}_{a}^{\text{update}}(\mathbf{X}_{i}),\text{ with }N_{a}=\sum_{i=1}^{N}I(A_{i}=a).$
+This correction aligns the predicted outcomes with the observed data within each treatment arm while retaining the flexibility of nonlinear GLMs, thereby ensuring that the AIPW estimator achieves efficiency gains without relying on the bias requirement in Theorem 1.
+We allow users to decide whether to apply linear calibration in practice by including the option lin.cal = TRUE or FALSE in the Coadvise() function of our Coadvise R package.
+When users are uncertain about whether their postulated nonlinear GLM models may introduce substantial bias, enabling this additional step is recommended.
 3.3 Post-selection inference
-Although Theorem 1 guarantees efficiency gain regardless of the choice of variable selection method, the uncertainty introduced by variable selection must be accounted for when conducting inference with the final AIPW estimator.
+Although Theorem 1 guarantees efficiency gain under established mild conditions on outcome models, regardless of the choice of variable selection method, the uncertainty introduced by variable selection must be accounted for when conducting inference with the final AIPW estimator.
 This challenge is commonly referred to as the post-selection inference problem.
@@
 For instance, stratified permuted block randomization often assigns participants within predefined blocks (e.g., based on age or clinical site) to treatment arms in a way that maintains a predetermined allocation ratio, thereby achieving better balance within each block (Zelen, 1974).
-Our second extension centers around the covariate-adaptive randomization (Bugni et al., 2018; Ye et al., 2022; Rafi, 2023; Bannick et al., 2023) including stratified permuted block randomization (Zelen, 1974) and Pocock and Simon’s minimization (Pocock and Simon, 1975; Taves, 1974) as special cases.
+Our second extension centers around the covariate-adaptive randomization (Bugni et al., 2018; Ye et al., 2022; Rafi, 2023; Bannick et al., 2025) including stratified permuted block randomization (Zelen, 1974) and Pocock and Simon’s minimization (Pocock and Simon, 1975; Taves, 1974) as special cases.
 In general, we let $\mathbf{Z}$ be a discrete baseline covariate vector used in covariate-adaptive randomization, where $\mathbf{Z}$ has $L<\infty$ distinct joint levels $\mathbf{z}_{1},\dots,\mathbf{z}_{L}$.
@@
 $\mathbf{Q}(\mathbf{z})=\text{diag}\left\{\pi_{a}^{-1}\mathbb{E}\{Y(a)-\theta_{a}\mid\mathbf{Z}=\mathbf{z}\}-\pi_{a}^{-1}\mathbb{E}[\mu_{a}(\mathbf{X})-\mathbb{E}\{\mu_{a}(\mathbf{X})\}\mid\mathbf{Z}=\mathbf{z}]\right\},\quad\text{for }a=0,1.$
-Furthermore, $\mathbf{\Omega}(\mathbf{z})=\text{diag}\left\{\pi_{0}(\mathbf{z}),\pi_{1}(\mathbf{z})\right\}-\left(\pi_{0}(\mathbf{z})~~\pi_{1}(\mathbf{z})\right)^{\prime}\left(\pi_{0}(\mathbf{z})~~\pi_{1}(\mathbf{z})\right)$ is a $\times 2$ 2 covariance matrix determined by the covariate-adaptive randomization design.
+Furthermore, $\mathbf{\Omega}(\mathbf{z})=\text{diag}\left\{\pi_{0}(\mathbf{z}),\pi_{1}(\mathbf{z})\right\}-\left(\pi_{0}(\mathbf{z})~~\pi_{1}(\mathbf{z})\right)^{\prime}\left(\pi_{0}(\mathbf{z})~~\pi_{1}(\mathbf{z})\right)$ is a $2\times 2$ covariance matrix determined by the covariate-adaptive randomization design.
 In contrast, $\mathbf{\Omega}_{\text{SR}}=\text{diag}\{\pi_{0},\pi_{1}\}-(\pi_{0}~~\pi_{1})^{\prime}(\pi_{0}~~\pi_{1})$ denotes the corresponding matrix under simple randomization.
@@
 We use a comprehensive data-generating process (DGP), detailed in Section 4.1, and generate two distinct types of datasets for each DGP.
-[…]
+First, we generate 20 independent super-population full datasets, each containing $5\times 10^{7}$ independent units, with both potential outcomes $(Y(0),Y(1))$ available for each unit.
+We then compute the mean of Y(1)Y(0)Y(1)-Y(0) across these 20 large datasets to obtain the true ATE.
+Given the large sample size and 20 replications, the associated uncertainty is negligible.
 Second, we generate $M=500$ independent random sample datasets of size $N$, with $N$ varying in {40, 100, 200, 500} to represent different sample size scenarios.
@@
   • CP%: The proportion of replications where $\tau$ falls within the 95% confidence interval constructed from $\widehat{\tau}$ and its estimated standard error, using a normal approximation.
-  • Power%: The proportion of replications where the p-value is 0.05$\leq$ 0.05≤ 0.05 for testing $H_{0}:\tau=0$ versus $H_{a}:\tau\neq 0$.
-With $M=500$ simulations, a CPis considered significantly different from the nominal 95% coverage level if it falls outside the range $95\%\pm 1.96\times\sqrt{95\%\times 5\%/500}=[93\%,97\%]$.
+  • Power%: The proportion of replications where the p-value is $\leq 0.05$ for testing $H_{0}:\tau=0$ versus $H_{a}:\tau\neq 0$.
+With $M=500$ simulations, a CP% is considered significantly different from the nominal 95% coverage level if it falls outside the range $95\pm 1.96\times\sqrt{95\times 5/500}=[93,97]$.
 4.4 Results
@@
 Nonlinear binary outcome Refer to caption
-Figure 3: Simulation results for the nonlinear and binary continuous outcomes with sample sizes (a) $N=40$ and (b) $N=200$.
-In the Bias boxplots, the black dots represent the medians.
-In the CPplots, the red dashed line indicates the 95% coverage level, while the gray dashed lines indicate 93% and 97% coverage levels.
-In the Powerplots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
-Missing values in the CPand Powerfor ANHECOVA and ANCOVA are due to the non-obtainable sandwich variance estimators in high-dimensional settings.
+Figure 3: Simulation results for nonlinear and binary outcomes with sample sizes (a) $N=40$ and (b) $N=200$.
+In the Bias boxplots, black dots indicate medians.
+In the CP% plots, the red dashed lines mark the 95% coverage level, while the gray dashed lines mark the 93% and 97% levels.
+In the Power% plots, the red dashed lines indicate 100% power, and the gray dashed lines indicate 80% power.
+Missing values in CP% and Power% for ANHECOVA and ANCOVA arise from the non-obtainable sandwich variance estimators in high-dimensional settings.
 Figure 4: Distributions of the AIPW estimates across 500 Monte Carlo replicates under nonlinear continuous and binary outcomes, with sample sizes $N=40$ and $N=200$.
@@
 The blue lines indicate the true ATE.
-Figure 3 presents the results of biases via boxplots in the upper panels, and the results of CPand Powerin the lower four panels under each outcome.
+Figure 3 presents the results of biases via boxplots in the upper panels, and the results of CP% and Power% in the lower four panels under each outcome.
 We used different colors to distinguish between variable selection methods: green represents the estimator without variable selection (fully adjusted including all covariates), cyan represents Lasso and adaptive Lasso, gray indicates selection based on marginal correlation (as described in Table 2 for pre-specified $k$ or $\xi$), orange represents preliminary testing, and purple represents the simple estimator (which does not include any covariates).
@@
 This is expected, since they include a relatively large number of covariates compared to the sample size.
-Across both sample sizes, the Simple, Pre-test (both $\alpha=0.05$ and 0.10.10.1) estimators exhibit wider bias box widths, with the Simple estimator widest.
+Across both sample sizes, the Simple, Pre-test (both $\alpha=0.05$ and $0.1$) estimators exhibit wider bias box widths, with the Simple estimator widest.
 The Corr method does not exhibit consistent performance across different parameter choices of $k$ and $\xi$.
@@
 These findings suggest that Simple, Pre-test, and Corr methods are generally less efficient in our simulation setting, particularly when compared to Lasso and adaptive Lasso.
-In terms of CPand power, all methods exhibit greater stability around the 95% nominal level at the larger sample size ($N=200$), reflecting improved consistency of the variance estimators with increasing sample size.
-However, even when $N=200$, we observe below nominal or missing CPfor all three covariate-adjusted estimators (due to the inability to obtain HW sandwich variance estimators in high-dimensional settings).
-Interestingly, across both sample sizes, the Simple and Pre-test (both $\alpha=0.05$ and 0.10.10.1) estimators result in the lowest power.
+In terms of CP% and power, all methods exhibit greater stability around the 95% nominal level at the larger sample size ($N=200$), reflecting improved consistency of the variance estimators with increasing sample size.
+However, even when $N=200$, we observe below nominal or missing CP% for all three covariate-adjusted estimators (due to the inability to obtain HW sandwich variance estimators in high-dimensional settings).
+Interestingly, across both sample sizes, the Simple and Pre-test (both $\alpha=0.05$ and $0.1$) estimators result in the lowest power.
 This finding for the Pre-test estimator is consistent with Zhao and Ding (2023), where they advise against its use in practice based on theory and simulations.
-Lasso and adaptive Lasso, when combined with all estimators, yield satisfactory CPunder the continuous outcome with $N=200$.
+Lasso and adaptive Lasso, when combined with all estimators, yield satisfactory CP% under the continuous outcome with $N=200$.
 However, under the binary outcome setting with $N=200$, they exhibit under-coverage when used with the AIPW estimator, though this performance improves when the sample size increases to $N=500$, as shown in our Online Supplemental Material.
@@
 The results for other two outcomes (linear continuous and linear binary) are similar to those shown in Figure 3 and our comments above.
-At the largest sample size ($N=500$), the variance estimators following Lasso and adaptive Lasso selections perform more reliably with similar efficiency gains, yielding CPvalues that are closer to the 95% nominal level across all outcomes.
+At the largest sample size ($N=500$), the variance estimators following Lasso and adaptive Lasso selections perform more reliably with similar efficiency gains, yielding CP% values that are closer to the 95% nominal level across all outcomes.
 However, the issue of non-obtainable sandwich variance estimators for ANHECOVA and ANCOVA without variable selection persists even at $N=500$.
@@
 The data includes 114 covariates, making it high-dimensional given the sample size ($N=169$).
-We consider two outcomes: (i) continuous SBP values (in mmHg); and (ii) binary SBP values, classifying SBP as high (130$\geq$ 130≥ 130 mmHg) or low (<130<130< 130 mmHg) based on a 130 mmHg threshold.
+We consider two outcomes: (i) continuous SBP values (in mmHg); and (ii) binary SBP values, classifying SBP as high ($\geq 130$ mmHg) or low (<130<130 mmHg) based on a 130 mmHg threshold.
 Both the outcome and covariates contain missing values.
@@
 Furthermore, our framework can be extended to a wide range of more general and complex settings.
-These include covariate-adaptive randomization (Bannick et al., 2023; Kernan et al., 1999; Broglio, 2018; Zhu et al., 2024; Yu et al., 2024), rerandomization (Zhu and Liu, 2023; Morgan and Rubin, 2012), time-to-event outcomes (Hua et al., 2024; Yi et al., 2020; Benkeser et al., 2021), restricted mean survival times (Han, 2023; Zhong et al., 2022; Karrison and Kocherginsky, 2018; Hanada et al., 2024), cluster-randomized trials (Murray, 1998; Donner et al., 2000; Stephens et al., 2012; Zhou et al., 2022; Balzer et al., 2023), and ordinal or composite outcomes with novel estimands such as the win ratio or generalized odds ratio (Agresti, 1980; Pocock et al., 2012; Barnhart et al., 2024; Matsouaka, 2022; Oakes, 2016; Zhang and Wu, 2024), as well as the Mann-Whitney rank-based statistic (Follmann et al., 2020).
-Other promising directions include conformal inference on individual or heterogeneous treatment effects (Yang et al., 2024; Wang et al., 2024; Gao et al., 2024a), the integration of multi-source RCT and observational data (Zhuang et al., 2024; Lee et al., 2024b, 2023; Yang et al., 2023; Han et al., 2025; Liu et al., 2024a; Yang and Wang, 2022; Liu et al., 2025), synthetic controls (Abadie et al., 2015, 2010; Ben-Michael et al., 2021), and external controls (Li et al., 2023; Lee et al., 2024c; Gao et al., 2023; Wang et al., 2025).
-Additionally, alternative strategies for estimating regression coefficients, such as debiased machine learning algorithms (Chernozhukov et al., 2018; Banik and Shao, 2022; Carlsson et al., 2014), debiased regressions (Lu et al., 2023; Chang et al., 2024; Chiang et al., 2023), and linear calibration (Bannick et al., 2023; Guo and Basse, 2023; Cohen and Fogarty, 2024) are of interest for future investigation.
+These include more general covariate-adaptive randomization (Bannick et al., 2025; Kernan et al., 1999; Broglio, 2018; Zhu et al., 2024; Yu et al., 2024), rerandomization (Zhu and Liu, 2023; Morgan and Rubin, 2012), time-to-event outcomes (Hua et al., 2024; Yi et al., 2020; Benkeser et al., 2021), restricted mean survival times (Han, 2023; Zhong et al., 2022; Karrison and Kocherginsky, 2018; Hanada et al., 2024), cluster-randomized trials (Murray, 1998; Donner et al., 2000; Stephens et al., 2012; Zhou et al., 2022; Balzer et al., 2023), and ordinal or composite outcomes with novel estimands such as the win ratio or generalized odds ratio (Agresti, 1980; Pocock et al., 2012; Barnhart et al., 2024; Matsouaka, 2022; Oakes, 2016; Zhang and Wu, 2024), as well as the Mann-Whitney rank-based statistic (Follmann et al., 2020).
+Other promising directions include conformal inference on individual or heterogeneous treatment effects (Yang et al., 2024; Wang et al., 2024; Gao et al., 2024a), the integration of multi-source RCT and observational data (Zhuang et al., 2024; Lee et al., 2024b, 2023; Yang et al., 2023; Han et al., 2025; Liu et al., 2024a; Yang and Wang, 2022; Liu et al., 2025), synthetic controls (Abadie et al., 2015, 2010; Ben-Michael et al., 2021), and external controls (Li et al., 2023; Lee et al., 2024c; Gao et al., 2023; Wang et al., 2025a).
+Additionally, alternative strategies for estimating regression coefficients, such as debiased machine learning algorithms (Chernozhukov et al., 2018; Banik and Shao, 2022; Wang et al., 2025b; Carlsson et al., 2014), debiased regressions (Lu et al., 2023; Chang et al., 2024; Chiang et al., 2023), and linear calibration (Bannick et al., 2025; Guo and Basse, 2023; Cohen and Fogarty, 2024) are of interest for future investigation.
 Data and code availability statement
@@
 Appendix S Appendix: R Code and Tutorial to COADVISE
-S.1 Summary of data imputation methods in COADVISE
-Imputation method
-Handles missing
-covariates
-Handles missing
-outcome
-Assumption
+S.1 Summary of missing data imputation methods in COADVISE
+Imputation method Handles missing covariates Handles missing outcome Assumption
 Complete-case analysis (cc) $\checkmark$ $\checkmark$ MCAR
-Multiple imputation by chained equations
-(mice; Van Buuren and Groothuis-Oudshoorn (2011))
-$\checkmark$ $\checkmark$ MAR
-Random forest
-(missForest; Stekhoven and Bühlmann (2012))
-$\checkmark$ $\checkmark$ MAR
-Inverse probability weighting
-(ipw; Robins et al. (1994))
-$\times$ $\checkmark$ MAR
-Missingness indicator
-(missInd; Zhao et al. (2024);
-Zhao and Ding (2022); Song et al. (2021))
-$\checkmark$ $\times$ MAR
+Multiple imputation by chained equations (mice; Van Buuren and Groothuis-Oudshoorn (2011)) $\checkmark$ $\checkmark$ MAR
+Random forest (missForest; Stekhoven and Bühlmann (2012)) $\checkmark$ $\checkmark$ MAR
+Inverse probability weighting (ipw; Robins et al. (1994)) $\times$ $\checkmark$ MAR
+Missingness indicator (missInd; Zhao et al. (2024); Zhao and Ding (2022); Song et al. (2021)) $\checkmark$ $\times$ MAR
 Table S.1: Summary of missing data imputation methods available in the current Coadvise R package.
@@
 out0.model.aipw $\mathbb{E}\{Y(0)\mid\mathbf{X}\}$ model in AIPW $\times$ "linear" "linear", "logit", "probit" "log", "cloglog", "identity"
+lin.cal Whether to use linear calibration in AIPW $\times$ FALSE TRUE, FALSE
 k Number of covariates with highest correlations with outcome $\times$ 1 Any positive integers less than the number of covariates
@@
 In this section, we present the proposed COADVISE framework, providing technical details in Algorithm 1. For illustration, we use Lasso for variable selection and AIPW for estimation.
-[…]
+Let $\mathcal{R}^{p}$ denote the $p$-dimensional Euclidean space $(p>0)$.
+Define $N_{a}=\sum_{i=1}^{N}I(A_{i}=a)$ as the number of units with treatment assignment $A_{i}=a$, $s_{N}^{2}(\mathcal{W})$ as the sample variance of the i.i.d.
+data $\mathcal{W}=\{W_{1},\dots,W_{N}\}$, and $\text{cov}_{N}(\mathcal{W},\mathcal{V})$ as the sample covariance of the data $\mathcal{W}=\{W_{1},\dots,W_{N}\}$ and $\mathcal{V}=\{V_{1},\dots,V_{N}\}$.
+Similarly, $s_{N_{a}}^{2}(\mathcal{W})$ represents the sample variance for all $W_{i}\in\mathcal{W}$ with $A_{i}=a$, and $\text{cov}_{N_{a}}(\mathcal{W},\mathcal{V})$ denotes the sample covariance for all $W_{i}\in\mathcal{W}$ and $V_{i}\in\mathcal{V}$ with $A_{i}=a$, where $a=0,1$.
+Denote $\mathbb{P}_{N}[f(W)]=N^{-1}\sum_{i=1}^{N}f(W_{i})$ the empirical mean of any function of the i.i.d.
+observed data $\mathcal{W}$.
 Algorithm 1 COADVISE by using Lasso for the variable selection and AIPW for ATE estimation
@@
 ?robincar_linear # the function in RobinCar we will use later
-#### We first use Coadvise to perform a variable selection, using the Corr.k method with k=5, and results are saved in the following "step1" object
+#### We first use Coadvise to perform variable selection, using the Corr.k method with k=5, and results are saved in the "step1" object
 step1 <- Coadvise(Y=Y, X=X, A=A, trt.name=1, ctrl.name=0,
@@
 seed=1335)
-#### We then use covariates selected in step1 for ANCOVA for further analyses
+#### We then use covariates selected in step1 for ANCOVA
 X.selected <- step1\$stage1.covars\$ANCOVA
@@
 #### The simple estimator by RobinCar
-rc.simple <- robincar_linear(df=df.RobinCar, treat_col="A", response_col="Y",
+rc.simple <- robincar_linear(df=df.RobinCar, treat_col="A",
+response_col="Y",
 contrast_h="diff", adj_method="ANOVA")
@@
 Then, it is natural to consider setting the derivative with respect to $({\boldsymbol{\beta}_{0}^{*}}^{\prime},{\boldsymbol{\beta}_{1}^{*}}^{\prime})^{\prime}$ to 0, i.e., to solve equations
-0 0 $=\frac{1-\pi_{1}}{\pi_{1}}\mathbb{E}\left\{g_{\boldsymbol{\beta}_{1}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{1}^{*})[Y(1)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{1}^{*})]\right\},$
-0 0 $=\frac{\pi_{1}}{1-\pi_{1}}\mathbb{E}\left\{g_{\boldsymbol{\beta}_{0}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}^{*})[Y(0)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}^{*})]\right\},$ (B.1.4)
+$0$ $=\frac{1-\pi_{1}}{\pi_{1}}\mathbb{E}\left\{g_{\boldsymbol{\beta}_{1}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{1}^{*})[Y(1)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{1}^{*})]\right\},$
+$0$ $=\frac{\pi_{1}}{1-\pi_{1}}\mathbb{E}\left\{g_{\boldsymbol{\beta}_{0}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}^{*})[Y(0)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}^{*})]\right\},$ (B.1.4)
 where $g_{\boldsymbol{\beta}}(\mathbf{X}^{\prime}\boldsymbol{\beta}^{*})$ denotes $\dfrac{\partial}{\partial\boldsymbol{\beta}}g(\mathbf{X}^{\prime}\boldsymbol{\beta})\bigg|_{\boldsymbol{\beta}=\boldsymbol{\beta}^{*}}$.
@@
 Consequently, the estimator in the class $\mathcal{G}_{\boldsymbol{\beta}}^{g}$, which minimizes the variance, must be at least as efficient as $\widehat{\tau}_{\text{simple}}$.
-[…]
+Furthermore, this result can easily be extended situations where a variable selection step is performed before AIPW estimation.
+Suppose we select $s$ variables from $\mathbf{X}$ for $Y(a)$ model, where $0<s<p$.
+When fitting the OLS estimator, the only difference is that $\boldsymbol{\beta}\in\mathcal{R}^{s}$ instead of $\mathcal{R}^{p}$ for estimators in $\mathcal{G}^{\boldsymbol{\beta}}_{g}$.
+Here, $\widehat{\tau}_{\text{simple}}$ remains a special case of estimators in $\mathcal{G}^{\boldsymbol{\beta}}_{g}$.
+Therefore, the proof remains straightforward.
 Moreover, this illustrates that with more covariates (i.e., higher dimensions of $\boldsymbol{\beta}_{a}$), there is a greater opportunity to achieve higher efficiency.
@@
 To further explore the conditions for efficiency gain, we examine the difference between the asymptotic variances of $\widehat{\tau}_{\text{simple}}$ and $\widehat{\tau}^{g}$, given by:
-$\mathbb{V}(\widehat{\tau}_{\text{simple}})-\mathbb{V}(\widehat{\tau}^{g})$ $=\mathbb{E}\left\{\frac{1-\pi_{1}}{\pi_{1}}Y(1)^{2}+\frac{\pi_{1}}{1-\pi_{1}}Y(0)^{2}\right\}-\mathbb{E}\left\{\frac{1-\pi_{1}}{\pi_{1}}\{Y(1)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{1}^{*})\}^{2}+\frac{\pi_{1}}{1-\pi_{1}}\{Y(0)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}^{*})\}^{2}\right\}$
+$\mathbb{V}(\widehat{\tau}_{\text{simple}})-\mathbb{V}(\widehat{\tau}^{g})$ $=\mathbb{E}\left\{\frac{1-\pi_{1}}{\pi_{1}}Y(1)^{2}+\frac{\pi_{1}}{1-\pi_{1}}Y(0)^{2}\right\}$
+$\qquad-\mathbb{E}\left\{\frac{1-\pi_{1}}{\pi_{1}}\{Y(1)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{1}^{*})\}^{2}+\frac{\pi_{1}}{1-\pi_{1}}\{Y(0)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}^{*})\}^{2}\right\}$
 $=\mathbb{E}\left\{\frac{1-\pi_{1}}{\pi_{1}}g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{1})\{2Y(1)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{1})\}+\frac{\pi_{1}}{1-\pi_{1}}g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{0})\{2Y(0)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{0})\}\right\}$
 $=\frac{1-\pi_{1}}{\pi_{1}}\underbrace{\mathbb{E}\left\{g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{1})[2Y(1)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{1})]\right\}}_{I_{1}}+\frac{\pi_{1}}{1-\pi_{1}}\underbrace{\mathbb{E}\left\{g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{0})[2Y(0)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{0})]\right\}}_{I_{0}}.$ (B.3.1)
-For both $a=0,1$, we note that $I_{a}=\mathbb{E}\left\{g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})[2Y(a)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})]\right\}=\mathbb{E}\left\{g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})[2\mathbb{E}(Y(a)\mid\mathbf{X})-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})]\right\}=\mathbb{E}\{I_{a}(\mathbf{X})\}$, where $I_{a}(\mathbf{X})=g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})[2\mathbb{E}(Y(a)\mid\mathbf{X})-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})]$.
+For $a=0,1$, we note that $I_{a}=\mathbb{E}\left\{g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})[2Y(a)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})]\right\}=\mathbb{E}\left\{g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})[2\mathbb{E}(Y(a)\mid\mathbf{X})-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})]\right\}=\mathbb{E}\{I_{a}(\mathbf{X})\}$, where $I_{a}(\mathbf{X})=g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})[2\mathbb{E}(Y(a)\mid\mathbf{X})-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})]$.
 Therefore, $I_{a}(\mathbf{X})$ achieves its maximum when $g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})=\mathbb{E}(Y(a)\mid\mathbf{X})$.
@@
 This means that if the bias can be controlled by the right-hand side almost surely, along with (i), the proof in Section S.2 regarding the minimizer still applies to logistic regression.
-[…]
+Figure S.3.1 illustrates the right-hand side of inequality in (ii).
+From the figure, if $\boldsymbol{\beta}_{a}^{*}$ is such that $g(\mathbf{X}^{\prime}\boldsymbol{\beta}^{*})\approx 1/2$, the (absolute value of) bias is allowed to be relatively large.
+However, when $g(\mathbf{X}^{\prime}\boldsymbol{\beta}^{*})$ is either smaller or larger, the bias must be smaller, approaching zero when $g(\mathbf{X}^{\prime}\boldsymbol{\beta}^{*})\approx 0$ or $1$.
+In other words, when the true proportion of $Y(a)$ is very small or vary large, the model must better approximate the true value.
+But when the true proportion is around $50\%$, more error in the posited model is acceptable.
+The intuition is when the model $g(\mathbf{X}^{\prime}\boldsymbol{\beta}^{*})$ approaches a medium value, the uncertainty regarding whether the true $Y(a)$ is 0 or $1$ increases.
 Figure S.3.1: An illustration for the bound of absolute bias in Condition S.3.2(ii) (right-hand side of the inequality).
-Panel (a) provides the absolute biases over $g(\mathbf{X}^{\prime}\boldsymbol{\beta})\in[0.1,0.9]$, which ranges from [0.11,)[0.11$\infty$)[ 0.11 , ∞ ).
+Panel (a) provides the absolute biases over $g(\mathbf{X}^{\prime}\boldsymbol{\beta})\in[0.1,0.9]$, which ranges from $0.11,\infty)$.
 When $g(\mathbf{X}^{\prime}\boldsymbol{\beta})=0.5$, the absolute bias is actually $\infty$; Panel (b) provides a further look for $g(\mathbf{X}^{\prime}\boldsymbol{\beta})\in[0.1,0.3]$.
@@
 $\boldsymbol{B}=\underbrace{\begin{pmatrix}0.10&0.10&\dots&0.10\\ 0.11&0.11&\dots&0.11\\ \vdots&\vdots&\ddots&\vdots\\ 0.59&0.59&\dots&0.59\end{pmatrix}}_{50\text{ columns}}+2\boldsymbol{I}_{50},$
-where $\boldsymbol{I}_{50}$ is the $\times 50$ 50 identity matrix.
+where $\boldsymbol{I}_{50}$ is the $50\times 50$ identity matrix.
 We specify the following two models for potential outcomes:
@@
 S.2 Additional simulation results of methods performance
-In Figures S.2.1S.2.16, we present full results (Bias, CPand Power%) of simulation in Section 4 of the main text.
+In Figures S.2.1S.2.16, we present full results (Bias, CP% and Power%) of simulation in Section 4 of the main text.
 Figure S.2.1: Simulation results under continuous outcome, linear $\delta(\mathbf{X})$ and $N=40$.
-In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
+In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
 In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
 Figure S.2.2: Simulation results under continuous outcome, linear $\delta(\mathbf{X})$ and $N=100$.
-In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
+In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
 In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
 Figure S.2.3: Simulation results under continuous outcome, linear $\delta(\mathbf{X})$ and $N=200$.
-In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
+In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
 In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
 Figure S.2.4: Simulation results under continuous outcome, linear $\delta(\mathbf{X})$ and $N=500$.
-In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
+In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
 In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
 Figure S.2.5: Simulation results under continuous outcome, nonlinear $\delta(\mathbf{X})$ and $N=40$.
-In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
+In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
 In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
 Figure S.2.6: Simulation results under continuous outcome, nonlinear $\delta(\mathbf{X})$ and $N=100$.
-In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
+In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
 In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
 Figure S.2.7: Simulation results under continuous outcome, nonlinear $\delta(\mathbf{X})$ and $N=200$.
-In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
+In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
 In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
 Figure S.2.8: Simulation results under continuous outcome, nonlinear $\delta(\mathbf{X})$ and $N=500$.
-In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
+In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
 In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
 Figure S.2.9: Simulation results under binary outcome, linear $\delta(\mathbf{X})$ and $N=40$.
-In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
+In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
 In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
 Figure S.2.10: Simulation results under binary outcome, linear $\delta(\mathbf{X})$ and $N=100$.
-In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
+In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
 In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
 Figure S.2.11: Simulation results under binary outcome, linear $\delta(\mathbf{X})$ and $N=200$.
-In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
+In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
 In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
 Figure S.2.12: Simulation results under binary outcome, linear $\delta(\mathbf{X})$ and $N=500$.
-In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
+In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
 In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
 Figure S.2.13: Simulation results under binary outcome, nonlinear $\delta(\mathbf{X})$ and $N=40$.
-In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
+In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
 In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
 Figure S.2.14: Simulation results under binary outcome, nonlinear $\delta(\mathbf{X})$ and $N=100$.
-In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
+In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
 In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
 Figure S.2.15: Simulation results under binary outcome, nonlinear $\delta(\mathbf{X})$ and $N=200$.
-In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
+In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
 In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
 Figure S.2.16: Simulation results under binary outcome, nonlinear $\delta(\mathbf{X})$ and $N=500$.
-In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
+In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels.
 In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.

### v6

@@
 Incorporating covariates into the analysis can enhance both the interpretability and efficiency of RCT results.
-While covariate balance can be pursued during both the design (Fisher, 1926; Taves, 1974; Pocock and Simon, 1975; Rosenberger and Sverdlov, 2008; Hu and Hu, 2012; Hu et al., 2014; Morgan and Rubin, 2012; Li et al., 2018b; Zhao and Ding, 2024) and analysis phases of an RCT, this paper focuses on methods for covariate adjustment during the analysis of already-conducted trials.
+While covariate balance can be pursued during both the design (Fisher, 1926; Taves, 1974; Pocock and Simon, 1975; Rosenberger and Sverdlov, 2008; Hu and Hu, 2012; Hu et al., 2014; Morgan and Rubin, 2012; Li et al., 2018b; Zhao and Ding, 2024b) and analysis phases of an RCT, this paper focuses on methods for covariate adjustment during the analysis of already-conducted trials.
 Regulatory agencies, including the U.S.
@@
 Several methods have been studied for data-driven variable selection procedures in RCTs.
-For example, the least absolute shrinkage and selection operator (Lasso) (Tibshirani, 1996) has been used in conjunction with ANHECOVA to select relevant covariates (Bloniarz et al., 2016; Schochet, 2022; Liu et al., 2023; Zhu et al., 2024).
+For example, the least absolute shrinkage and selection operator (Lasso) (Tibshirani, 1996) has been used in conjunction with ANHECOVA to select relevant covariates (Bloniarz et al., 2016; Schochet, 2022; Liu et al., 2023; Zhu et al., 2025).
 Another approach involves selecting the top $k$ covariates with the highest marginal correlations with the outcome, where $k<p$ (Senn, 1994).
-Alternatively, covariates can be screened using preliminary significance tests, such as two-sample $t$-tests or $z$-tests, to identify those most strongly associated with the treatment assignment or outcome (Permutt, 1990; Zhao and Ding, 2023).
+Alternatively, covariates can be screened using preliminary significance tests, such as two-sample $t$-tests or $z$-tests, to identify those most strongly associated with the treatment assignment or outcome (Permutt, 1990; Zhao and Ding, 2024c).
 These methods have primarily focused on continuous outcomes and the application of ANCOVA and ANHECOVA estimators.
@@
 In the context of RCTs, where bias is controlled, covariate adjustment primarily aims to enhance efficiency.
-Therefore, consistent with previous research (Cho and Yang, 2023; Yang et al., 2020), we recommend selecting covariates that are strong prognostic factors for the outcome.
+Therefore, consistent with previous research (Cho and Yang, 2024; Yang et al., 2020), we recommend selecting covariates that are strong prognostic factors for the outcome.
 For binary or categorical outcomes, research on covariate adjustment has also made significant progress (Ye et al., 2023a; Moore and van der Laan, 2009; Jiang et al., 2017; Guo and Basse, 2023; Cohen and Fogarty, 2024), though variable selection has not been a primary focus in these studies.
@@
 The open-source Coadvise package includes detailed documentation and reproducible examples, ensuring accessibility for a broad audience.
-Variable selection Lasso (Tibshirani, 1996), Adaptive Lasso (Zou, 2006), Marginal correlation (Permutt, 1990), Preliminary testing (Senn, 1994; Zhao and Ding, 2023)
+Variable selection Lasso (Tibshirani, 1996), Adaptive Lasso (Zou, 2006), Marginal correlation (Permutt, 1990), Preliminary testing (Senn, 1994; Zhao and Ding, 2024c)
 Covariate-adjusted estimator Model-based variance estimator (see Section 2) Robust to model misspecification (see Section 2) Nonlinear adjustment (see Section 3) Efficiency gain guarantee (see Section 3)
@@
 3.2 Theoretical results
-We establish theoretical results for efficiency gains via the AIPW estimator, with detailed proofs provided in Web Appendix A of our Online Supplemental Material.
+We establish theoretical results for efficiency gains via the AIPW estimator, with detailed proofs provided in Web Appendix A of Online Supplemental Material.
 The key results are summarized in the following Theorem 1.
@@
 For example, with positive-valued outcomes, conditions (a) and (b) can be simplified to $0\leq g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})\leq 2\mathbb{E}(Y(a)\mid\mathbf{X})$, which is reasonable for some positive categorical outcomes, such as binary outcomes modeled by logistic regression.
-Further illustration and discussion of these conditions are provided in Web Appendix A of our Online Supplemental Material.
+Further illustration and discussion of these conditions are provided in Web Appendix A of Online Supplemental Material.
 Comparing (i) and (ii), linear models with OLS ensure efficiency gains without additional bias constraints.
@@
 In our framework, variable selection is applied only to the outcome models used in the covariate-adjusted estimators.
-For all estimators considered—ANCOVA, ANHECOVA, and AIPW with Lasso-based variable selection—Condition 1 ensures that the residuals in their HW sandwich or model-based variance estimators capture the unexplained variation, including uncertainty due to variable selection (Liu et al., 2023; Bloniarz et al., 2016; Zhu et al., 2024).
+For all estimators considered—ANCOVA, ANHECOVA, and AIPW with Lasso-based variable selection—Condition 1 ensures that the residuals in their HW sandwich or model-based variance estimators capture the unexplained variation, including uncertainty due to variable selection (Liu et al., 2023; Bloniarz et al., 2016; Zhu et al., 2025).
 As a result, the robust variance estimators constructed from these residuals implicitly account for this uncertainty.
@@
 for $a=0,1$, where $\epsilon\sim\mathcal{N}(0,1)$ i.i.d., and $\delta(\mathbf{X})$ characterizes the individual treatment effect.
-The details of $\boldsymbol{\beta}_{0}$, different cases of $\delta(\mathbf{X})$ (linear and nonlinear), and the distributions used for $\mathbf{X}$ and $\mathbf{V}$ can be found in our Online Supplemental Material (Section B) for reproducibility.
+The details of $\boldsymbol{\beta}_{0}$, different cases of $\delta(\mathbf{X})$ (linear and nonlinear), and the distributions used for $\mathbf{X}$ and $\mathbf{V}$ can be found in Web Appendix B of Online Supplemental Material for reproducibility.
 4.2 Methods for comparison
@@
 where the test is rejected, with $\alpha=0.05,0.10$
-(Senn, 1994; Zhao and Ding, 2023)
+(Senn, 1994; Zhao and Ding, 2024c)
 4.3 Performance criteria
@@
 For brevity, we present the simulation results for the nonlinear continuous and nonlinear outcomes under sample sizes of 40 and 200 in this section.
-Complete simulation results, including figures, are available in our Online Supplemental Material (Section B).
+Complete simulation results, including figures, are available in Web Appendix B of Online Supplemental Material.
 Nonlinear continuous outcome Refer to caption
@@
 Interestingly, across both sample sizes, the Simple and Pre-test (both $\alpha=0.05$ and $0.1$) estimators result in the lowest power.
-This finding for the Pre-test estimator is consistent with Zhao and Ding (2023), where they advise against its use in practice based on theory and simulations.
+This finding for the Pre-test estimator is consistent with Zhao and Ding (2024c), where they advise against its use in practice based on theory and simulations.
 Lasso and adaptive Lasso, when combined with all estimators, yield satisfactory CP% under the continuous outcome with $N=200$.
-However, under the binary outcome setting with $N=200$, they exhibit under-coverage when used with the AIPW estimator, though this performance improves when the sample size increases to $N=500$, as shown in our Online Supplemental Material.
+However, under the binary outcome setting with $N=200$, they exhibit under-coverage when used with the AIPW estimator, though this performance improves when the sample size increases to $N=500$, as shown in Web Appendix B of Online Supplemental Material.
 This empirical finding suggests that, for binary outcomes using logistic regression as the outcome model in AIPW, it is valuable to investigate more robust variance estimation methods for smaller sample sizes in future research.
@@
 Finally, to empirically assess the asymptotic normality of post-selection estimators following Lasso and adaptive Lasso, we present histograms of AIPW estimates after variable selection in Figure 4.
-Additional results for ANCOVA, ANHECOVA, and other simulation scenarios are provided in the Online Supplemental Material.
+Additional results for ANCOVA, ANHECOVA, and other simulation scenarios are provided in Web Appendix B of Online Supplemental Material.
 As shown in Figure 4, the empirical distributions closely align with the corresponding theoretical normal densities (indicated by red dashed lines).
@@
 Moreover, although logistic regression is commonly used for propensity score estimation (Thomas et al., 2020; Hirano and Imbens, 2001; Connors et al., 1996) it remains uncertain whether it consistently enhances efficiency or if alternative methods with clearer regularity conditions might be preferable.
-Another potential extension involves prognostic covariate adjustment (PROCOVA), which employs machine learning techniques to generate prognostic scores based on baseline characteristics (Harrell Jr et al., 1984, 1996; Vanderbeek et al., 2022; Rubin and Thomas, 2000; Schuler et al., 2021; Li et al., 2024).
+Another potential extension involves prognostic covariate adjustment (PROCOVA), which employs machine learning techniques to generate prognostic scores based on baseline characteristics (Harrell Jr et al., 1984, 1996; Vanderbeek et al., 2022; Rubin and Thomas, 2000; Li et al., 2024).
 Future work could extend our variable selection methods to PROCOVA or explore the joint use of propensity and prognostic scores for covariate adjustment (Yang and Zhang, 2023; Zhang et al., 2022).
 Furthermore, our framework can be extended to a wide range of more general and complex settings.
-These include more general covariate-adaptive randomization (Bannick et al., 2025; Kernan et al., 1999; Broglio, 2018; Zhu et al., 2024; Yu et al., 2024), rerandomization (Zhu and Liu, 2023; Morgan and Rubin, 2012), time-to-event outcomes (Hua et al., 2024; Yi et al., 2020; Benkeser et al., 2021), restricted mean survival times (Han, 2023; Zhong et al., 2022; Karrison and Kocherginsky, 2018; Hanada et al., 2024), cluster-randomized trials (Murray, 1998; Donner et al., 2000; Stephens et al., 2012; Zhou et al., 2022; Balzer et al., 2023), and ordinal or composite outcomes with novel estimands such as the win ratio or generalized odds ratio (Agresti, 1980; Pocock et al., 2012; Barnhart et al., 2024; Matsouaka, 2022; Oakes, 2016; Zhang and Wu, 2024), as well as the Mann-Whitney rank-based statistic (Follmann et al., 2020).
-Other promising directions include conformal inference on individual or heterogeneous treatment effects (Yang et al., 2024; Wang et al., 2024; Gao et al., 2024a), the integration of multi-source RCT and observational data (Zhuang et al., 2024; Lee et al., 2024b, 2023; Yang et al., 2023; Han et al., 2025; Liu et al., 2024a; Yang and Wang, 2022; Liu et al., 2025), synthetic controls (Abadie et al., 2015, 2010; Ben-Michael et al., 2021), and external controls (Li et al., 2023; Lee et al., 2024c; Gao et al., 2023; Wang et al., 2025a).
-Additionally, alternative strategies for estimating regression coefficients, such as debiased machine learning algorithms (Chernozhukov et al., 2018; Banik and Shao, 2022; Wang et al., 2025b; Carlsson et al., 2014), debiased regressions (Lu et al., 2023; Chang et al., 2024; Chiang et al., 2023), and linear calibration (Bannick et al., 2025; Guo and Basse, 2023; Cohen and Fogarty, 2024) are of interest for future investigation.
-Data and code availability statement
-The BestAIR trial data used in Section 5 is available upon reasonable request at https://sleepdata.org.
-The R package Coadvise implementing the methodology of this paper is available at https://github.com/yiliu1998/Coadvise.
-Acknowledgments
+These include more general covariate-adaptive randomization (Bannick et al., 2025; Kernan et al., 1999; Broglio, 2018; Zhu et al., 2025; Yu et al., 2024), rerandomization (Zhu and Liu, 2023; Morgan and Rubin, 2012), time-to-event outcomes (Hua et al., 2024; Yi et al., 2020; Benkeser et al., 2021), restricted mean survival times (Han, 2023; Zhong et al., 2022; Karrison and Kocherginsky, 2018; Hanada et al., 2024), cluster-randomized trials (Murray, 1998; Donner et al., 2000; Stephens et al., 2012; Zhou et al., 2022; Balzer et al., 2023), and ordinal or composite outcomes with novel estimands such as the win ratio or generalized odds ratio (Agresti, 1980; Pocock et al., 2012; Barnhart et al., 2025; Matsouaka, 2022; Oakes, 2016; Zhang and Wu, 2024), as well as the Mann-Whitney rank-based statistic (Follmann et al., 2020).
+Other promising directions include conformal inference on individual or heterogeneous treatment effects (Yang et al., 2024; Wang et al., 2024; Gao et al., 2024a), the integration of multi-source RCT and observational data (Zhuang et al., 2025; Lee et al., 2024b, 2023; Yang et al., 2023; Han et al., 2025; Liu et al., 2024a; Yang and Wang, 2022; Liu et al., 2025), synthetic controls (Abadie et al., 2015, 2010; Ben-Michael et al., 2021), and external controls (Li et al., 2023; Lee et al., 2024c; Gao et al., 2025; Wang et al., 2025a).
+Additionally, alternative strategies for estimating regression coefficients, such as debiased machine learning algorithms (Chernozhukov et al., 2018; Wang et al., 2025b; Carlsson et al., 2014), debiased regressions (Lu et al., 2023; Chang et al., 2024; Chiang et al., 2023), and linear calibration (Bannick et al., 2025; Guo and Basse, 2023; Cohen and Fogarty, 2024) are of interest for future investigation.
+Acknowledgment
 The authors are grateful to Marlena Bannick for insightful discussions during the early stages of this research endeavor.
+Funding
 Yi Liu was supported by the National Heart, Lung, and Blood Institute (NHLBI) of the National Institutes of Health (NIH) under Award Number T32HL079896.
@@
 Government.
+Data and Code Availability Statement
+The BestAIR trial data used in Section 5 is available upon reasonable request at https://sleepdata.org.
+The R package Coadvise implementing the methodology of this paper is available at https://github.com/yiliu1998/Coadvise.
 Appendix S Appendix: R Code and Tutorial to COADVISE
@@
 Inverse probability weighting (ipw; Robins et al. (1994)) $\times$ $\checkmark$ MAR
-Missingness indicator (missInd; Zhao et al. (2024); Zhao and Ding (2022); Song et al. (2021)) $\checkmark$ $\times$ MAR
+Missingness indicator (missInd; Zhao et al. (2024); Zhao and Ding (2024a); Song et al. (2021)) $\checkmark$ $\times$ MAR
 Table S.1: Summary of missing data imputation methods available in the current Coadvise R package.
@@
 var.sel.method Variable selection method $\times$ "Lasso" "No", "Lasso", "A.Lasso", "Corr.k", "Corr.xi", "Pre.test"
-lasso.family Family used in Lasso regression $\times$ "gaussian" "gaussian", "binomial", "poisson", "multinomial" (Same as the glmnet package (Friedman et al., 2021))
+lasso.family Family used in Lasso regression $\times$ "gaussian" "gaussian", "binomial", "poisson", "multinomial" (Same as the glmnet package)
 A.lasso.family Family used in adaptive Lasso regression $\times$ "gaussian" "gaussian", "binomial", "poisson", "multinomial"
@@
 We also need to specify the model family via lasso.family if "Lasso" is used.
-The default is "gaussian" for using linear model with Gaussian random error assumption in Lasso regression, but users can choose other families, such as "binomial" for binary outcome if using logistic regression, and any families provided in the glmnet package.Friedman et al.
-(2021); Hastie et al.
-(2021) The arguments out0.model.aipw and out1.model.aipw are for specifying the link functions used in outcome models $\mu_{0}(\mathbf{X})$ and $\mu_{1}(\mathbf{X})$, respectively, in the AIPW estimator.
+The default is "gaussian" for using linear model with Gaussian random error assumption in Lasso regression, but users can choose other families, such as "binomial" for binary outcome if using logistic regression, and any families provided in the glmnet package (Friedman et al., 2021; Hastie et al., 2021).
+The arguments out0.model.aipw and out1.model.aipw are for specifying the link functions used in outcome models $\mu_{0}(\mathbf{X})$ and $\mu_{1}(\mathbf{X})$, respectively, in the AIPW estimator.
 For example, in the binary outcome case, we use "logit" to specify logistic regression models for both potential outcomes.
@@
 Nevertheless, we can still improve efficiency through “appropriate augmentation,” even without precise model specification, compared to $\widehat{\tau}_{\text{simple}}$.
-This is our objective in leveraging covariate information to gain efficiency (Bannick et al., 2023).
+This is our objective in leveraging covariate information to gain efficiency (Bannick et al., 2025).
 In practice, we focus on a subclass of $\mathcal{G}$, denoted as $\mathcal{G}_{\boldsymbol{\beta}}^{g}\subset\mathcal{G}$, where $g_{a}(\mathbf{X})=g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})$ follows a generalized linear model (GLM) for $Y(a)$, for $a=0,1$.